A Complete Bibliography of the Journal of the Royal Statistical Society, Series A family: 1950–1959

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

09 November 2023 Version 1.02

Title word cross-reference

0 ≤ x ≤ 1 [396]. 100 [447]. 2 × 2 [728, 465]. 50 [447]. ^{x} [448]. C_{n}(x) [396]. K_{0}(x) [396]. K_{1}(x) [396]. n = 26 [204]. S_{n}(x) [396]. sinh^{−1} \sqrt{x} [24]. t [256, 872]. X [371]. x = 10 [672, 673]. x = 2 [672]. x = 2.5 [673]. Y_{0}(x) [396]. Y_{1}(x) [396].

-Distribution [872]. -Tests [256].

1 [69, 177, 160]. 125th [1020]. 13 [237]. 17th [1023]. 1869 [734]. 18th [915, 967]. 1900 [496]. 1919 [530]. 1930’s [451]. 1938 [346, 114]. 1939 [83]. 1941 [679]. 1945 [8]. 1946 [553]. 1947 [119]. 1948 [190, 49, 86]. 1948/1949 [190]. 1949 [190, 575, 812, 343, 33, 289, 143]. 1950 [69, 39, 678, 758, 186, 449, 533, 756, 347]. 1951 [650, 786, 344, 349]. 1952 [902, 736, 490, 554, 342]. 1953 [527, 484, 705]. 1954 [843, 983, 818]. 1955 [809]. 1956 [683]. 1957 [763, 779, 954]. 1959 [1023]. 19th [803].

2 [525]. 20th [697]. 21st [763, 70, 69]. 22nd [590, 39]. 23rd [483]. 24th [386]. 25 [207]. 25-Årsmeldinger [207]. 26th [279]. 27th [178, 177, 229]. 28th [413].

31st [69, 177].

41 [431, 162]. 42 [452]. 43 [211]. 49 [627].

50 [532]. 51 [576]. 52 [757, 529]. 54 [903, 571]. 57 [813].

6th [763].

75 [30].

A. [502]. År [207]. Årsmeldinger [207]. Abdel [569]. Abdel-Aty [569]. Abel [649]. Abel-Smith [649]. Abilities [81, 80, 799]. Ability [80]. Abraham [134, 596]. Abramovitz [815]. Abramowitz [145]. Abrams [235, 87]. Abstract [115]. Acceptance [742, 140]. Accident [247]. Accidents [247, 692]. Account [492]. Accountants [752]. Accounting [800, 399, 189]. Accounts [736, 450, 782, 784]. Accuracy [164]. Acheson [284, 1027]. Ackoff [778]. Action [235]. Active [290, 285]. Activity [496]. Actuaries [683, 342]. Adams [530, 645, 711]. Addendum [571]. Adding [801, 722]. Adding-Up [801, 722]. Addison [702]. Additions [16, 37, 59, 91, 125, 152, 170, 194, 219, 243, 268, 297, 325, 354, 376, 406, 436, 458, 474, 504, 539, 560, 580, 608, 635, 658, 686, 714, 740, 765, 792, 824, 858, 883, 909, 935, 961, 989, 1016, 1041]. Adler [760]. Administration [382, 809, 275]. Admissions [679]. Advanced [942, 307]. Advances [176, 84]. Advertising [227, 490]. Affairs [186, 262]. Africa [410]. Agatha [346]. Age [755]. Aggregate [468]. Aggregation [566]. Aging [879]. Agrarian [319]. Agrégation [920]. Agricultural [787, 564, 99, 332, 693, 870]. Agriculture [532, 98, 744, 118, 366]. Air [383]. Airy [951]. Aitchison [805]. Alan [927, 264]. Albert [758, 147, 368]. Alder [318]. Alexander [79]. Algebra [391]. Allais [648]. Allan [627, 1000]. Allen [699, 1030, 774]. Allocation [975]. Alternative [800]. Alva [852]. America [117, 52]. Americaine [648]. American [628, 734, 655, 51, 262, 430, 846, 817].

Analyses [664]. Analysis [247, 876, 772, 1044, 1007, 443, 281, 515, 563, 795, 418, 1033, 369, 949, 83, 640, 388, 842, 102, 302, 484, 485, 614, 512, 562, 931, 847, 797, 1006, 991, 520, 727, 231, 585, 395, 834, 423, 252, 916, 944, 945, 314, 917, 918, 671, 182, 556, 337, 201, 430, 868, 647, 750, 102, 853]. Analytical [833]. Anderson [918]. Andrew [550]. Anne [453]. Anniversary [1020]. Annotated [345]. Annual [1022, 575, 70, 69, 178, 177, 279, 386, 483, 590, 697, 803, 915, 1023, 705, 843, 143]. Antilogarithms [448]. Appleyard [753]. Applicable [598]. Applicances [190]. Application [255, 809, 768, 638, 591, 318, 974, 668, 285]. Applications [1021, 1025, 259, 309, 797, 157, 333, 1030, 889, 593, 308, 133, 46, 520, 486, 77]. Applied [1007, 399, 727, 946]. Appraisal [980]. Approach [640, 699]. Approaches [319]. April [349, 344]. Archives [344]. Arctan [371]. Area [745, 904, 430, 292]. Areas [50]. Arguments [952, 448, 569, 521]. Arithmetic [973]. Arley [45]. Armsen [728]. Army [462]. Arne [598, 107, 108]. Arnoff [778]. Arrow [997, 1033, 948, 283]. Art [260]. Arthur [763, 310, 470, 731, 206, 987, 711]. Articles [15, 36, 58, 90, 124, 151, 169, 193, 218, 242, 267, 296, 324, 353, 375, 405, 435, 457, 473, 503, 538, 559, 579, 607, 634, 657, 685, 713, 739, 764, 791, 823, 857, 882, 908, 934, 960, 988, 1015, 1040]. Asia [117]. Asking [260]. Aspects [613, 844, 917, 638, 258]. Assay [306]. Assess [1035]. Assessment [586]. Asset [812]. Assets [719, 541]. Associate [808]. Associated [886, 363]. Assumption [694]. Assurance [372]. Astronomy [836, 364]. Atkinson [812]. Atlantic [493]. Atmospheric [796, 797]. Attempts [1035]. Aty [569]. Audience [662]. Auditing [188]. Auditors [752]. Augustus [322]. Australia [532]. Authorities [450]. Authors [770].

Automatic [923, 807, 397]. Automation [1031]. Available [545, 449]. Avery [711]. Axiomes [648].

B [109, 442, 202, 876, 997, 81, 421, 464, 976, 469, 13, 204, 788, 102, 232, 213, 549, 616, 853, 533, 209, 567, 419, 395, 838, 336, 629]. B. [537]. Back [18, 60, 93, 221, 244, 269, 299, 327, 355, 378, 407, 659, 688, 716, 860]. Backwardness [900]. Bailey [866, 783, 1024]. Balance [184, 3, 85]. Balanced [808]. Balances [497]. Balances-of-Payments [497]. Bancroft [548]. Band [795]. Banking [643]. Barger [734]. Barker [706]. Bartlett [591]. Base [448]. Based [247]. Basic [645]. Basis [644, 819]. Batchelor [345]. Battin [807]. Bauer [931]. Beach [897]. Bearing [514]. Beer [137]. before [701, 39]. Beginnings [969]. Behavior [785, 702]. Behavioral [547, 748]. Behaviour [530]. Bell [466]. Belonging [863]. Benjamin [679, 876, 976]. Berkeley [280, 836]. Bernard [618]. Bernardo [234]. Berry [1030]. Bessel [341, 521, 396, 465]. Bethlem [650, 903]. Between [832, 769, 479, 448, 906, 952, 331]. Bibliography [176, 165, 954, 157, 333, 345, 781]. Bickley [465]. Bill [360]. Binomial [521, 674, 139, 447]. Bioassay [392]. Biological [137, 1025, 464, 306, 1010, 104, 870, 103]. Biologique [76]. Biology [836, 839, 548, 1024]. Biomathematics. [464]. Biometric [307]. Biometrika [518]. Biometry [583, 522]. Bizley [806]. Blacker [650, 903, 350]. Blackwell [516]. Blank [814]. Bliss [522, 392]. Block [808]. Board [783]. Boddington [255]. Bond [984]. Book [75, 109, 759, 554, 653, 678, 529, 785, 787, 814, 982, 136, 442, 594, 212, 137, 141, 165, 202, 252, 256, 365, 392, 522, 597, 968, 27, 32, 103, 591, 644, 724, 811, 650, 679, 810, 844, 851, 876, 903, 927, 1008, 1025, 417, 772, 866, 916, 944, 997, 113, 233, 286, 287, 290, 350, 571, 572, 649, 782, 845, 873, 877, 923, 978, 1037, 953, 954, 400, 784, 848, 203, 338, 420, 547, 670, 730, 929, 346, 530, 981, 666, 1009, 487, 921, 969, 239, 573, 255, 280, 284, 310, 370, 623, 211, 238, 52]. Book [73, 81, 364, 757, 816, 846, 875, 880, 983, 184, 189, 523, 646, 849, 425, 427, 628, 682, 815, 1007, 264, 754, 491, 318, 443, 698, 752, 928, 1002, 897, 984, 185, 234, 281, 319, 421, 428, 464, 467, 488, 515, 517, 553, 625, 755, 758, 867, 879, 889, 891, 900, 947, 956, 976, 1026, 1034, 893, 971, 398, 190, 210, 147, 316, 528, 101, 180, 181, 621, 138, 148, 469, 575, 627, 704, 706, 812, 813, 819, 898, 925, 930, 401, 492, 974, 805, 972, 53, 345, 418, 945, 116, 372, 683, 76, 551, 1033, 598, 576, 30, 550, 306, 369]. Book [399, 453, 452, 470, 557, 574, 489, 527, 111, 230, 546, 648, 725, 869, 753, 1010, 315, 566, 707, 343, 431, 532, 445, 896, 949, 112, 786, 11, 13, 31, 79, 83, 84, 957, 552, 568, 726, 204, 257, 259, 314, 341, 342, 924, 680, 878, 253, 388, 620, 747, 737, 776, 778, 842, 895, 1000, 788, 102, 104, 183, 498, 78, 970, 922, 144, 232, 394, 592, 593, 645, 700, 773, 809, 836, 890, 892, 917, 918, 213, 308, 8, 9, 71, 119, 208, 235, 260, 107, 108, 133, 393, 484, 485, 749, 188, 734, 142, 288, 920, 187]. Book [282, 902, 736, 363, 466, 134, 251, 307, 339, 389, 415, 549, 596, 616, 699, 746, 804, 835, 943, 841, 853, 979, 450, 526, 1004, 655, 677, 10, 28, 33, 45, 46, 72, 77, 451, 996, 1005, 931, 619, 701, 705, 806, 843, 901, 904, 905, 977, 1036, 490, 760, 236, 667, 669, 735, 837, 852, 946, 973, 531, 820, 600, 671, 807, 100, 182, 283, 444, 106, 347, 1043, 48, 49, 50, 51, 668, 140, 309, 311, 289, 468, 847, 117, 143, 160, 186, 261, 262, 285, 449, 493, 495, 555, 556, 681, 624, 47, 82, 214, 595, 337, 54]. Book [74, 397, 980, 906, 162, 781, 201, 703, 263, 654, 340, 366, 533, 114, 601, 149, 402, 207, 429, 783, 626, 733, 209, 292, 146, 258, 426, 164, 237, 291, 317, 494, 702, 774, 899, 817, 12, 80, 85, 86, 548, 1024, 497, 163, 430, 496, 731, 850, 161, 348, 349, 1003, 105, 567, 839, 840, 865, 868, 894, 1035, 926, 732, 1006, 312, 524, 367, 115, 651, 652, 756, 955, 1011, 118, 779, 368, 390, 391, 419, 519, 520, 521, 647, 672, 673, 674, 675, 676, 727, 728, 750, 777, 808, 870, 871, 872, 950, 951, 952, 998, 999]. Book [1030, 1031, 874, 29, 135, 179, 205, 206, 231, 416, 516, 525, 617, 751, 775, 888, 948, 975, 1028, 1029, 1032, 422, 87, 518, 942, 818, 748, 110, 344, 395, 454, 599, 729, 780, 1001, 834, 838, 919, 139, 254, 313, 336, 371, 396, 423, 424, 446, 447, 448, 465, 486, 569, 570, 618, 622, 145, 629, 1027, 806, 135]. Books [706]. Booth [397, 923, 550, 397]. Borstal [623]. Bound [214]. Bourne [13, 13]. Bowden [419]. Bowley [763, 310]. Bowman [853]. Braddock [984]. Bradford [47, 595]. Braithwaite [442, 616]. Bray [782]. Brian [649]. Brinley [929, 493]. Bristol [454]. Britain [553, 627, 493, 906, 533, 44, 437, 912, 156, 41, 911, 276, 223, 561, 113, 815, 489, 874]. British [81, 698, 210, 53, 737, 1004, 54, 850, 198, 719, 158, 228, 831, 678, 526, 451, 555]. Brockmeyer [107]. Bronchitis [877]. Brooke [902]. Brookes [202, 567, 567, 336]. Brooks [393]. Brown [805, 835]. Buch [45]. Buckingham [922]. Buckland [779]. Budgets [691, 723, 197, 847]. Bulletin [143]. Burchardt [986]. Bureau [185, 704, 783, 570]. Burn [103, 1004]. Burns [470]. Bush [617]. Business [2, 32, 147, 640, 144, 760, 236, 852, 263, 317, 494, 1003, 777, 1028, 145, 203, 369]. Butterbaugh [165]. Butters [32].

C [75, 202, 392, 522, 650, 903, 286, 350, 873, 670, 846, 928, 421, 464, 488, 1034, 813, 805, 399, 452, 431, 896, 84, 778, 842, 394, 393, 307, 339, 526, 1005, 236, 669, 309, 311, 209, 146, 258, 426, 702, 80, 567, 868, 367, 521, 675, 775, 87, 729, 254, 336, 424]. C.B [39]. C.B. [168]. C.B.E [217]. C.H. [537]. Cahiers [707].

Cairncross [492]. Calcolo [78]. Calcul [77, 181]. Calculating [111]. Calculation [735]. Calculations [156, 370]. Calculator [923]. Calculators [397]. Calculus [567]. Calhoun [522]. Callbacks [477]. Calvin [487]. Cambridge [424]. Campion [39]. Canada [86]. Canadian [652]. Cancer [1035, 927]. Canning [777]. Cans [272]. Capital [529, 238, 52, 508, 402, 704]. Caradog [928]. Card [994, 864]. Cardiff [906]. Cardpack [592, 74]. Cards [259, 1030]. Carlo [726]. Carnap [251]. Carr [928]. Carr-Saunders [928]. Carrier [449]. Carruthers [393]. Cars [830, 438]. Carter [526]. Cartter [627]. Case [863]. Casey [259, 1030]. Cassels [430]. Cattell [395]. Causal [612]. Causes [9]. Cecil [987]. Census [185, 553, 142, 285, 980, 343, 1011]. Censuses [600, 261]. Central [627, 543, 872]. Ceylon [880]. Chained [479]. Chair [39]. Chaloner [678, 451]. Chambers [183]. Chance [75, 725, 592, 74]. Chang [184]. Change [494]. Changes [743, 885, 114]. Changing [982]. Chapman [346]. Chapters [775]. Characteristics [861, 983, 758]. Charles [113, 288, 599, 633]. Charnes [444]. Chart [629]. Charting [420]. Chayes [750]. Chebyshev [396]. Cheeseman [83, 84]. Chemical [137, 398, 132]. Chemistry [839]. Chemists [253]. Chenery [997]. Chernoff [1026]. Chessboard [592, 74]. Chester [210]. Chew [947]. Childbirth [737]. Children [756, 81]. Chintschin [205]. Choice [725, 592, 283, 74]. Choix [648]. Chorafas [998]. Chronic [877]. Chun [184]. Chung [232, 549]. Churchman [778]. CIBA [729]. Cinema [437]. Cinema-Going [437]. Cinemas [437]. Circle [387]. Clarence [982]. Clark [956, 148, 394]. Clarke [113]. Class [863, 289]. Classes [808]. Classical [775]. Classification [861, 9]. Clatworthy [808]. Clay [666, 501]. Clifford [396]. Clinical [586]. Clothing [249]. Clustering [477]. Clyde [599]. Co [26]. Co-Operative [26]. Coal [957, 768, 41]. Coastwise [214]. Cobb [599]. Cochran [867, 71]. Cocoa [239]. Coe [372]. Coefficients [519, 521]. Coffee [239]. Cohen [646]. Coking [768]. Cole [723]. Colerus [567]. Colin [148]. Collected [83]. Colliery [584]. Colombo [234]. Colonial [53, 39]. Colour [743]. Columbia [737]. Columns [257]. Combination [892]. Combinatorial [944]. Comment [382]. Commentary [875]. Commerce [682, 255]. Commercial [147]. Commission [97, 10, 186, 343]. Committee [704, 342, 1036, 940, 349]. Committees [481]. Communication [417]. Communities [758]. Community [720, 769]. Companies [116]. Company [941, 820]. Comparative [703]. Comparison [990, 304, 784, 469, 832, 154]. Competition [1032]. Compilation [467]. Compleat [552]. Complementi [78]. Complex [569]. Complications [737]. Comportant [648]. Comprehensive [954]. Computation [1031]. Computer [965, 841, 570]. Computing [257, 342, 419]. Comrie [123]. Conan [292]. Concentration [898, 640]. Concentrazione [749]. Concept [247]. Concepts [645]. Concerning [741]. Conclusion [967]. Conditions [928, 148, 98]. Conference [848, 704, 527, 481, 51, 291, 161, 510, 187, 263, 237, 317]. Conferences [313]. Confidence [232]. Congestion [563]. Congress [809, 51, 695]. Construction [735, 357]. Consumer [109, 530, 530, 783]. Consumption [849, 156, 347, 468, 246, 884]. Contiguity [225]. Contingency [728, 465]. Continuation [80]. Continuing [819]. Continuity [694]. Contribution [3, 179]. Contributions [972, 950, 808, 416]. Control [165, 284, 138, 930, 401, 1000, 999, 1031, 1027, 776, 807, 524, 629]. Controle [422]. Controversy [1035]. Convened [967]. Conversion [676]. Coontz [845]. Cooper [444]. Coppinger [733]. Corlett [530]. Cornell [700]. Corporate [211, 984]. Corrado [749]. Correlated [675]. Correlation [621, 100]. Correlational [945]. Cosecants [519]. Cosine [569]. Cosines [672]. Cost [719, 286, 649, 273, 885, 114, 692]. Costs [842]. Cotton [262, 358, 5]. Council [69, 177, 278, 385, 482, 589, 696, 802, 914, 1022]. Countries [52, 983, 832, 261, 744, 931, 402, 262]. County [926]. Course [916, 921, 969, 72, 946, 344, 1043]. Courts [701]. Covariance [664]. Cowden [776]. Cox [141, 968, 571, 867, 71]. Cramér [593]. Crawford [532]. Credibility [865]. Credit [783]. Crew [147]. Criminal [361]. Criminology [128]. Criterion [382, 801, 275, 722]. Critical [813, 888]. Criticism [247]. Critique [648]. Croxton [1025]. Crump [209]. Cumulate [674]. Currencies [469]. Current [56, 35, 89, 122, 167, 192, 216, 241, 266, 295, 321, 352, 374, 404, 433, 456, 472, 500, 535, 578, 603, 631, 709, 762, 790, 822, 855, 933, 959, 1013, 1039, 510]. Currently [825]. Curtis [399]. Curves [990]. Cushion [977]. Cuthbert [851]. Cycle [452]. Cycles [144, 236, 494, 145, 263]. Cyclical [184]. Cyril [337]. Cyrus [1043].

D [787, 982, 252, 522, 968, 32, 103, 644, 876, 877, 203, 322, 530, 239, 928, 553, 956, 1034, 210, 528, 550, 306, 111, 753, 552, 204, 314, 1000, 232, 994, 288, 902, 466, 804, 841, 526, 1005, 977, 397, 366, 783, 774, 231, 424]. D. [994]. D.Sc [710]. Dahlberg [731]. Dalenius [972]. Daniel [547]. dans [76]. Darrell [498]. Data [927, 83, 767, 285, 777, 612, 399, 980, 999, 423]. Data. [520]. David [814, 516, 465, 365]. Davidson [683]. Davies [515]. Davis [212]. Day [95, 967]. DC [570]. Deadly [225].

Dean [1031]. Death [9]. Debate [813, 1034]. Debt [146]. December [69, 177, 967]. Decennial [754, 904, 905]. Decimal [952, 448]. Decision [1026, 924, 134]. Decisions [888, 1028, 516]. d’Econometrie [707]. Definition [495]. degli [622]. Degree [519]. Delens [110]. Delivered [344]. Della [234]. DeLury [204, 232]. Demand [814, 992, 438, 940, 817, 20, 780, 443]. Demands [1033]. Deming [101]. Demographers [185]. Demographic [572, 49, 53]. Demographie [428]. Demography [571, 880, 732, 141]. Densities [175]. Department [186, 262, 207]. Dependencies [850]. Depopulation [786]. Depreciation [830]. Derivation [664, 575, 583]. Derman [1043]. Desai [431]. Descending [673]. Descriptive [11]. Design [886, 644, 281, 893, 388, 102, 104, 668, 330, 231, 515, 418]. Designs [947, 808, 871, 867, 102, 71]. Determinants [963, 652]. Determination [950]. Developed [52, 425, 744]. Development [81, 911, 117, 131, 1018, 231, 570, 572, 12, 1006]. Developments [571, 829]. Deviates [675, 568]. Devices [342]. Devons [698, 149]. Diagnosis [965]. Dick [202]. Dictionary [779]. Diego [622]. Diet [289]. Difference [1001]. Differences [226, 129]. Different [22]. Differential [257, 863, 390]. Differentials [787]. Digital [923, 397, 419, 841]. Digits [568]. Dimitris [998]. Discussion [97, 542, 798, 767, 39]. Disease [679]. Diseases [769, 9]. Distribution [109, 805, 480, 598, 939, 114, 942, 876, 232, 872, 734, 674, 139]. Distributions [583, 549, 770]. Divided [512]. Dividend [941]. Division [186]. Dixon [201]. Dobrovolsky [211]. Doctor [95]. Domestic [190, 347]. Donald [233, 146, 312]. Doob [389]. Dornbusch [669]. Dorothy [723]. Dr. [536]. Drei [205]. DuBois [945]. Dudding [629]. Dudley [758, 776, 114]. Duncan [758, 1004, 888, 284, 1027]. d’une [648]. Durée [290]. During [83, 743, 10, 236, 571, 52]. Dutch [652]. Dutch-Canadian [652]. Dynamic [730]. Dynamics [144]. DYSEAC [570].

Earnings [360, 98]. East [117]. Ecole [648]. Ecological [624]. Econometrics [197, 836, 443, 340]. Econometriques [707]. Economic [15, 36, 58, 90, 124, 151, 169, 193, 218, 242, 267, 296, 324, 353, 375, 405, 435, 457, 473, 503, 538, 559, 579, 607, 634, 657, 685, 713, 739, 764, 791, 823, 857, 882, 908, 934, 960, 988, 1015, 1040, 851, 572, 845, 784, 861, 573, 52, 829, 425, 698, 879, 900, 148, 704, 813, 453, 470, 557, 574, 566, 144, 302, 414, 526, 852, 117, 186, 262, 493, 783, 426, 164, 494, 702, 983, 897, 900, 813, 898, 931, 973, 106, 117, 162, 494]. Economic-Demographic [572]. Economical [742]. Economically [285]. Economics [203, 929, 1007, 264, 528, 369, 157, 333, 1018, 1028, 981, 1002, 805, 598, 315, 774, 899, 1037, 1009, 147]. Économiquement [290]. Economists [946]. Economy [198, 492, 512, 734, 736, 130, 210, 493]. Edged [357]. Edrick [524]. Education [463, 330, 418, 105]. Educational [81, 700]. Edward [138, 898, 369, 736]. Edwards [101, 759]. Edwin [924]. Effects [627]. Efficiency [382, 766, 957, 994, 864, 275, 1005, 979, 1006]. Egmont [567]. Eigenvalues [950]. Eighteenth [279]. Eileen [902]. Elaine [729]. Eldridge [30]. Electrical [466]. Electricity [66, 156]. Electronic [965, 767, 777]. Elementary [338, 310, 424, 1025, 666, 1026, 852, 312, 254]. Elements [593, 488, 976, 311]. Eliasberg [815]. Elizabeth [400]. Ely [149]. Emphasis [1006]. Empire [53]. Empirical [924]. Employment [815, 329, 98, 276, 982].

End [612]. Ended [69, 177]. Ending [69]. Engineering [809, 308, 46, 871, 466]. Engineers [1043]. England [876, 927, 649, 875, 754, 928, 786, 8, 904, 905, 1011]. English [926]. Enquiry [10, 874]. Environment [886]. Epidemics [866, 83]. Epidemiological [705, 843]. Epidemiology [31]. Epps [168]. Equations [950, 257, 390, 1001]. Equipment [767]. Eric [557]. Erik [348]. Erlang [107]. Ernie [1044, 991]. Error [865]. Errors [274, 359]. Esperimenti [622]. Essay [411]. Essays [494, 163]. Essential [979]. Essentials [700]. Estadistica.Part [525]. Estimates [912, 575, 223]. Estimating [691, 723, 744]. Eugene [318, 1031]. Europe [186, 601, 186]. European [681, 757]. Evelyn [318]. Everett [679]. Eversley [1034]. Everyman [567]. Evidence [28, 1035]. Examination [561]. Exceptional [332]. Exchange [832]. Exchanges [209]. Exchequer [753]. Exclusive [795]. Executives [760]. Exercise [274]. Exercises [895, 446]. Expansion [301]. Expenditure [554, 530, 691, 723, 753, 490, 347, 223, 6, 874]. Expenditures [190]. Experience [861, 984]. Experienced [226, 129]. Experiences [703]. Experiment [450, 871, 388]. Experimental [477, 381, 644, 867, 947, 893, 71, 668]. Experimentation [367, 76, 137, 398, 620]. Experiments [418, 611, 102, 231, 968, 281, 515]. Experts [261]. Explanation [442, 112]. Exponential [673]. Exports [52]. Exposition [897]. Extension [465]. Extent [130]. External [449]. Extrapolation [46]. Extrasensory [729]. Extreme [520, 423]. Extreme-Value [423]. Extremes [970]. Eysenck [523].

F [202, 365, 782, 203, 487, 364, 897, 418, 470, 949, 112, 257, 700, 393, 526, 1036, 669, 626, 733, 163, 1003, 567, 870, 999, 87, 465]. F. [986]. F.R.S. [217]. F.R.S.E. [710].

Fabricant [496]. Fabrications [422]. Facilitate [256]. Factor [484, 614, 395]. Factorial [871]. Factors [730, 941, 1010, 871]. Factory [842]. Faculty [683]. Families [350, 846]. Family [553, 197, 10, 847, 816]. Faning [878, 10]. Far [117]. Farm [460, 543, 23]. Fascination [840]. Faster [419]. Fats [43]. Federal [783]. Federer [668]. Fein [1009]. Feldman [599]. Felix [750]. Feller [889, 133]. Fellows [56, 913, 993]. Female [154]. Ferber [468]. Fertility [319, 553, 1034, 833, 10, 1011]. Festinger [547]. Fetter [853]. Field [384, 611, 620, 367]. Fields [227, 4, 43, 65, 441, 44, 99, 545, 643, 66, 174, 158, 588, 335, 228, 26, 305, 277, 463, 383, 361, 414, 67, 362, 514, 25, 42, 132, 64, 587, 249, 5, 565]. Fieller [675]. Fifth [1023]. Fifty [10]. Figures [519]. Filters [795]. Finance [317]. Financial [69, 177, 812, 813, 514, 163]. Findlay [606]. Finite [953]. Finney [103, 465, 252, 644, 306, 366]. First [590, 679, 921, 969, 878, 809, 481, 72, 1043, 48, 143, 349]. Fiscal [627]. Fisher [870, 594, 27, 724, 1035]. Five [350, 262, 952, 448]. Fixed [719, 541]. Fleming [839]. Flow [529]. Fluctuations [162, 144]. Fondements [648]. Food [305, 347, 6, 20]. Force [982]. Forces [81]. Ford [214]. Forecasting [573]. Forecasts [30]. Foreign [209]. Forestry [343]. Form [801, 722, 728]. Formation [704, 402]. Formula [833]. Formulae [465]. Formulas [308]. Fortegnelse [207]. Fortet [181]. Foundation [729]. Foundations [546, 251]. Fourth [915]. Fox [521]. Fractional [871]. Framework [316]. Franco [556]. Frank [201]. Frankel [425]. Franklin [999]. Fraser [804]. Fred [878]. Frederick [759, 1025, 949, 426, 617]. Freedom [411]. French [861]. Frequency [583]. Freudenthal [208]. Freund [1003, 338]. Friedman [849, 1029]. Front [1, 19, 38, 62, 94, 127, 153, 172, 196, 222, 245, 271, 300, 328, 356, 380, 409, 507, 610, 637, 660, 690, 718, 828, 938]. Frontiers [470]. Frowen [813]. Fruit [367]. Fryer [488]. Fuels [190]. Function [442, 20, 569, 849]. Functions [341, 41, 521, 396, 423, 465, 134, 468, 951]. Fundamental [105]. Fundamentals [1031, 189, 943]. Further [723, 950].

G [109, 103, 916, 877, 1037, 953, 969, 73, 491, 421, 517, 867, 621, 723, 576, 230, 725, 532, 620, 183, 700, 213, 71, 835, 450, 1036, 100, 311, 774, 899, 85, 651, 779, 777, 87, 942, 729, 446, 447, 465]. Gains [238]. Galenson [628]. Gambling [224, 646]. Games [616, 667, 416, 516, 339, 179, 1032, 888]. Gamma [569]. Gap [508]. Gas [66]. Gass [974]. Geiringer [869]. Gene [999]. General [70, 69, 178, 177, 279, 386, 483, 590, 697, 803, 915, 1023, 967, 875, 801, 754, 8, 863, 904, 905, 977, 722]. Generalization [833]. Generalized [569]. Genetic [894]. Geographic [50]. Geographical [876, 787]. Geometrical [664]. Geometry [638]. George [814, 747, 168, 783, 872, 217, 606]. Gerald [872]. Gerhard [340]. Gertrude [71]. Gilbert [653, 469, 429]. Gilt [357]. Gilt-Edged [357]. Gini [749]. Giorgio [290]. Girshick [516]. Giuseppe [78, 622]. Gjennom [207]. Glasgow [492]. Glass [553]. Glassey [81]. Glenday [856]. Glick [846]. Gnedenko [549]. Going [437]. Gold [303]. Goldberg [1001]. Goldsmith [575, 575]. Goncalves [856]. Good [28]. Goodman [810]. Goods [461, 109]. Goodwin [103]. Gore [650, 398]. Gottfried [263]. Goulden [337]. Government [545, 627, 496, 1006, 344]. Governmental [189]. Governments [682]. Governors [783]. Grabbe [1031]. Graduation [833]. Grant [165]. Grants [753]. Graphic [487]. Graphical [597]. Great [815, 553, 493, 874, 44, 437, 912, 156, 41, 911, 276, 223]. Grebenik [553]. Greek [770]. Greenwald [489]. Grenander [772]. Griffin [979]. Gross [479]. Group [261]. Grouped [480]. Groups [1021, 400, 702]. Growth [815, 704, 493, 900, 494]. Grundfragen [282]. Grundy [878]. Grunhut [701]. Guide [26, 760]. Guilford [105]. Gumbel [520, 970].

H [75, 678, 136, 103, 845, 923, 670, 523, 425, 318, 421, 467, 488, 1026, 813, 345, 945, 551, 230, 753, 566, 388, 620, 842, 1000, 102, 232, 593, 208, 107, 363, 943, 853, 451, 619, 946, 807, 309, 847, 449, 82, 337, 397, 702, 39, 567, 524, 676, 808, 29, 135, 179, 416, 948, 518, 110, 599, 834, 569]. H. [632]. Haberler [263]. Hackett [189]. Hagenbuch [981]. Hald [308, 308]. Halstrøm [107]. Handbook [185, 453, 487, 393, 1031]. Hans [556, 751]. Hansel [646]. Hansen [445]. Happens [236]. Harbury [1005]. Harold [364, 734, 746]. Harrison [87]. Harry [699, 447]. Hart [599]. Hartley [518]. Hartree [111, 314]. Harvey [572, 900]. Hawtrey [85]. Haycocks [619]. Hazard [886]. Hazelwood [953]. Head [22]. Heady [1010]. Health [649, 737, 836, 481, 843, 286, 680, 901]. Heasman [1010]. Heating [190]. Hegeland [551]. Heights [462]. Held [70, 69, 178, 177, 279, 915, 1023, 967, 39, 386, 483, 590, 697, 803]. Helen [415]. Henderson [451, 678, 957, 444]. Hendricks [773]. Henry [501, 956, 531]. Herbert [726, 599]. Herdan [725]. Herman [443, 485, 655]. Hermann [623]. Hickman [702, 984]. Hicks [316, 780]. High [319]. Hilda [869]. Hill [47, 595]. Hillmann [813]. Hirofumi [997]. Hirsch [837]. Historical [131, 115]. History [683, 452, 557, 665]. Hoel [517]. Hoffenberg [1033]. Hoffman [678]. Hogben [592, 74, 865]. Hogg [753]. Holland [6]. Hollis [997]. Home [87]. Hope [30]. Hospital [650, 903, 233]. Hospitals [679]. Household [691, 723, 289, 6, 874]. Housing [44, 681, 817, 1018]. Houthakker [847]. Howard [888]. Huff [498]. Hugh [467]. Hultgren [489]. Human [10, 82, 80]. Hundred [70, 69, 178, 177, 279, 386, 483, 590, 697, 803, 915, 1023]. Hundredths [519]. Hurstfield [401]. Hurwicz [997]. Hurwitz [445]. Hutchinson [756]. Hygiene [70, 69, 178, 177, 279, 386, 483, 590, 697, 803, 915, 1023, 967]. Hyman [599]. Hyperbolic [24, 672]. Hypergeometric [232]. Hypothesis [182].

I. [530]. Ian [31]. II [575, 341, 836, 51, 993, 694, 895, 592, 429, 416, 955]. III [875, 836, 429]. Ilersic [255]. Illness [1008, 1009]. Illustrated [928]. Immediate [86]. Immediately [967]. Immigrants [756]. Impact [425]. Implications [879]. Import [318]. Imports [52]. Inbreeding [27]. Income [982, 811, 564, 400, 848, 211, 1007, 691, 627, 723, 576, 512, 187, 939, 980, 114, 429, 237, 291, 494, 731, 850, 161, 349, 587, 331, 32, 848, 146, 653, 655, 429, 348]. Incomes [329, 556]. Incomplete [808]. Independent [549]. Index [273, 199, 318, 677, 745, 735, 781, 357, 663, 694, 160]. Index-Numbers [273]. Indexes [479]. India [212, 431, 12, 512, 229, 978]. Indian [654, 613]. Indices [661, 173, 920]. Individual [283]. Individuals [812, 702]. Induction [230]. Industrial [1021, 886, 730, 284, 515, 183, 836, 1005, 531, 160, 258, 12, 884, 998, 975, 1019, 1027, 360, 600, 261]. Industries [654]. Industry [678, 613, 719, 158, 461, 277, 1004, 67, 132, 262, 64, 5, 565, 628, 947, 957, 259, 777, 1030, 680]. Inexperienced [226, 129]. Infant [1010]. Inference [1043, 612, 597, 724, 415, 746, 835]. Inflation [963, 86, 930]. Influence [10, 155, 225]. Information [227, 4, 43, 65, 61, 171, 270, 384, 379, 408, 476, 506, 582, 689, 717, 794, 827, 910, 937, 1017, 44, 99, 545, 643, 66, 174, 158, 588, 335, 228, 26, 305, 277, 463, 383, 361, 414, 67, 362, 514, 25, 42, 132, 64, 587, 249, 5, 344, 565, 996]. Ingenieure [206]. Injuries [9]. Input [479]. Inquiries [350]. Inquiry [741]. Instalment [783]. Institute [413, 229, 51]. Instruments [111]. Insurance [113, 116, 228, 497, 264]. Insurances [497]. Intake [462]. Integer [341, 521]. Integrals [204, 257, 569, 951]. Intelligence [331]. Inter [52, 745, 51]. Inter-American [51]. Inter-Area [745]. Inter-War [52]. Interdependence [198]. Interest [583]. Interindustry [1033]. Intermediate [806, 135]. Internal [287]. International [529, 784, 929, 861, 52, 983, 469, 809, 9, 383, 155, 51, 285, 556, 413, 481, 229, 495, 497]. Interpolation [46, 519]. Interpretation [782, 845]. Interviewer [741]. Interviewers [226, 129]. Interviewing [599]. Introduction [591, 370, 73, 698, 1002, 889, 575, 1000, 394, 133, 531, 444, 894, 888, 1028, 395, 202, 944, 517, 778, 890, 918, 339, 45, 837, 201, 366, 732, 525, 1001]. Introductory [136]. Inventories [891, 145, 145]. Inventory [948]. Inversione [234]. Investigation [850, 896]. Investigations [358, 619]. Investment [924, 514, 601, 541, 86]. Investor [984]. Ireland [851]. Iron [64]. Irvine [491]. Irving [469]. Isaac [555]. Isles [851]. Issues [813]. IV [836, 51, 465, 653].

J [814, 137, 252, 32, 103, 644, 844, 866, 997, 877, 530, 284, 370, 81, 364, 757, 816, 523, 646, 1002, 185, 758, 316, 930, 401, 805, 116, 1033, 306, 453, 546, 1010, 532, 949, 11, 13, 552, 259, 253, 620, 776, 712, 970, 144, 232, 645, 890, 208, 288, 339, 389, 450, 28, 72, 531, 807, 283, 847, 449, 214, 201, 703, 366, 626, 548, 1024, 497, 1003, 105, 840, 868, 520, 521, 872, 205, 948, 422, 599, 780, 424, 629, 1027]. J. [723]. Jackson [417]. Jacob [30, 599]. Jaffe [288, 185]. Jamaica [873]. James [109, 287, 345, 957, 533, 1030]. Jan [144]. January [763]. Jasny [118]. Jeffery [287, 449]. Jefferys [533]. Jeffreys [109, 746]. Jenks [400]. Jennett [629]. Jensen [598, 107, 108]. Jerzy [280, 313]. Jevons [605]. John [944, 338, 318, 752, 786, 123, 924, 979, 733, 430, 999, 710]. Johnson [997, 29, 135, 895]. Jones [928, 906, 641, 80]. Jordan [775]. Joseph [415, 733]. Juiste [208]. Julius [555]. June [70, 69, 178, 177, 279, 386, 483, 590, 697, 803, 915, 1023]. Jureen [443]. Juvenile [701].

k-Statistics [569]. K.B.E [168]. Kafka [112]. Karlin [997, 948]. Karmel [946]. Katz [547]. Keeling [753]. Keith [32]. Kempthorne [548, 281, 894]. Kendall [779, 942, 465, 916, 73, 621, 100, 446]. Kenneth [997, 283]. Kent [1030]. Kew [795]. Kingdom [346, 530, 4, 384, 304, 441, 99, 199, 174, 329, 963, 662, 277, 22, 964, 661, 63, 541, 246]. Kingsley [212]. Klein [1043, 162]. Knight [346]. Knowledge [470]. Known [583]. Kolmogorov [549]. Kopal [647]. Kosambi [994]. Kravis [469]. Kuczynski [53]. Kuh [924]. Kuhn [179, 416, 702]. Kullback [996]. Kulldorff [651]. Kuznets [400, 429, 494].

L [109, 202, 103, 1037, 953, 670, 346, 752, 515, 1026, 971, 398, 546, 257, 778, 895, 104, 208, 260, 107, 188, 389, 549, 806, 671, 309, 906, 536, 899, 548, 839, 524, 521, 29, 135, 1029, 838]. Labor [628]. Laboratory [871]. Labour [4, 982, 966, 562, 262, 744, 478, 358, 1005]. Lambe [311]. Lancelot [592, 74, 865]. Land [384]. Langrangian [519]. Language [725]. Laning [807]. Large [521]. Lars [443]. Latent [257]. Latin [52, 117, 262]. Latin-American [262]. Latscha [465]. Lavers [213]. Law [442]. Lawrence [238, 752, 188, 162]. Learning [617]. Lectures [344, 313]. Lee [679]. Lees [753].

Lefeber [975]. Lehoczky [737]. Leibenstein [572, 900]. Lending [116, 703]. Length [412, 770]. Lengyel [497]. Leo [147]. Leon [547]. Leonard [778, 842]. Leonid [997]. Leser [258]. Leslie [623, 190, 123, 1014]. Lessons [210]. Lester [255]. Lev [415]. Level [542]. Levels [873, 901, 495, 871]. Levens [79]. Levinson [75]. Levitt [109]. Lewis [106, 675, 1002, 369, 453, 878, 10]. Lewis-Faning [878, 10]. L’Héritier [76]. Librarianship [344]. Library [16, 37, 59, 91, 125, 152, 170, 194, 219, 243, 268, 297, 325, 354, 376, 406, 436, 458, 474, 504, 539, 560, 580, 608, 635, 658, 686, 714, 740, 765, 792, 824, 825, 858, 883, 909, 935, 961, 989, 1016, 1041]. Lie [498]. Lieberman [872]. Life [754, 412, 116, 372, 107, 492, 878, 54]. Limit [549]. Limitation [10]. Limits [232]. Linder [206]. Lindley [424]. Lindquist [203, 418]. Linéaire [920]. Linear [997, 768, 667, 444, 950, 919, 974, 566, 391]. Lipsey [496]. List [16, 37, 59, 91, 125, 152, 170, 194, 219, 243, 268, 297, 325, 354, 376, 406, 436, 458, 474, 504, 539, 560, 580, 608, 635, 658, 686, 714, 740, 765, 792, 824, 825, 858, 883, 909, 935, 961, 989, 1016, 1041]. Little [147]. Living [873, 273, 431, 885, 114, 85, 495]. Ll [80]. Lloyd [189]. Local [382, 1021, 545, 275, 450, 225, 753]. Location [975]. Logan [876, 902, 977]. Logarithms [952, 448]. Logical [251]. Lognormal [805]. London [624, 344, 70, 69, 178, 177, 279, 386, 483, 590, 697, 803, 915, 1023, 967, 40, 1008, 33, 818]. Long [886, 939, 982]. Long-Term [886, 939]. Losses [238]. Lottery [585]. Louis [817, 975]. Loveday [921]. Luce [888]. Lundberg [348]. Lush [548]. Lyle [842, 604]. Lyon [763].

M [814, 502, 136, 597, 591, 679, 916, 997, 649, 238, 73, 816, 646, 928, 867, 891, 180, 621, 138, 627, 372, 1033, 648, 1010, 445, 957, 388, 104, 892, 71, 235, 902, 363, 415, 943, 806, 669, 852, 100, 654, 733, 430, 779, 1030, 1031, 516, 1029, 87, 942, 834, 446, 465]. Maat [208]. Maccoll [109]. Machines [111, 419]. Mackay [233]. Mackworth [82]. Maclean [11]. MacMillan [508]. Madeline [1030]. Madow [445]. Magazines [819]. Mainland [312]. Mainly [906]. Maintenance [694, 891]. Major [926]. Malcolm [757]. Male [743, 154]. Malthusian [1034]. Malzburg [679]. Man [80]. Management [491, 1000, 853, 626, 733, 998]. Managerial [1028]. Manford [702]. Mann [102]. Mannheim [623]. Manpower [288]. Manual [310, 395, 9]. Manufacturer [1036, 145]. Manufacturing [719, 629]. March [484, 39]. Margaret [109]. Marine [3]. Market [759, 760, 1032, 110, 87]. Marketing [410, 491]. Markets [703]. Marris [973]. Marschak [997]. Martin [453, 1032, 844]. Mary [287, 420]. Mason [898]. Massey [201]. Masuyama [597]. Materials [246, 401]. Mathematical [176, 542, 695, 866, 1044, 280, 517, 411, 769, 773, 836, 638, 157, 333, 868, 991, 948, 313, 341, 774]. Mathematics [1002, 464, 548, 567, 11, 839, 727]. Mathematische [838]. Matrices [257]. Matrix [391]. Matter [18, 60, 93, 1, 19, 38, 62, 94, 127, 153, 172, 221, 244, 269, 299, 196, 222, 245, 271, 327, 355, 378, 407, 300, 328, 356, 380, 409, 507, 659, 688, 716, 610, 637, 660, 690, 718, 860, 828, 938]. Matthews [452]. Maudsley [650, 903]. Maverick [106]. Max [701, 760]. Maxwell [893]. McCarthy [949, 890]. McCloskey [626, 733]. McCracken [841]. McKean [1006]. McKenzie [927]. McKinsey [339]. Meade [930]. Means [795, 342, 612]. Measles [720]. Measure [886]. Measurement [785, 966, 873, 530, 81, 1007, 766, 315, 495, 82, 885, 971, 901].

Measurements [925, 363]. Measuring [513, 511]. Medical [387, 615, 441, 893, 965, 47, 595, 312, 870]. Medicine [69, 177, 803, 1025, 441, 70, 178, 279, 386, 483, 590, 697, 915, 1023, 967, 84, 312]. Mediziner [206]. Meeting [96, 509, 639, 967, 70, 69, 178, 177, 279, 386, 483, 590, 697, 803, 915, 1023]. Megalithic [561]. Mehta [654]. Melman [730]. Members [492]. Men [147]. Mental [679, 1009, 1008]. Mercantile [3]. Merton [1043]. Meteorology [393]. Meterological [527]. Method [597, 306, 257, 188, 562, 155, 744, 202, 735]. Methoden [206]. Méthodes [76]. Methods [594, 137, 591, 724, 547, 623, 185, 398, 974, 972, 445, 896, 253, 776, 393, 466, 307, 804, 160, 1024, 1029, 665, 621, 138, 550, 726, 747, 922, 450, 946, 100, 811, 755, 369, 142, 671, 337]. Methorst [632]. Metodos [525]. Meyer [924, 726]. Michael [926]. Migrants [983]. Migration [449, 652, 955, 287, 929, 555, 679, 493, 651]. Miles [912]. Millar [729]. Miller [788, 655, 146, 521, 424]. Million [568]. Mills [426]. Milne [390]. Milton [653, 849, 469, 429]. Ministry [330]. Mises [869]. Mitchell [87, 236]. Mobility [513]. Modal [750]. Mode [480]. Model [598]. Modeles [707]. Models [897, 617]. Modern [837, 338, 760, 1003]. Modernes [422]. Modigliani [556]. Moe [108]. Mogey [816]. Monopoly [898]. Monte [726]. Moore [969]. Moral [616]. Morbidity [233, 927, 878, 977]. Morey [189]. Morgenstern [164]. Morse [891]. Mortalité [290]. Mortality [927, 830, 878, 737, 154, 1010, 619, 904, 905]. Mortara [290]. Mortgage [116, 703]. Mortimer [732]. Morton [703]. Moscow [695]. Moser [873, 928, 896]. Moses [815, 145, 1026]. Mosteller [617]. Mothes [422]. Motor [830, 438, 158]. Motor-Cars [830, 438]. Motoring [992]. Mounsey [370]. Move [454]. Movement [184, 652]. Movements [52, 964, 63]. Mr. [39]. Mrs. [712]. Multiple [834]. Multiplier [551]. Multivariate [916, 917, 918, 945]. Munroe [180]. Murray [772]. Musicians [412]. Mutually [795].

N [365, 851, 866, 953, 880, 210, 372, 895, 712, 917, 393, 549, 45, 449, 82, 626, 12, 1024, 1006, 998, 29, 135, 465]. Napolitani [622]. Natalità [234]. National [554, 649, 264, 469, 704, 819, 576, 512, 481, 783, 494, 850, 349, 587, 570, 811, 556, 731]. Nations [510, 1007, 48, 49, 50, 51, 186, 262]. Natural [952, 448]. Nature [227, 4, 43, 65, 384, 441, 44, 99, 545, 238, 643, 66, 174, 575, 158, 588, 335, 228, 26, 305, 277, 463, 383, 361, 414, 67, 362, 514, 25, 42, 132, 64, 587, 249, 5, 565]. Naturwissenschafter [206]. Naum [118]. Need [586]. Neighbourhood [816]. Neisser [556]. nella [234]. Net [479, 541]. Neter [752]. Neville [676]. Nevin [736]. Newcastle [877]. Newcastle-Upon-Tyne [877]. Newell [877]. Newspapers [819]. Newton [287]. Neyman [280, 72, 313]. Nielsen [671]. Nigeria [576]. Nine [519]. Nineteenth [386]. No [707, 1036, 160]. Nombres [920]. Nomenclature [50]. Nomography [79]. Non [997, 872]. Non-Central [872]. Non-Linear [997]. Nonparametric [804, 748]. Norbert [46]. Norges [207]. Normal [568, 439, 675, 423]. Norman [209]. Norris [1008]. Northern [851]. Note [89, 167, 216, 374, 404, 456, 472, 578, 762, 1039, 966, 695, 1044, 832, 994, 130, 331, 665]. Notes [56, 35, 14, 34, 55, 88, 122, 192, 121, 150, 166, 191, 241, 266, 295, 215, 240, 265, 294, 321, 352, 320, 351, 373, 403, 433, 500, 432, 455, 471, 499, 535, 603, 534, 558, 577, 602, 631, 709, 630, 656, 684, 708, 790, 822, 738, 761, 789, 821, 855, 933, 854, 881, 907, 932, 959, 1013, 958, 985, 1012, 1038, 796, 723, 336]. Notices [56]. November [763]. Number [769, 357, 694]. Numbers [273, 199, 112, 745, 160, 799, 663, 694, 567, 585, 840, 781]. Numerical [257, 671, 550, 314, 922, 390, 647, 727]. Nurkse [402].

O [678, 281, 515, 452, 451, 731, 518]. O.E.E.C [574]. Observational [612]. Observations [439, 892, 164]. Observatory [795]. Obstetrical [737]. Obtained [226]. Occupational [905]. Oceanography [797]. Offenders [701]. Offisielle [207]. Ogborn [372]. Ogilvie [877]. Oils [43]. Ökonometrie [282]. Old [799, 694]. Older [956]. Oligopoly [863, 1032]. One [70, 69, 178, 177, 279, 386, 483, 590, 697, 803, 915, 1023, 396]. Operations [954, 778, 1019, 842, 345, 626, 733, 998, 1029]. Operative [26]. Operatives [360]. Opionions [949]. Order [341]. Ordered [439]. Orders [521, 396]. Ordinary [967]. Organization [843]. Orthogonal [204]. Orwin [633]. Oscar [548, 894]. Oskar [164]. Ostle [618]. Other [120, 293, 941, 619, 163, 367]. Otis [758]. Outline [522]. Outlines [461]. Output [479, 22]. Over-Populated [744]. Overseas [451, 601]. Owen [548]. Ownership [812].

P [141, 650, 844, 876, 903, 287, 350, 978, 969, 211, 81, 189, 421, 517, 625, 891, 30, 753, 393, 902, 655, 931, 977, 946, 214, 105, 756, 521, 728, 729, 424, 629]. P. [625]. Pacifying [225]. Paden [203]. Pairing [439]. Paish [163]. Pakistan [431, 512, 212]. Pandurang [486]. Paper [1044, 723, 994]. Papers [798, 596, 1036, 10]. Paradine [421]. Parental [331]. Parry [1002, 453, 81]. Part [875, 341, 895, 302, 465]. Partially [808]. Particular [510, 277]. Parts [22]. Pass [795]. Passenger [40]. Past [1007, 10].

Patients [769, 737]. Pattern [318, 553, 812, 512]. Patterns [54]. Paul [846, 1007]. Pay [95]. Payments [3, 85, 184, 497]. Payne [260]. Peacock [264]. Pearce [367]. Pearson [518, 675]. Peasant [744]. Peck [953]. People [655]. Perception [729]. Perennial [367]. Performance [82, 118]. Period [52, 743, 86]. Periodical [17, 92, 126, 195, 220, 298, 326, 377, 459, 475, 505, 540, 581, 609, 636, 687, 715, 793, 826, 859, 936, 962, 1042]. Periodicals [15, 36, 58, 90, 124, 151, 169, 193, 218, 242, 267, 296, 324, 353, 375, 405, 435, 457, 473, 503, 538, 559, 579, 607, 634, 657, 685, 713, 739, 764, 791, 823, 857, 882, 908, 934, 960, 988, 1015, 1040]. Periodicity [720]. Perks [619]. Perlen [205]. Perry [1030, 259, 104]. Personnel [81, 814]. Peter [571, 182, 624]. Petersen [652]. Petrographic [750]. Petrol [992]. Petroleum [788, 277]. Ph [76]. Pharmacy [839]. Phenomenon [478]. Philip [945, 949, 842, 890, 604]. Philosopher [616]. Physics [836]. Piano [622]. Pigou [987]. Place [137, 734]. Places [952, 448]. Planned [652]. Planning [611, 186, 788, 968, 149]. Plans [118]. Plants [367]. Pleasure [1037]. Polak [144]. Polar [676]. Police [979]. Policy [941, 931]. Politics [523]. Polynomials [204, 396]. Pompilj [622, 78]. Populated [744]. Population [212, 510, 861, 175, 427, 625, 755, 956, 30, 588, 10, 439, 97, 845, 978, 467, 142, 285]. Populations [879]. Ports [214]. Positive [648, 341]. Post [555, 86, 163]. Post-War [555, 86, 163]. Postulats [648]. Postwar [627]. Pound [303]. Poverty [213]. Power [469, 303, 832]. Practical [520]. Practice [372, 620, 131, 972, 977, 149]. Prais [847]. Prediction [966, 623]. Preliminary [345]. Prescribing [844]. Present [1007]. Presentation [638, 487, 820]. Presentations [800]. Presented [69, 177].

President [39]. Prest [530, 576, 850]. Price [286, 303, 155, 885]. Prices [68, 159, 250, 334, 440, 544, 642, 771, 887, 995, 304, 52, 302, 863, 661, 173, 677, 735]. Primer [669, 365]. Principal [117]. Principle [108]. Principles [464, 614, 312, 254, 969, 620, 47, 595, 110]. Private [529, 992, 601]. Probabilità [78]. Probabilités [77, 181]. Probabilities [651]. Probability [442, 889, 593, 809, 836, 133, 45, 72, 865, 520, 674, 139, 423, 280, 180, 230, 251, 596, 775, 313, 869, 835, 28, 806, 1043, 1028]. Probit [252]. Problem [978, 332, 965, 163, 898, 350]. Problems [510, 319, 836, 885, 368, 1029, 784, 117, 402]. Procedure [949]. Proceedings [70, 7, 96, 178, 279, 386, 483, 509, 590, 639, 697, 803, 915, 1023, 967, 836, 51, 727, 280, 527, 809, 836, 51]. Processes [591, 807, 389]. Processing [767, 777]. Product [554, 199, 63]. Production [842, 853, 41, 964, 160, 884, 975, 427, 948]. Productive [766]. Productivity [628, 964, 262, 358, 730, 426]. Products [460, 469]. Professor [763]. Program [627]. Programme [886, 920]. Programming [974, 768, 997, 841, 667, 444, 919, 923]. Progress [236, 148, 426]. Projection [755]. Proneness [247]. Propensity [831]. Proportionality [694]. Proportions [743]. Proposals [677]. Proudfoot [757]. Psychological [484]. Psychologist [395]. Psychology [523, 893, 418, 105, 183]. Psychometry [836]. Psychoses [926]. Public [815, 130, 146]. Publications [120, 293, 825]. Publishing [174]. Punched [259, 1030]. Purchasing [469, 303, 832]. Purification [257]. Purposes [50].

Quality

[165, 984, 776, 1000, 284, 138, 524, 1027]. Quantitative [556, 312]. Quarrels [225]. Quenouille [136, 388, 363, 943, 834]. Questions [260]. Queueing [953]. Queues [891]. Quota [248, 381].

R [442, 141, 968, 27, 287, 571, 953, 530, 666, 255, 318, 316, 181, 706, 812, 53, 116, 683, 576, 399, 452, 111, 431, 259, 314, 924, 712, 922, 213, 920, 616, 853, 45, 807, 468, 449, 162, 292, 774, 12, 85, 850, 839, 779, 391, 870, 999, 617, 888, 110, 780, 465]. Radhakrishna [307]. Ragazzini [999]. Ragnar [402]. Raiffa [888]. Rail [42]. Ramamohana [994]. Ramo [1031]. Random [568, 549, 675, 807]. Randomization [994, 864]. Rank [621, 100]. Rao [994, 307]. Rashi [1009]. Rates [273, 129, 832, 833, 721]. Raw [401, 246]. Raymond [575, 395]. Readership [819]. Readings [759, 919]. Real [199, 63]. Reappraisal [641]. Reasoning [890, 943]. Received [825]. Recente [234]. Recession [528]. Recommendations [861]. Recording [342]. Recovery [574]. Rectangular [676]. Rectangular-Polar [676]. Redistribution [627]. Reference [392, 591, 510, 873, 805, 277, 885, 86, 145]. Refugees [757]. Regional [329, 414, 848]. Registrar [875, 754, 8, 904, 905]. Registrar-General [875, 754]. Regression [842]. Regulations [387, 1021]. Reichmann [840]. Reiss [758]. Rejoinder [801]. Related [587]. Relating [384, 927, 277, 64]. Relation [623, 769, 331]. Relations [479, 566, 531]. Relationship [865]. Relative [799, 52]. Rene [707]. Repairing [276]. Replacement [719]. Report [69, 177, 278, 385, 482, 589, 696, 802, 914, 1022, 650, 903, 553, 574, 342, 10, 261, 940, 349, 236, 677, 1011, 874]. Reports [97, 820, 104]. Requirements [460]. Research [594, 954, 848, 491, 345, 737, 836, 862, 187, 1036, 760, 186, 783, 626, 733, 237, 291, 317, 161, 998, 1029, 344, 704, 759, 778, 307, 870, 231, 110, 599, 618, 547]. Researches [82]. Reserve [783]. Resnikoff [872]. Resources [788, 288, 1006, 625].

Response [129]. Results [226]. Retail [304, 677, 735, 533]. Retention [211]. Returns [17, 92, 126, 195, 220, 298, 326, 377, 459, 475, 505, 540, 581, 609, 636, 687, 715, 793, 826, 859, 936, 962, 1042]. Review [75, 109, 759, 554, 653, 678, 529, 785, 787, 814, 982, 136, 442, 594, 212, 137, 141, 165, 202, 252, 256, 365, 392, 522, 597, 968, 27, 32, 103, 591, 644, 724, 811, 650, 679, 810, 844, 851, 876, 903, 927, 1008, 1025, 417, 772, 866, 916, 944, 997, 113, 233, 286, 287, 290, 350, 571, 572, 649, 782, 845, 873, 877, 923, 978, 1037, 953, 954, 400, 784, 848, 203, 338, 420, 547, 670, 730, 929, 346, 530, 981, 666, 1009, 487, 921, 969, 239, 573, 255, 280, 284, 310, 370, 623, 211, 238, 52]. Review [73, 81, 364, 757, 816, 846, 875, 880, 983, 184, 189, 523, 646, 849, 425, 427, 628, 682, 815, 1007, 264, 754, 491, 318, 443, 698, 752, 928, 1002, 897, 984, 185, 234, 281, 319, 421, 428, 464, 467, 488, 515, 517, 553, 625, 755, 758, 867, 879, 889, 891, 900, 947, 956, 976, 1026, 1034, 893, 971, 398, 190, 210, 147, 316, 528, 101, 180, 181, 621, 138, 148, 469, 575, 627, 704, 706, 812, 813, 819, 898, 925, 930, 401, 492, 974, 805, 972, 53, 345, 418, 945, 116, 372, 683, 76, 551, 1033, 598, 576, 30, 550, 306, 369]. Review [399, 453, 452, 470, 557, 574, 489, 527, 111, 230, 546, 648, 725, 869, 753, 1010, 315, 566, 707, 343, 431, 532, 445, 896, 949, 112, 786, 11, 13, 31, 79, 83, 84, 957, 552, 568, 726, 204, 257, 259, 314, 341, 342, 924, 680, 878, 253, 388, 620, 747, 737, 776, 778, 842, 895, 1000, 788, 102, 104, 183, 498, 78, 970, 922, 144, 232, 394, 592, 593, 645, 700, 773, 809, 836, 890, 892, 917, 918, 213, 308, 8, 9, 71, 119, 208, 235, 260, 107, 108, 133, 393, 484, 485, 749, 188, 734, 142, 288, 920, 187]. Review [282, 902, 736, 363, 466, 134, 251, 307, 339, 389, 415, 549, 596, 616, 699, 746, 804, 835, 943, 841, 853, 979, 450, 526, 1004, 655, 677, 10, 28, 33, 45, 46, 72, 77, 451, 996, 1005, 931, 619, 701, 705, 806, 843, 901, 904, 905, 977, 1036, 490, 760, 236, 667, 669, 735, 837, 852, 946, 973, 531, 820, 600, 671, 807, 100, 182, 283, 444, 106, 347, 1043, 48, 49, 50, 51, 668, 140, 309, 311, 289, 468, 847, 117, 143, 160, 186, 261, 262, 285, 449, 493, 495, 555, 556, 681, 624, 47, 82, 214, 595, 337, 54]. Review [74, 397, 980, 906, 162, 781, 201, 703, 263, 654, 340, 366, 533, 114, 601, 149, 402, 207, 429, 783, 626, 733, 209, 292, 146, 258, 426, 164, 237, 291, 317, 494, 702, 774, 899, 817, 12, 80, 85, 86, 548, 1024, 497, 163, 430, 496, 731, 850, 161, 348, 349, 1003, 105, 567, 839, 840, 865, 868, 894, 1035, 926, 732, 1006, 312, 524, 367, 115, 651, 652, 756, 955, 1011, 118, 779, 368, 390, 391, 419, 519, 520, 521, 647, 672, 673, 674, 675, 676, 727, 728, 750, 777, 808, 870, 871, 872, 950, 951, 952, 998, 999]. Review [1030, 1031, 874, 29, 135, 179, 205, 206, 231, 416, 516, 525, 617, 751, 775, 888, 948, 975, 1028, 1029, 1032, 422, 87, 518, 942, 818, 748, 110, 344, 395, 454, 599, 729, 780, 1001, 834, 838, 919, 139, 254, 313, 336, 371, 396, 423, 424, 446, 447, 448, 465, 486, 569, 570, 618, 622, 612, 145, 629, 1027]. Revision [780]. Revival [528]. Richard [653, 810, 649, 530, 869, 315, 777, 641]. Richardson [257, 531]. Richter [751]. Riordan [944]. Rios [525]. Risk [508, 646]. Risque [648]. Rivett [421]. Road [912, 461, 42, 692, 223]. Robert [997, 921, 364, 189, 496, 391, 1030, 617, 1028]. Roberts [753, 699]. Robin [973]. Rodney [490]. Roland [1006]. Role [315]. Roll [557]. Rome [413]. Romig [447]. Ronald [594, 724, 1035, 344]. Roose [528]. Rosander [254]. Rose [346]. Rosenblatt [772]. Rostas [536]. Rowe [530]. Rowntree [537, 213]. Rowson [57]. Roy [707, 917, 920, 526, 856]. Royal [1023, 650, 903, 10, 70, 69, 97, 178, 177, 279, 386, 483, 590, 697, 803, 915, 1020, 967, 39]. Rudolf [251]. Run [912]. Rural [304, 758, 1005, 786]. Russell [778, 434]. Russia [665].

S [591, 851, 1037, 425, 427, 682, 898, 30, 753, 79, 259, 917, 260, 804, 667, 669, 847, 899, 12, 497, 367, 675, 1030, 231, 948, 518, 919, 569]. Sainsbury [624]. Salaries [346]. Same [967, 863]. Sample [332, 477, 972, 693, 445]. Sampled [999]. Sampled-Data [999]. Samples [332, 439]. Sampling [886, 752, 21, 742, 450, 248, 381, 140, 330, 131, 665, 101, 773, 972, 949, 486]. Samuel [997, 1001]. Sanders [620]. Sarkar [880]. Sasieni [1029]. Sastry [12]. Saul [974]. Saulnier [116]. Saunders [928, 839]. Savage [546]. Save [831]. Saville [786]. Saving [691, 575, 723]. Savings [400]. Scarf [948]. Schemes [742]. Schlaifer [1028]. Schlesinger [318]. Schlote [451]. Schmid [487, 669]. School [70, 69, 178, 177, 279, 386, 483, 590, 697, 803, 915, 1023, 967, 344]. Schools [200, 969, 83, 336]. Schrock [138]. Science [75, 259, 366, 1030, 442, 464]. Sciences [1025, 547, 893, 748]. Scientific [814, 724, 835, 442, 188, 746]. Scientist [395]. Scotch [565]. Scotland [683, 543, 258]. Scottish [492, 23]. SEAC [570]. Secants [519]. Second [697, 280]. Secretariat [50]. Section [615, 1021]. Sections [7]. Sector [130]. Seebohm [537, 213]. Seers [114]. Selected [983, 494, 596, 955]. Selection [81]. Selmer [997]. Seltzer [238]. Seminaire [707]. Senders [971]. Sentence [770]. Sentence-Length [770]. Sentralbyrå [207]. September [342, 51]. Sequential [256]. Sergei [211]. Serial [825]. Series [653, 318, 1033, 302, 46, 182, 429, 797, 348, 772, 485, 834]. Service [649]. Services [586, 441, 545, 863]. ses [77]. Session [413, 229]. Sessional [69, 177]. Seventeenth [178, 177]. Sewell [782]. Sex [755]. Sexagesimal [519]. Seymour [730]. Shackle [1037, 899]. Shares [400]. Sheldon [956]. Shelter [817]. Shepherd [926]. Ship [721, 276]. Ship-Repairing [276]. Shipbuilding [276, 906]. Shipping [214].

Ships [863]. Shirras [606]. Short [521, 573]. Short-Term [573]. Shortage [508]. Shubik [1032]. Shuffling [994, 864]. Shunter [247]. Sickness [286, 902]. Sidney [748]. Siegel [30, 748]. Significance [728, 465, 31]. Significant [519]. Silverman [490]. Simon [400, 429, 494, 1031, 57]. Simple [567]. Simulation [1019]. Simultaneous [950]. Since [318, 964, 114, 734, 496, 955]. Sine [569]. Sines [672]. Singapore [119]. Sir [763, 501, 168]. Sites [561]. Sittig [208]. Situation [681]. Six [8]. Sixteen [952, 448]. Sixteenth [70, 69]. Sixth [574]. Sixto [525]. Size [720, 941, 939]. Sizes [925]. Slower [795]. Small [214]. Smart [892]. Smith [32, 649, 464, 1005]. Smoking [1036, 1035]. Smoothing [46]. Snedecor [747]. Social [861, 800, 928, 879, 316, 21, 896, 31, 84, 119, 235, 736, 669, 513, 652, 395, 599, 844, 113, 981, 758, 1034, 1010, 235, 450, 283]. Socialized [118]. Societies [425, 319]. Society [70, 69, 178, 177, 279, 386, 483, 590, 697, 803, 915, 1020, 1023, 967, 39, 342]. Solomon [996, 496]. Solow [997]. Solution [390, 950]. Some [613, 318, 257, 156, 583, 593, 638, 885, 797, 799, 663, 694, 1035, 520, 101, 917, 258]. Sources [227, 4, 43, 65, 384, 811, 441, 44, 99, 545, 643, 66, 174, 158, 588, 335, 228, 26, 305, 277, 463, 383, 361, 414, 67, 362, 514, 25, 42, 132, 64, 587, 249, 5, 565, 453]. Southall [502]. Soviet [628, 695]. Space [975, 817]. Spangler [788]. Spear [420]. Spearman [80]. Special [227, 4, 43, 65, 384, 967, 392, 591, 441, 873, 44, 99, 545, 643, 66, 174, 805, 158, 588, 335, 228, 26, 305, 277, 463, 383, 361, 414, 67, 362, 514, 25, 42, 132, 885, 86, 64, 587, 249, 5, 565, 145]. Specification [796, 629]. Spencer [835]. Spiegelman [732]. Spinning [358]. Sprowls [666]. Squares [13]. Staff [492]. Stafleu [208]. Standard [85, 431]. Standardization [103]. Standards [104, 285, 495, 570]. Standing [1036].

Stanley [605]. State [489, 213]. States [318, 575, 30, 162, 601, 496, 115, 438, 787, 956, 600, 115]. Stationary [772, 485, 46]. Statistical [227, 4, 70, 69, 15, 36, 58, 90, 43, 65, 14, 34, 55, 88, 178, 177, 124, 151, 169, 193, 121, 150, 166, 191, 279, 218, 242, 267, 296, 215, 240, 265, 294, 386, 324, 353, 375, 405, 384, 320, 351, 373, 403, 413, 483, 435, 457, 473, 503, 432, 455, 471, 499, 590, 586, 538, 559, 579, 607, 534, 558, 577, 602, 697, 634, 657, 685, 713, 630, 656, 684, 708, 803, 739, 764, 791, 823, 738, 761, 789, 821, 915, 857, 882, 908, 934]. Statistical [854, 881, 907, 932, 1020, 1023, 967, 960, 988, 1015, 1040, 958, 985, 1012, 1038, 137, 165, 202, 365, 597, 644, 650, 903, 1044, 44, 99, 545, 969, 370, 875, 185, 563, 643, 66, 138, 174, 492, 158, 369, 588, 640, 335, 228, 26, 645, 890, 918, 305, 308, 8, 9, 393, 277, 463, 383, 307, 943, 361, 414, 67, 362, 514, 946, 25, 42, 132, 1043, 48, 50, 51, 229, 337, 201, 366, 12, 39, 64, 561, 115, 991, 358, 779, 20, 871, 587, 516, 249, 5, 565, 424, 594, 724, 772, 364, 752, 398, 306, 253]. Statistical [747, 776, 895, 1000, 308, 466, 134, 415, 1024, 865, 368, 520, 870, 231]. Statistician [611]. Statisticians [940, 518]. Statistics [227, 4, 43, 65, 2, 176, 200, 301, 387, 384, 542, 392, 811, 876, 927, 1025, 410, 441, 44, 99, 545, 280, 829, 983, 928, 467, 643, 66, 174, 158, 453, 869, 588, 128, 335, 228, 737, 26, 583, 836, 305, 862, 277, 463, 383, 584, 596, 361, 414, 67, 362, 705, 843, 977, 1036, 514, 946, 25, 42, 132, 143, 449, 638, 224, 272, 157, 333, 39, 64, 105, 839, 312, 115, 587, 1028, 249, 942, 748, 5, 565, 313, 336, 481, 136, 810, 233, 338, 420, 666, 921, 284, 310, 73, 1007, 698, 488, 517, 976]. Statistics [971, 546, 498, 394, 700, 804, 526, 45, 72, 996, 669, 837, 852, 106, 309, 311, 47, 595, 1003, 894, 254, 446, 569, 1027, 554, 203, 670, 255, 421, 399, 112, 970, 699, 979, 548, 29, 135, 618]. Statistik [207, 838]. Statistiques [76].

Statistische [206]. Statistisk [207]. Statistiske [207]. Statutory [410]. Staveley [344]. Steel [64, 884]. Stember [599]. Stephan [949]. Stephen [813]. Sterling [303, 292, 430]. Stevens [1014]. Stevenson [84]. Stewart [633, 576, 288]. Stigler [814]. Stillbirths [31]. Stochastic [591, 389, 617]. Stoll [391]. Stone [530, 653, 315, 712]. Strategy [1032]. Strategyst [552]. Strength [574]. Structure [613, 1004, 258, 1032, 654]. Stuart [942]. Studenski [1007]. Students [1002, 464, 147]. Studies [187, 160, 237, 955, 1019, 997, 291, 161, 948]. Study [387, 442, 679, 1007, 443, 575, 627, 477, 452, 144, 485, 531, 155, 381, 468, 449, 493, 12, 926, 924, 624]. Sugar [67]. Suicide [624]. Suivant [290]. Sukhatme [486]. Summary [927]. Sums [549]. Supplement [954, 754, 904, 905, 333, 115, 165]. Supply [814, 460, 788, 940]. Surplus [744]. Survey [851, 52, 928, 819, 972, 53, 913, 993, 21, 445, 949, 737, 119, 902, 33, 117, 186, 131, 818, 693, 888, 896]. Surveys [332, 477, 235, 486]. Survival [990]. Sutherland [31]. Sweden [972]. Sydney [845]. Symmetric [465]. Symposia [727]. Symposium [280, 836, 484, 419, 729, 1004, 726, 140]. System [40, 783]. Systems [1006, 999].

T [650, 866, 623, 264, 190, 147, 753, 918, 931, 806, 668, 548, 1024, 434, 675]. Table [24, 521, 728, 448, 465, 569, 521, 728, 465, 672, 673, 952, 371]. Tables [575, 341, 308, 368, 519, 520, 870, 518, 396, 423, 569, 465, 953, 676, 424, 447, 256, 754, 519, 674, 872, 139, 396, 423, 448, 569]. Tax [238]. Taxable [32]. Taxes [146]. Taylor [323]. Tea [239]. Teach [810]. Teaching [200, 542, 336]. Technique [629]. Techniques [969, 868, 422]. Technological [309]. Technologists [421]. Technology [788]. Telecommunications [335]. Telephone [809]. Television [662].

Temperature [795]. Ten [799, 952]. Ten-Year-Old [799]. Tendenza [234]. Term [886, 939, 573]. Terms [779]. Territories [601]. Testing [182]. Tests [728, 104, 256, 465]. Tetley [29, 135]. Text [806, 135]. Text-Book [806, 135]. Textile [262]. Theil [566]. Their [31, 259, 820, 756, 1030, 255, 318, 467]. Theodore [323]. Theoretical [446]. Theorie [648]. Theories [845, 1034, 972, 131]. Theory [442, 27, 866, 572, 73, 849, 1007, 889, 101, 411, 593, 773, 809, 836, 308, 133, 339, 45, 996, 667, 668, 638, 663, 694, 865, 520, 179, 416, 775, 948, 1032, 942, 486, 417, 1026, 551, 445, 391, 942, 780, 180, 616, 516]. Third [803, 695, 836]. Thomas [997, 812, 929, 493]. Thomson [1044]. Thos [489]. Thought [419, 557]. Throop [32]. Thursday [279, 967]. Tibbitts [956]. Timber [362]. Time [1033, 302, 46, 182, 797, 834, 772, 485, 899]. Time-Series [772, 485]. Time-Series-Part [302]. Tinbergen [144]. Tins [272]. Tintner [340]. Tippett [670, 842, 309]. Tisdale [30]. Titmuss [649]. Tobacco [65, 1036]. Todd [1036]. Tool [616]. Tore [972]. Toronto [527]. Total [912]. Toxicity [104]. Trade [318, 452, 489, 451, 745, 155, 556]. Trading [533]. Training [623]. Transactions [497]. Transfers [146]. Transformation [24]. Transport [461, 768, 383, 40, 143, 223, 975, 489]. Tranter [868]. Travel [33, 818]. Tre [207]. Treatise [230]. Treatment [238, 788, 359]. Trees [367]. Trefethen [626]. Trend [553, 939, 496]. Trends [876, 154]. Triennial [650, 903]. Tropical [70, 69, 178, 177, 279, 386, 483, 590, 697, 803, 915, 1023, 967, 850]. Trumpler [364]. Truth [869]. Tse [184]. Tuberculosis [876]. Tucker [179, 416]. Turbulence [796, 797]. Turnover [966, 562, 478]. Tuttle [852]. Twentieth [483]. Twenty [590, 697, 803, 915, 1023]. Twenty-Fifth [1023]. Twenty-First [590]. Twenty-Fourth [915]. Twenty-Second [697]. Twenty-Third [803]. Twins [799]. Two [906, 808, 871]. Tyne [877]. Typenlehre [428]. Types [743].

U [772]. U. [994]. U.K. [65, 3, 64]. U.S. [185]. U.S.A. [20]. U.S.S.R [118]. Udny [73, 217]. Uncertainty [788, 1028]. Under-Developed [52, 425, 744]. Underdeveloped [931, 402]. Understanding [638]. Unemployment [511, 785]. Unified [583]. Unit [745]. United [787, 346, 530, 318, 956, 575, 30, 600, 50, 51, 162, 601, 496, 115, 4, 384, 304, 441, 510, 99, 199, 174, 329, 963, 662, 438, 277, 22, 964, 661, 63, 541, 246, 48, 49, 186, 262]. Universities [911, 969, 704]. University [542, 492, 344]. Upon [247, 877]. Upper [400]. Uppsala [484]. Urban [304, 175, 758, 289, 1018, 116, 703]. Use [2, 384, 767, 638, 817, 839]. Used [330]. Uses [805]. Using [965]. Utilization [190, 288]. Utting [723, 450]. Uzawa [997].

V [923, 239, 553, 947, 836, 994, 549, 699, 397, 258, 419, 424, 486]. Vacation [344]. Vajda [667, 919]. Value [745, 520, 423]. Values [357, 520, 283, 204]. Vance [752, 188]. Variabilita [749]. Variability [741]. Variables [549]. Variance [664, 102]. Variations [795, 329]. Vehicles [912]. Vera [1008, 815]. Vernon [81]. VI [104]. Vice [39]. Vice-President [39]. Vie [290]. Villars [231]. Virginia [971]. Visit [695]. Visual [112, 638]. Visualised [731]. Vital [976, 588, 481, 705, 843]. Vitamins [392]. Vol [530, 341, 592, 836, 10, 51, 237, 291, 161, 1031, 518, 942]. Volume [61, 171, 270, 379, 408, 476, 506, 582, 689, 717, 794, 827, 910, 937, 1017, 575, 745, 416].

W [678, 787, 202, 522, 876, 203, 530, 632, 81, 427, 682, 984, 867, 889, 398, 101, 575, 753, 445, 259, 253, 747, 1000, 104, 892, 918, 282, 902, 699, 451, 619, 977, 444, 548, 163, 840, 779, 390, 808, 1030, 179, 416, 599, 729, 465, 629]. W. [1044, 434]. W.C [69, 177]. Waerden [838]. Wage [273, 963]. Wage-Rates [273]. Wages [787, 543, 360, 23, 346]. Wahrscheinlichkeitstheorie [751]. Wald [134, 596]. Wales [876, 927, 875, 754, 928, 786, 8, 904, 905, 1011, 649]. Walker [415]. Wallis [699]. Walter [981, 628, 773, 668]. War [52, 210, 555, 86, 163, 955]. Wars [906]. Wartime [532]. Washington [570]. Wastage [721]. Water [1006]. Watson [753]. Wattal [978]. Waugh [368]. Wealth [653, 187, 429, 237, 291, 161, 348, 848, 187, 291, 161]. Weatherford [787]. Weaver [364]. Webb [322]. Wednesday [70, 178, 386, 483, 590, 697, 803, 915, 1023]. Weeks [81]. Weighing [28]. Weights [462]. Welfare [213]. Well [583]. Well-Known [583]. Welsh [736, 453]. Werner [451, 837]. Wesley [236]. West [778, 410]. Westerborg [318]. Western [261]. Whelpton [30]. Whisky [565]. Whittle [182]. Wholesale [68, 159, 250, 334, 440, 544, 642, 771, 887, 995, 661, 173]. Wickizer [239]. Wiener [46]. Wilfred [201, 1014]. Wilhelm [428]. Wilkins [623, 190]. William [489, 71, 133, 652, 599]. Williams [552, 1003]. Willis [417]. Winkler [428, 282]. Winnick [817]. Winslow [286]. Wisconsin [812]. Wishart [710, 620]. Without [112]. Wold [443, 485]. Wolfenden [467]. Wolstenholme [729]. Women [925]. Woodlands [343]. Wooldridge [1031]. Worker [344, 183]. Workers [594]. Working [289]. Working-Class [289]. Works [107]. World [510, 460, 843, 51, 906, 955, 427, 682, 625]. Woytinsky [427, 682, 427, 682]. Wright [230]. Wynn [80].

X [341]. XI [161].

Yaspan [1029]. Yastremsky [833]. Yates [870]. Year [69, 177, 875, 878, 48, 49, 143, 799, 48]. Years [813, 83, 8, 130, 10]. York [679]. Youden [253]. Yourself [810]. Yule [73, 217].

Z [837]. Zahlentheorie [205]. Zdenek [647]. Zealand [532, 693]. Zero [396, 448]. Zygmunt [737].

References

Anonymous:1950:FMa

[1] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 113(1):??, ???? 1950. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2980796.

Anonymous:1950:USB

[2] Anonymous. The use of statistics in business. Journal of the Royal Statistical Society. Series A (General), 113 (1):1–8, ???? 1950. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2980797.

Kendall:1950:UKM

[3] M. G. Kendall. The U.K. mercantile marine and its contribution to the balance of payments. Journal of the Royal Statistical Society. Series A (General), 113(1):9–36, ???? 1950. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2980798.

Ainsworth:1950:SNS

[4] R. B. Ainsworth. The sources and nature of statistical information in special fields of statistics United Kingdom labour statistics. Journal of the Royal Statistical Society. Series A (General), 113(1):37–49, ???? 1950. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2980799.

Wild:1950:SNS

[5] A. C. Wild. The sources and nature of statistical information in special fields of statistics: Statistics of the cotton industry. Journal of the Royal Statistical Society. Series A (General), 113(1):50– 58, ???? 1950. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2980800.

Stuvel:1950:HEF

[6] G. Stuvel and S. F. James. Household expenditure on food in Holland. Journal of the Royal Statistical Society. Series A (General), 113(1):59–80, ???? 1950. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2980801.

Anonymous:1950:PS

[7] Anonymous. Proceedings of sections. Journal of the Royal Statistical Society. Series A (General), 113(1):81–86, ???? 1950. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2980802.

K:1950:BRBa

[8] J. K. Book review: The Registrar General’s Statistical Review of England and Wales for the Six Years 1940–1945. Journal of the Royal Statistical Society. Series A (General), 113(1):87–90, ???? 1950. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2980803.

K:1950:BRBb

[9] J. K. Book review: Manual of the International Statistical Classification of Diseases, Injuries and Causes of Death. Journal of the Royal Statistical Society. Series A (General), 113(1):90–92, ???? 1950. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2980804.

M:1950:BRBa

[10] W. J. M. Book review: Papers of the Royal Commission on Population, Vol.

1. Report of an Enquiry Into Family Limitation and Its Influence on Human Fertility During the Past Fifty Years, by E. Lewis-Faning. Journal of the Royal Statistical Society. Series A (General), 113(1):92–93, ???? 1950. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2980805.

H:1950:BRBa

[11] H. O. H. Book review: Descriptive Mathematics, by J. Maclean. Journal of the Royal Statistical Society. Series A (General), 113(1):93–94, ???? 1950. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2980806.

S:1950:BRBa

[12] G. F. S. Book review: A Statistical Study of India’s Industrial Development, by N. S. R. Sastry. Journal of the Royal Statistical Society. Series A (General), 113(1):94–95, ???? 1950. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2980807.

H:1950:BRBb

[13] H. O. H. Book review: Bourne’s Book of Squares, by J. B. Bourne. Journal of the Royal Statistical Society. Series A (General), 113(1):95, ???? 1950. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2980808.

Anonymous:1950:SNa

[14] Anonymous. Statistical notes. Journal of the Royal Statistical Society. Series A (General), 113(1):96–98, ???? 1950. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2980809.

Anonymous:1950:SEAa

[15] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society. Series A (General), 113(1):99–102, ???? 1950. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2980810.

Anonymous:1950:LALa

[16] Anonymous. List of additions to the library. Journal of the Royal Statistical Society. Series A (General), 113 (1):103–109, ???? 1950. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2980811.

Anonymous:1950:PRa

[17] Anonymous. Periodical returns. Journal of the Royal Statistical Society. Series A (General), 113(1):110–112, ???? 1950. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2980812.

Anonymous:1950:BMa

[18] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 113(1):??, ???? 1950. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2980813.

Anonymous:1950:FMb

[19] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 113(2):??, ???? 1950. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981037.

Tobin:1950:SDF

[20] James Tobin. A statistical demand function for food in the U.S.A. Journal of the Royal Statistical Society. Series A (General), 113(2):113–149, ???? 1950. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981038.

Gray:1950:SSS

[21] P. G. Gray and T. Corlett. Sampling for the social survey. Journal of the Royal Statistical Society. Series A (General), 113(2):150–206, ???? 1950. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981039.

Leser:1950:OPH

[22] C. E. V. Leser. Output per head in different parts of the United Kingdom. Journal of the Royal Statistical Society. Series A (General), 113(2):207– 219, ???? 1950. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2981040.

Molland:1950:SFW

[23] R. Molland and G. Evans. Scottish farm wages from 1870 to 1900. Journal of the Royal Statistical Society. Series A (General), 113(2):220–227, ???? 1950. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2981041.

Anscombe:1950:THT

[24] F. J. Anscombe. Table of the hyperbolic transformation sinh^{−1} \sqrt{x}. Journal of the Royal Statistical Society. Series A (General), 113(2):228–229, ???? 1950. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2981042.

McMullen:1950:SNS

[25] L. McMullen. The sources and nature of statistical information in special fields of statistics. Journal of the Royal Statistical Society. Series A (General), 113(2):230–237, ???? 1950. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2981043.

Hough:1950:SNS

[26] J. A. Hough. The sources and nature of statistical information in special fields of statistics: A guide to co-operative statistics. Journal of the Royal Statistical Society. Series A (General), 113 (2):238–248, ???? 1950. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2981044.

B:1950:BRBa

[27] M. S. B. Book review: The Theory of Inbreeding, by R. A. Fisher. Journal of the Royal Statistical Society. Series A (General), 113(2):249–250, ????

1950. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981045.

M:1950:BRBb

[28] P. A. M. Book review: Probability and the Weighing of Evidence, by I. J. Good. Journal of the Royal Statistical Society. Series A (General), 113 (2):250, ???? 1950. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2981046.

V:1950:BRB

[29] S. V. Book review: Statistics, by N. L. Johnson; H. Tetley. Journal of the Royal Statistical Society. Series A (General), 113(2):250–251, ???? 1950. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981047.

G:1950:BRB

[30] E. G. Book review: Forecasts of the Population of the United States, 1945– 75, by P. K. Whelpton; Hope Tisdale Eldridge; Jacob S. Siegel. Journal of the Royal Statistical Society. Series A (General), 113(2):251–252, ???? 1950. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981048.

H:1950:BRBc

[31] J. P. W. H. Book review: Stillbirths: Their Epidemiology and Social Significance, by Ian Sutherland. Journal of the Royal Statistical Society. Series A (General), 113(2):252–253, ???? 1950. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981049.

B:1950:BRBb

[32] F. S. B. Book review: Taxable and Business Income, by D. Throop Smith; J. Keith Butters. Journal of the Royal Statistical Society. Series A (General), 113(2):253–254, ???? 1950. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2981050.

M:1950:BRBc

[33] C. A. M. Book review: London Travel Survey, 1949. Journal of the Royal Statistical Society. Series A (General), 113 (2):254–255, ???? 1950. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2981051.

Anonymous:1950:SNb

[34] Anonymous. Statistical notes. Journal of the Royal Statistical Society. Series A (General), 113(2):256–259, ???? 1950. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2981052.

Anonymous:1950:CNa

[35] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 113(2):260, ???? 1950. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2981053.

Anonymous:1950:SEAb

[36] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society. Series A (General), 113(2):261–264, ???? 1950. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2981054.

Anonymous:1950:LALb

[37] Anonymous. List of additions to the library. Journal of the Royal Statistical Society. Series A (General), 113 (2):265–270, ???? 1950. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981055.

Anonymous:1950:FMc

[38] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 113(3):??, ???? 1950. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2980999.

Searle:1950:CSD

[39] W. F. Searle, E. J. Phillips, and C. J. Martin. Colonial statistics: A discussion before the Royal Statistical Society held on March 22nd, 1950, Mr. H. Campion, C.B., Vice-President, in the Chair. Journal of the Royal Statistical Society. Series A (General), 113(3): 271–298, ???? 1950. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2981000.

Menzler:1950:LPT

[40] F. A. A. Menzler. London and its passenger transport system. Journal of the Royal Statistical Society. Series A (General), 113(3):299–345, ???? 1950. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981001.

Lomax:1950:CPF

[41] K. S. Lomax. Coal production functions for Great Britain. Journal of the Royal Statistical Society. Series A (General), 113(3):346–351, ????

1950. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981002.

Menzler:1950:SNS

[42] F. A. A. Menzler. The sources and nature of statistical information in special fields of statistics: Rail and road statistics. Journal of the Royal Statistical Society. Series A (General), 113 (3):352–375, ???? 1950. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981003.

Anonymous:1950:SNSa

[43] Anonymous. The sources and nature of statistical information in special fields of statistics: Oils and fats statistics. Journal of the Royal Statistical Society. Series A (General), 113(3):376– 395, ???? 1950. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2981004.

Bowley:1950:SNS

[44] Marian Bowley. The sources and nature of statistical information in special fields of statistics: The housing statistics of Great Britain. Journal of the Royal Statistical Society. Series A (General), 113(3):396–411, ???? 1950. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981005.

M:1950:BRBd

[45] P. A. M. Book review: Introduction to the Theory of Probability and Statistics, by N. Arley; K. R. Buch. Journal of the Royal Statistical Society. Series A (General), 113(3):412, ???? 1950. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981006.

M:1950:BRBe

[46] P. A. M. Book review: The Extrapolation, Interpolation and Smoothing of Stationary Time Series, with Engineering Applications, by Norbert Wiener. Journal of the Royal Statistical Society. Series A (General), 113(3):413– 414, ???? 1950. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2981007.

O:1950:BRBa

[47] P. D. O. Book review: Principles of Medical Statistics, by A. Bradford Hill. Journal of the Royal Statistical Society. Series A (General), 113(3):414, ???? 1950. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981008.

N:1950:BRBa

[48] J. W. N. Book review: United Nations: Statistical Year Book 1948, First Year. Journal of the Royal Statistical Society. Series A (General), 113(3): 414–416, ???? 1950. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2981009.

N:1950:BRBb

[49] J. W. N. Book review: United Nations, Demographic Year Book, 1948. Journal of the Royal Statistical Society. Series A (General), 113(3):416–417, ???? 1950. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981010.

N:1950:BRBc

[50] J. W. N. Book review: Nomenclature of Geographic Areas for Statistical Purposes, by United Nations Secretariat.

Journal of the Royal Statistical Society. Series A (General), 113(3):417, ???? 1950. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981011.

N:1950:BRBd

[51] J. W. N. Book review: Proceedings of International Statistical Conference. (September 1947.) Vol. II: United Nations, World Statistical Congress; Proceedings of International Statistical Conference. (September 1947.) Vol. IV: Inter-American Statistical Institute. Journal of the Royal Statistical Society. Series A (General), 113(3): 418, ???? 1950. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2981012.

C:1950:BRBa

[52] C. F. C. Book review: Relative Prices of Exports and Imports of Under-Developed Countries. International Capital Movements during the Inter-War Period. Economic Survey of Latin America. Journal of the Royal Statistical Society. Series A (General), 113(3):418–419, ???? 1950. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2981013.

E:1950:BRB

[53] M. J. E. Book review: Demographic Survey of the British Colonial Empire, by R. R. Kuczynski. Journal of the Royal Statistical Society. Series A (General), 113(3):419–420, ???? 1950. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981014.

P:1950:BRBa

[54] A. R. P. Book review: Patterns of British Life. Journal of the Royal Statistical Society. Series A (General), 113 (3):420–421, ???? 1950. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981015.

Anonymous:1950:SNc

[55] Anonymous. Statistical notes. Journal of the Royal Statistical Society. Series A (General), 113(3):422–423, ???? 1950. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981016.

Anonymous:1950:CNN

[56] Anonymous. Current notes and notices to Fellows. Journal of the Royal Statistical Society. Series A (General), 113 (3):424, ???? 1950. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2981017.

I:1950:SR

[57] L. I. Simon Rowson. Journal of the Royal Statistical Society. Series A (General), 113(3):425, ???? 1950. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981018.

Anonymous:1950:SEAc

[58] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society. Series A (General), 113(3):426–429, ???? 1950. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981019.

Anonymous:1950:LALc

[59] Anonymous. List of additions to the library. Journal of the Royal Statistical Society. Series A (General), 113 (3):430–434, ???? 1950. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2981020.

Anonymous:1950:BMb

[60] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 113(3):??, ???? 1950. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2981021.

Anonymous:1950:VI

[61] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (General), 113(4):613– 615, ???? 1950. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2980864.

Anonymous:1950:FMd

[62] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 113(4):??, ???? 1950. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2980865.

Reddaway:1950:MRP

[63] W. B. Reddaway. Movements in the real product of the United Kingdom, 1946–1949. Journal of the Royal Statistical Society. Series A (General), 113 (4):435–463, ???? 1950. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2980866.

Shone:1950:SNS

[64] Robert M. Shone. The sources and nature of statistical information in special fields of statistics: Statistics relating to the U.K. iron and steel industry. Journal of the Royal Statistical Society. Series A (General), 113(4):464–486, ???? 1950. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2980867.

Anonymous:1950:SNSb

[65] Anonymous. The sources and nature of statistical information in special fields of statistics: U.K. tobacco statistics. Journal of the Royal Statistical Society. Series A (General), 113(4):487– 508, ???? 1950. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2980868.

Daniel:1950:SNS

[66] G. H. Daniel. The sources and nature of statistical information in special fields of statistics: Electricity and gas statistics. Journal of the Royal Statistical Society. Series A (General), 113 (4):509–530, ???? 1950. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2980869.

Lyle:1950:SNS

[67] Philip Lyle. The sources and nature of statistical information in special fields of statistics: The sugar industry. Journal of the Royal Statistical Society. Series A (General), 113(4):531–543, ???? 1950. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2980870.

Anonymous:1950:WP

[68] Anonymous. Wholesale prices in 1949. Journal of the Royal Statistical Society. Series A (General), 113(4):544– 558, ???? 1950. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2980871.

Anonymous:1950:RCF

[69] Anonymous. Report of the Council for the financial year ended December 31st, 1949, and for the sessional year ending June 21st, 1950, presented at the one hundred and sixteenth Annual General Meeting of the Royal Statistical Society, held at the London School of Hygiene and Tropical Medicine, W.C. 1, on June 21st, 1950. Journal of the Royal Statistical Society. Series A (General), 113(4): 559–575, ???? 1950. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2980872.

Anonymous:1950:POH

[70] Anonymous. Proceedings of the one hundred and sixteenth Annual General Meeting of the Royal Statistical Society, held at the London School of Hygiene and Tropical Medicine, on Wednesday, June 21st, 1950. Journal of the Royal Statistical Society. Series A (General), 113(4):576, ???? 1950. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2980873.

K:1950:BRBc

[71] M. G. K. Book review: Experimental Designs, by William G. Cochran; Gertrude M. Cox. Journal of the Royal Statistical Society. Series A (General), 113(4):577–578, ???? 1950. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2980874.

M:1950:BRBf

[72] P. A. M. Book review: A First Course in Probability and Statistics, by J. Neyman. Journal of the Royal Statistical Society. Series A (General), 113(4): 578, ???? 1950. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2980875.

C:1950:BRBb

[73] C. F. C. Book review: An Introduction to the Theory of Statistics, by G. Udny Yule; M. G. Kendall. Journal of the Royal Statistical Society. Series A (General), 113(4):578–580, ???? 1950. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2980876.

P:1950:BRBb

[74] R. L. P. Book review: Chance and Choice by Cardpack and Chessboard, by Lancelot Hogben. Journal of the Royal Statistical Society. Series A (General), 113(4):580–582, ???? 1950. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2980877.

A:1950:BRB

[75] P. A. Book review: The Science of Chance, by H. C. Levinson. Journal of the Royal Statistical Society. Series A (General), 113(4):582, ???? 1950. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2980878.

F:1950:BRB

[76] D. J. F. Book review: Les Méthodes Statistiques dans L’Expérimentation Biologique, by Ph. L’Héritier. Journal of the Royal Statistical Society. Series A (General), 113(4):583, ???? 1950. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2980879.

M:1950:BRBg

[77] P. A. M. Book review: Le Calcul des Probabilités et ses Applications. Journal of the Royal Statistical Society. Series A (General), 113(4):583–584, ???? 1950. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2980880.

J:1950:BRB

[78] N. L. J. Book review: Complementi di Calcolo Delle Probabilità, by Giuseppe Pompilj. Journal of the Royal Statistical Society. Series A (General), 113 (4):584, ???? 1950. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2980881.

H:1950:BRBd

[79] H. O. H. Book review: Nomography, by Alexander S. Levens. Journal of the Royal Statistical Society. Series A (General), 113(4):584, ???? 1950. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2980882.

S:1950:BRBb

[80] B. B. S. Book review: Human Ability: A Continuation of the Abilities of Man, by C. Spearman; Ll. Wynn Jones. Journal of the Royal Statistical Society. Series A (General), 113(4): 584–586, ???? 1950. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2980883.

C:1950:BRBc

[81] V. R. C. Book review: The Educational Development of Children, by W. Glassey; E. J. Weeks; The Measurement of Abilities. by P. E. Vernon; Personnel Selection in the British Forces. by P. E. Vernon; J. B. Parry. Journal of the Royal Statistical Society. Series A (General), 113(4):586–588, ???? 1950. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2980884.

O:1950:BRBb

[82] P. D. O. Book review: Researches on the Measurement of Human Performance, by N. H. Mackworth. Journal of the Royal Statistical Society. Series A (General), 113(4):588–589, ???? 1950. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2980885.

H:1950:BRBe

[83] J. P. W. H. Book review: Epidemics in Schools: An Analysis of the Data Collected During the Years 1935 to 1939, by E. A. Cheeseman. Journal of the Royal Statistical Society. Series A (General), 113(4):589–590, ???? 1950. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2980886.

H:1950:BRBf

[84] J. P. W. H. Book review: Recent Advances in Social Medicine, by A. C. Stevenson; E. A. Cheeseman. Journal of the Royal Statistical Society. Series A (General), 113(4):590–591, ???? 1950. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2980887.

S:1950:BRBc

[85] C. T. S. Book review: The Balance of Payments and the Standard of Living, by R. G. Hawtrey. Journal of the Royal Statistical Society. Series A (General), 113(4):591–592, ???? 1950. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2980888.

S:1950:BRBd

[86] D. S. Book review: Investment and Inflation, with Special Reference to the Immediate Post-War Period, Canada, 1945–1948. Journal of the Royal Statistical Society. Series A (General), 113 (4):592–593, ???? 1950. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2980889.

W:1950:BRB

[87] L. T. W. Book review: The Home Market, by G. Harrison; F. C. Mitchell; M. A. Abrams. Journal of the Royal Statistical Society. Series A (General), 113 (4):593–594, ???? 1950. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2980890.

Anonymous:1950:SNd

[88] Anonymous. Statistical notes. Journal of the Royal Statistical Society. Series A (General), 113(4):595–597, ???? 1950. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2980891.

Anonymous:1950:CNb

[89] Anonymous. Current note. Journal of the Royal Statistical Society. Series A (General), 113(4):598, ???? 1950. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2980892.

Anonymous:1950:SEAd

[90] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society. Series A (General), 113(4):599–603, ???? 1950. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2980893.

Anonymous:1950:LALd

[91] Anonymous. List of additions to the library. Journal of the Royal Statistical Society. Series A (General), 113 (4):604–607, ???? 1950. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2980894.

Anonymous:1950:PRb

[92] Anonymous. Periodical returns. Journal of the Royal Statistical Society. Series A (General), 113(4):608–611, ???? 1950. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2980895.

Anonymous:1950:BMc

[93] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 113(4):??, ???? 1950. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2980896.

Anonymous:1951:FMa

[94] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 114(1):??, ???? 1951. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2980897.

Anonymous:1951:DDP

[95] Anonymous. The Doctor’s day and pay. Journal of the Royal Statistical Society. Series A (General), 114(1):1–34, ???? 1951. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2980898.

Anonymous:1951:PM

[96] Anonymous. Proceedings of the meeting. Journal of the Royal Statistical Society. Series A (General), 114(1):34– 37, ???? 1951. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2980899.

Anonymous:1951:DRR

[97] Anonymous. A discussion on the reports of the Royal Commission on Population. Journal of the Royal Statistical Society. Series A (General), 114(1):38– 49, ???? 1951. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2980900.

Palca:1951:ECE

[98] H. Palca and I. G. R. Davies. Earnings and conditions of employment in agriculture. Journal of the Royal Statistical Society. Series A (General), 114(1):50– 58, ???? 1951. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2980901.

Britton:1951:SNS

[99] D. K. Britton and K. E. Hunt. The sources and nature of statistical information in special fields of statistics: The agricultural statistics of the United Kingdom. Journal of the Royal Statistical Society. Series A (General), 114 (1):59–98, ???? 1951. CODEN JS-SAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2980902.

Moran:1951:BRBa

[100] P. A. P. Moran. Book review: Rank Correlation Methods, by M. G. Kendall. Journal of the Royal Statistical Society. Series A (General), 114 (1):99–100, ???? 1951. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2980903.

David:1951:BRBa

[101] F. N. David. Book review: Some Theory of Sampling, by W. Edwards Deming. Journal of the Royal Statistical Society. Series A (General), 114(1):100, ???? 1951. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2980904.

Irwin:1951:BRBa

[102] J. O. Irwin. Book review: Analysis and Design of Experiments: Analysis of Variance and Analysis of Variance Designs, by H. B. Mann. Journal of the Royal Statistical Society. Series A (General), 114(1):100–101, ???? 1951. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2980905.

Bailey:1951:BRB

[103] Norman T. J. Bailey. Book review: Biological Standardization, by J. H. Burn; D. J. Finney; L. G. Goodwin. Journal of the Royal Statistical Society. Series A (General), 114(1):101–102, ???? 1951. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2980906.

Irwin:1951:BRBb

[104] J. O. Irwin. Book review: Reports on Biological Standards. VI. The Design of Toxicity Tests, by W. L. M. Perry. Journal of the Royal Statistical Society. Series A (General), 114(1):102, ???? 1951. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2980907.

Smith:1951:BRB

[105] B. Babington Smith. Book review: Fundamental Statistics in Psychology and Education, by J. P. Guilford. Journal of the Royal Statistical Society. Series A (General), 114(1):102–103, ???? 1951. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2980908.

Moser:1951:BRB

[106] C. A. Moser. Book review: Economic Statistics, by Lewis A. Maverick. Journal of the Royal Statistical Society. Series A (General), 114(1):103, ???? 1951. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2980909.

Kendall:1951:BRBa

[107] D. G. Kendall. Book review: The Life and Works of A. K. Erlang, by E. Brockmeyer; H. L. Halstrøm; Arne Jensen. Journal of the Royal Statistical Society. Series A (General), 114 (1):103–104, ???? 1951. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2980910.

Kendall:1951:BRBb

[108] D. G. Kendall. Book review: Moe’s Principle, by Arne Jensen. Journal of the Royal Statistical Society.

Series A (General), 114(1):104, ???? 1951. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2980911.

Abrams:1951:BRB

[109] M. Abrams. Book review: The Distribution of Consumer Goods, by James B. Jeffreys; Margaret Maccoll; G. L. Levitt. Journal of the Royal Statistical Society. Series A (General), 114 (1):104–105, ???? 1951. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2980912.

Wilkins:1951:BRB

[110] L. T. Wilkins. Book review: Principles of Market Research, by A. H. R. Delens. Journal of the Royal Statistical Society. Series A (General), 114 (1):105–106, ???? 1951. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2980913.

Good:1951:BRB

[111] I. J. Good. Book review: Calculating Instruments and Machines, by D. R. Hartree. Journal of the Royal Statistical Society. Series A (General), 114 (1):106–107, ???? 1951. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2980914.

Grebenik:1951:BRB

[112] E. Grebenik. Book review: Statistics Without Numbers: Visual Explanation, by F. Kafka. Journal of the Royal Statistical Society. Series A (General), 114 (1):108, ???? 1951. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2980915.

Benjamin:1951:BRB

[113] B. Benjamin. Book review: Social Insurance in Britain, by Charles E. Clarke. Journal of the Royal Statistical Society. Series A (General), 114 (1):108–109, ???? 1951. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2980916.

Reddaway:1951:BRB

[114] W. B. Reddaway. Book review: Changes in the Cost of Living and the Distribution of Income Since 1938, by Dudley Seers. Journal of the Royal Statistical Society. Series A (General), 114 (1):109–110, ???? 1951. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2980917.

Thomas:1951:BRB

[115] Brinley Thomas. Book review: Historical Statistics of the United States, 1789–1945: A Supplement to the Statistical Abstract of the United States. Journal of the Royal Statistical Society. Series A (General), 114(1):110, ???? 1951. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2980918.

Elderton:1951:BRB

[116] W. P. Elderton. Book review: Urban Mortgage Lending by Life Insurance Companies, by R. J. Saulnier. Journal of the Royal Statistical Society. Series A (General), 114(1):110–112, ???? 1951. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2980919.

Nixon:1951:BRBa

[117] J. W. Nixon. Book review: Economic Survey of Asia and the Far East, 1949; The Economic Development of Latin America and Its Principal Problems. Journal of the Royal Statistical Society. Series A (General), 114(1):112, ???? 1951. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2980920.

Thompson:1951:BRB

[118] R. J. Thompson. Book review: The Socialized Agriculture of the U.S.S.R.: Plans and Performance, by Naum Jasny. Journal of the Royal Statistical Society. Series A (General), 114(1): 113–114, ???? 1951. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2980921.

Kemsley:1951:BRB

[119] W. F. F. Kemsley. Book review: A Social Survey of Singapore, 1947. Journal of the Royal Statistical Society. Series A (General), 114(1):114–115, ???? 1951. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2980922.

Anonymous:1951:ONP

[120] Anonymous. Other new publications. Journal of the Royal Statistical Society. Series A (General), 114(1):115, ???? 1951. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2980923.

Anonymous:1951:SNa

[121] Anonymous. Statistical notes. Journal of the Royal Statistical Society. Series A (General), 114(1):116–118, ???? 1951. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2980924.

Anonymous:1951:CNa

[122] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 114(1):119, ???? 1951. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2980925.

Hartley:1951:LJC

[123] H. O. Hartley. Leslie John Comrie. Journal of the Royal Statistical Society. Series A (General), 114(1):120– 121, ???? 1951. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2980926.

Anonymous:1951:SEAa

[124] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society. Series A (General), 114(1):122–128, ???? 1951. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2980927.

Anonymous:1951:LALa

[125] Anonymous. List of additions to the library. Journal of the Royal Statistical Society. Series A (General), 114 (1):129–134, ???? 1951. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2980928.

Anonymous:1951:PRa

[126] Anonymous. Periodical returns. Journal of the Royal Statistical Society. Series A (General), 114(1):135–137, ???? 1951. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2980929.

Anonymous:1951:FMb

[127] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 114(2):??, ???? 1951. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2980973.

Grunhut:1951:SC

[128] M. Grünhut. Statistics in criminology. Journal of the Royal Statistical Society. Series A (General), 114(2):139– 162, ???? 1951. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2980974.

Durbin:1951:DRR

[129] J. Durbin and A. Stuart. Differences in response rates of experienced and inexperienced interviewers. Journal of the Royal Statistical Society. Series A (General), 114(2):163–206, ???? 1951. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2980975.

Ridley:1951:NEP

[130] T. M. Ridley. Note on the extent of the public sector of the economy in recent years. Journal of the Royal Statistical Society. Series A (General), 114 (2):207–213, ???? 1951. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2980976.

Seng:1951:HSD

[131] You Poh Seng. Historical survey of the development of sampling theories and practice. Journal of the Royal Statistical Society. Series A (General), 114 (2):214–231, ???? 1951. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2980977.

Morrell:1951:SNS

[132] A. J. H. Morrell, F. P. Stevens, and E. W. Talbot. The sources and nature of statistical information in special fields of statistics: Statistics of the chemical industry. Journal of the Royal Statistical Society. Series A (General), 114(2):232–248, ???? 1951. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2980978.

Kendall:1951:BRBc

[133] David G. Kendall. Book review: An Introduction to Probability Theory and its Applications, by William Feller. Journal of the Royal Statistical Society. Series A (General), 114(2):249–250, ???? 1951. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2980979.

Lindley:1951:BRB

[134] D. V. Lindley. Book review: Statistical Decision Functions, by Abraham Wald. Journal of the Royal Statistical Society. Series A (General), 114(2): 250–253, ???? 1951. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2980980.

Vajda:1951:BRBa

[135] S. Vajda. Book review: Statistics: An Intermediate Text-Book, by N. L. Johnson; H. Tetley. Journal of the Royal Statistical Society. Series A (General), 114(2):253–254, ???? 1951. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2980981.

Anscombe:1951:BRB

[136] F. J. Anscombe. Book review: Introductory Statistics, by M. H. Quenouille.

Journal of the Royal Statistical Society. Series A (General), 114(2):254, ???? 1951. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2980982.

Armitage:1951:BRBa

[137] P. Armitage. Book review: The Place of Statistical Methods in Biological and Chemical Experimentation, by E. J. de Beer. Journal of the Royal Statistical Society. Series A (General), 114(2): 255, ???? 1951. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2980983.

Davies:1951:BRB

[138] O. L. Davies. Book review: Quality Control and Statistical Methods, by Edward M. Schrock. Journal of the Royal Statistical Society. Series A (General), 114(2):255–256, ???? 1951. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2980984.

Wishart:1951:BRB

[139] J. Wishart. Book review: Tables of the Binomial Probability Distribution. Journal of the Royal Statistical Society. Series A (General), 114(2):256– 257, ???? 1951. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2980985.

Newman:1951:BRB

[140] Dennis Newman. Book review: Acceptance Sampling: A Symposium. Journal of the Royal Statistical Society. Series A (General), 114(2):257, ???? 1951. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2980986.

Armitage:1951:BRBb

[141] P. Armitage. Book review: Demography, by P. R. Cox. Journal of the Royal Statistical Society. Series A (General), 114(2):257–258, ???? 1951. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2980987.

Knowelden:1951:BRB

[142] J. Knowelden. Book review: Population Census Methods. Journal of the Royal Statistical Society. Series A (General), 114(2):258–260, ???? 1951. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2980988.

Nixon:1951:BRBb

[143] J. W. Nixon. Book review: Annual Bulletin of Transport Statistics. First Year: 1949. Journal of the Royal Statistical Society. Series A (General), 114 (2):260, ???? 1951. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2980989.

Johnson:1951:BRB

[144] Harry G. Johnson. Book review: The Dynamics of Business Cycles; A Study in Economic Fluctuations, by Jan Tinbergen; J. J. Polak. Journal of the Royal Statistical Society. Series A (General), 114(2):260–261, ???? 1951. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2980990.

Zentler:1951:BRB

[145] A. P. Zentler. Book review: Inventories and Business Cycles, with Special Reference to Manufacturers’ Inventories, by Moses Abramowitz. Journal of the Royal Statistical Society. Series A (General), 114(2):261–263, ???? 1951. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2980991.

Rostas:1951:BRB

[146] L. Rostas. Book review: Taxes, the Public Debt and Transfers of Income, by Donald C. Miller. Journal of the Royal Statistical Society. Series A (General), 114(2):263–264, ???? 1951. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2980992.

Daniell:1951:BRB

[147] P. A. Daniell. Book review: Economics for Commercial Students and Business Men, by Albert Crew; Leo T. Little. Journal of the Royal Statistical Society. Series A (General), 114(2):265, ???? 1951. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2980993.

Deane:1951:BRB

[148] P. M. Deane. Book review: The Conditions of Economic Progress, by Colin Clark. Journal of the Royal Statistical Society. Series A (General), 114(2): 265–266, ???? 1951. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2980994.

Richardson:1951:BRB

[149] J. Henry Richardson. Book review: Planning in Practice, by Ely Devons. Journal of the Royal Statistical Society. Series A (General), 114(2):266– 268, ???? 1951. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2980995.

Anonymous:1951:SNb

[150] Anonymous. Statistical notes. Journal of the Royal Statistical Society. Series A (General), 114(2):269–272, ???? 1951. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2980996.

Anonymous:1951:SEAb

[151] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society. Series A (General), 114(2):273–279, ???? 1951. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2980997.

Anonymous:1951:LALb

[152] Anonymous. List of additions to the library. Journal of the Royal Statistical Society. Series A (General), 114 (2):280–286, ???? 1951. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2980998.

Anonymous:1951:FMc

[153] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 114(3):??, ???? 1951. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2980778.

Martin:1951:CTM

[154] W. J. Martin. A comparison of the trends of male and female mortality. Journal of the Royal Statistical Society. Series A (General), 114(3):287– 306, ???? 1951. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2980779.

Morgan:1951:IPI

[155] D. J. Morgan and W. J. Corlett. The influence of price in international trade: A study in method. Journal of the Royal Statistical Society. Series A (General), 114(3):307–358, ???? 1951. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2980780.

Houthakker:1951:SCE

[156] H. S. Houthakker. Some calculations on electricity consumption in Great Britain. Journal of the Royal Statistical Society. Series A (General), 114 (3):359–371, ???? 1951. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2980781.

Scott:1951:BAM

[157] A. D. Scott. Bibliography of applications of mathematical statistics to economics, 1943–1949. Journal of the Royal Statistical Society. Series A (General), 114(3):372–393, ???? 1951. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2980782.

Ford:1951:SNS

[158] C. V. Ford. The sources and nature of statistical information in special fields of statistics: Statistics of the British motor industry. Journal of the Royal Statistical Society. Series A (General), 114(3):394–407, ???? 1951. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2980783.

Anonymous:1951:WP

[159] Anonymous. Wholesale prices in 1950. Journal of the Royal Statistical Society. Series A (General), 114(3):408– 422, ???? 1951. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2980784.

Nixon:1951:BRBc

[160] J. W. Nixon. Book review: Index Numbers of Industrial Production: Studies in Methods No. 1. Journal of the Royal Statistical Society. Series A (General), 114(3):423, ???? 1951. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2980785.

Seers:1951:BRB

[161] Dudley Seers. Book review: Studies in Income and Wealth. Vol. XI: Conference on Research in Income an Wealth. Journal of the Royal Statistical Society. Series A (General), 114(3):423– 425, ???? 1951. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2980786.

Penrice:1951:BRB

[162] G. Penrice. Book review: Economic, Fluctuations in the United States 1921– 41, by Lawrence R. Klein. Journal of the Royal Statistical Society. Series A (General), 114(3):425–427, ???? 1951. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2980787.

Saunders:1951:BRB

[163] C. T. Saunders. Book review: The Post-War Financial Problem and Other Essays, by F. W. Paish. Journal of the Royal Statistical Society. Series A (General), 114(3):427–428, ???? 1951. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2980788.

Roy:1951:BRB

[164] A. D. Roy. Book review: On the Accuracy of Economic Observations, by Oskar Morgenstern. Journal of the Royal Statistical Society. Series A (General), 114(3):428–430, ???? 1951. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2980789.

Armitage:1951:BRBc

[165] P. Armitage. Book review: A Bibliography of Statistical Quality Control: Supplement, by Grant I. Butterbaugh. Journal of the Royal Statistical Society. Series A (General), 114(3):430, ???? 1951. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2980790.

Anonymous:1951:SNc

[166] Anonymous. Statistical notes. Journal of the Royal Statistical Society. Series A (General), 114(3):431–434, ???? 1951. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2980791.

Anonymous:1951:CNb

[167] Anonymous. Current note. Journal of the Royal Statistical Society. Series A (General), 114(3):434, ???? 1951. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2980792.

M:1951:SGE

[168] F. A. A. M. Sir George Epps, K.B.E., C.B. Journal of the Royal Statistical Society. Series A (General), 114(3): 435, ???? 1951. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2980793.

Anonymous:1951:SEAc

[169] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society. Series A (General), 114(3):436–441, ???? 1951. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2980794.

Anonymous:1951:LALc

[170] Anonymous. List of additions to the library. Journal of the Royal Statistical Society. Series A (General), 114 (3):442–446, ???? 1951. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2980795.

Anonymous:1951:VI

[171] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (General), 114(4):605– 607, ???? 1951. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2981084.

Anonymous:1951:FMd

[172] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 114(4):??, ???? 1951. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2981085.

Stafford:1951:IWP

[173] J. Stafford. Indices of wholesale prices. Journal of the Royal Statistical Society. Series A (General), 114(4):447– 467, ???? 1951. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2981086.

Deane:1951:SNS

[174] Marjorie Deane. The sources and nature of statistical information in special fields of statistics: United Kingdom publishing statistics. Journal of the Royal Statistical Society. Series A (General), 114(4):468–489, ???? 1951. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981087.

Clark:1951:UPD

[175] Colin Clark. Urban population densities. Journal of the Royal Statistical Society. Series A (General), 114(4):490– 496, ???? 1951. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2981088.

Anonymous:1951:RAM

[176] Anonymous. Recent advances in mathematical statistics: Bibliography, 1943–47. Journal of the Royal Statistical Society. Series A (General), 114 (4):497–558, ???? 1951. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981089.

Anonymous:1951:RCF

[177] Anonymous. Report of the Council for the financial year ended December 31st, 1950, and for the sessional year ended June 27th, 1951, presented at the one hundred and seventeenth Annual General Meeting of the Royal Statistical Society, held at the London School of Hygiene and Tropical Medicine, W.C. 1, on June 27th, 1951. Journal of the Royal Statistical Society. Series A (General), 114(4):559– 575, ???? 1951. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2981090.

Anonymous:1951:POH

[178] Anonymous. Proceedings of the one hundred and seventeenth Annual General Meeting of the Royal Statistical Society, held at the London School of Hygiene and Tropical Medicine, on Wednesday, June 27th, 1951. Journal of the Royal Statistical Society. Series A (General), 114(4):576, ???? 1951. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2981091.

Vajda:1951:BRBb

[179] S. Vajda. Book review: Contribution to the Theory of Games, by H. W. Kuhn; A. W. Tucker. Journal of the Royal Statistical Society. Series A (General), 114 (4):577–578, ???? 1951. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2981092.

David:1951:BRBb

[180] F. N. David. Book review: Theory of Probability, by M. E. Munroe. Journal of the Royal Statistical Society. Series A (General), 114(4):578, ???? 1951. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2981093.

David:1951:BRBc

[181] F. N. David. Book review: Calcul des Probabilités, by R. Fortet. Journal of the Royal Statistical Society. Series A (General), 114(4):578–579, ???? 1951. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2981094.

Moran:1951:BRBb

[182] P. A. Moran. Book review: Hypothesis Testing in Time Series Analysis, by Peter Whittle. Journal of the Royal Statistical Society. Series A (General), 114(4):579, ???? 1951. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981095.

Irwin:1951:BRBc

[183] J. O. Irwin. Book review: Psychology and the Industrial Worker, by E. G. Chambers. Journal of the Royal Statistical Society. Series A (General), 114 (4):579–580, ???? 1951. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981096.

Carter:1951:BRBa

[184] C. F. Carter. Book review: Cyclical Movement in the Balance of Payments, by Tse Chun Chang. Journal of the Royal Statistical Society. Series A (General), 114(4):580–581, ???? 1951. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981097.

Cox:1951:BRB

[185] P. R. Cox. Book review: U.S. Bureau of the Census. Handbook of Statistical Methods for Demographers, by A. J. Jaffe. Journal of the Royal Statistical Society. Series A (General), 114 (4):581–582, ???? 1951. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981098.

Nixon:1951:BRBd

[186] J. W. Nixon. Book review: United Nations Department of Economic Affairs. Economic Survey of Europe in 1950, by Research and Planning Division of the Economic Commission for Europe. Journal of the Royal Statistical Society. Series A (General), 114 (4):582–583, ???? 1951. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2981099.

Langley:1951:BRB

[187] Kathleen M. Langley. Book review: Conference on Research in Income and Wealth. Studies in Income and Wealth. Journal of the Royal Statistical Society. Series A (General), 114(4):583– 584, ???? 1951. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2981100.

Kenny:1951:BRB

[188] T. Kenny. Book review: Scientific Method for Auditing, by Lawrence L. Vance. Journal of the Royal Statistical Society. Series A (General), 114 (4):584–586, ???? 1951. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2981101.

Carter:1951:BRBb

[189] C. F. Carter. Book review: Fundamentals of Governmental Accounting, by Lloyd Morey; Robert P. Hackett. Journal of the Royal Statistical Society. Series A (General), 114(4):586–587, ???? 1951. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2981102.

Daniel:1951:BRB

[190] G. H. Daniel. Book review: Domestic Utilization of Heating Applicances and Expenditures on Fuels in 1948/1949, by Leslie T. Wilkins. Journal of the Royal Statistical Society. Series A (General), 114(4):587, ???? 1951. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2981103.

Anonymous:1951:SNd

[191] Anonymous. Statistical notes. Journal of the Royal Statistical Society. Series A (General), 114(4):588–590, ???? 1951. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981104.

Anonymous:1951:CNc

[192] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 114(4):591, ???? 1951. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981105.

Anonymous:1951:SEAd

[193] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society. Series A (General), 114(4):592–595, ???? 1951. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981106.

Anonymous:1951:LALd

[194] Anonymous. List of additions to the library. Journal of the Royal Statistical Society. Series A (General), 114 (4):596–600, ???? 1951. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981107.

Anonymous:1951:PRb

[195] Anonymous. Periodical returns. Journal of the Royal Statistical Society. Series A (General), 114(4):601–604, ???? 1951. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981108.

Anonymous:1952:FMa

[196] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 115(1):??, ???? 1952. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2981109.

Houthakker:1952:EFB

[197] H. S. Houthakker. The econometrics of family budgets. Journal of the Royal Statistical Society. Series A (General), 115(1):1–28, ???? 1952. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2981110.

Barna:1952:IBE

[198] T. Barna. The interdependence of the British economy. Journal of the Royal Statistical Society. Series A (General), 115(1):29–81, ???? 1952. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2981111.

Carter:1952:INR

[199] C. F. Carter. Index numbers of the real product of the United Kingdom. Journal of the Royal Statistical Society. Series A (General), 115(1):82–125, ???? 1952. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2981112.

Anonymous:1952:TSS

[200] Anonymous. The teaching of statistics in schools. Journal of the Royal Statistical Society. Series A (General), 115 (1):126–137, ???? 1952. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2981113.

Plackett:1952:BRB

[201] R. L. Plackett. Book review: Introduction to Statistical Analysis, by Wilfred J. Dixon; Frank J. Massey. Journal of the Royal Statistical Society. Series A (General), 115(1):138, ???? 1952. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981114.

Armitage:1952:BRBa

[202] P. Armitage. Book review: Introduction to Statistical Method, by B. C. Brookes; W. F. L. Dick. Journal of the Royal Statistical Society. Series A (General), 115(1):139, ???? 1952. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981115.

Booker:1952:BRB

[203] H. S. Booker. Book review: Statistics for Economics and Business, by D. W. Paden; E. F. Lindquist. Journal of the Royal Statistical Society. Series A (General), 115(1):139–140, ???? 1952. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981116.

Hartley:1952:BRBa

[204] H. O. Hartley. Book review: Values and Integrals of the Orthogonal Polynomials up to n = 26, by D. B. DeLury. Journal of the Royal Statistical Society. Series A (General), 115(1):140–141, ???? 1952. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981117.

Vajda:1952:BRBa

[205] S. Vajda. Book review: Drei Perlen der Zahlentheorie, by A. J. Chintschin.

Journal of the Royal Statistical Society. Series A (General), 115(1):141– 142, ???? 1952. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2981118.

Vajda:1952:BRBb

[206] S. Vajda. Book review: Statistische Methoden für Naturwissenschafter, Mediziner und Ingenieure, by von Arthur Linder. Journal of the Royal Statistical Society. Series A (General), 115(1):142, ???? 1952. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981119.

Ridley:1952:BRB

[207] T. M. Ridley. Book review: Det Statistiske Department, 1920–50; Fortegnelse over Norges Offisielle Statistik, 1828–1950; Statistisk Sentralbyrå Gjennom 75 År. Tre 25-Årsmeldinger. Journal of the Royal Statistical Society. Series A (General), 115(1):143, ???? 1952. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981120.

Kemsley:1952:BRBa

[208] W. F. F. Kemsley. Book review: De Juiste Maat, by J. Sittig; H. Freudenthal; L. Stafleu. Journal of the Royal Statistical Society. Series A (General), 115(1):143–144, ???? 1952. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2981121.

Rix:1952:BRBa

[209] M. S. Rix. Book review: The A B C of the Foreign Exchanges, by Norman Crump. Journal of the Royal Statistical Society. Series A (General), 115 (1):144–145, ???? 1952. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981122.

Daniel:1952:BRB

[210] G. H. Daniel. Book review: Lessons of the British War Economy, by D. N. Chester. Journal of the Royal Statistical Society. Series A (General), 115 (1):145–146, ???? 1952. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981123.

Butler:1952:BRBa

[211] I. G. Butler. Book review: Corporate Income Retention 1915–43, by Sergei P. Dobrovolsky. Journal of the Royal Statistical Society. Series A (General), 115(1):146–147, ???? 1952. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2981124.

Anstey:1952:BRB

[212] Vera Anstey. Book review: The Population of India and Pakistan, by Kingsley Davis. Journal of the Royal Statistical Society. Series A (General), 115 (1):147–149, ???? 1952. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981125.

Jones:1952:BRB

[213] D. Caradog Jones. Book review: Poverty and the Welfare State, by B. Seebohm Rowntree; G. R. Lavers. Journal of the Royal Statistical Society. Series A (General), 115(1):149– 150, ???? 1952. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2981126.

Oades:1952:BRB

[214] R. C. Oades. Book review: Coastwise Shipping and the Small Ports, by P. Ford; J. A. Bound. Journal of the Royal Statistical Society. Series A (General), 115(1):150–151, ???? 1952. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2981127.

Anonymous:1952:SNa

[215] Anonymous. Statistical notes. Journal of the Royal Statistical Society. Series A (General), 115(1):152–155, ???? 1952. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2981128.

Anonymous:1952:CNa

[216] Anonymous. Current note. Journal of the Royal Statistical Society. Series A (General), 115(1):155, ???? 1952. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2981129.

Kendall:1952:GUY

[217] M. G. Kendall. George Udny Yule C.B.E., F.R.S. Journal of the Royal Statistical Society. Series A (General), 115(1):156–161, ???? 1952. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2981130.

Anonymous:1952:SEAa

[218] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society. Series A (General), 115(1):162–167, ???? 1952. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2981131.

Anonymous:1952:LALa

[219] Anonymous. List of additions to the library. Journal of the Royal Statistical Society. Series A (General), 115 (1):168–173, ???? 1952. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981132.

Anonymous:1952:PRa

[220] Anonymous. Periodical returns. Journal of the Royal Statistical Society. Series A (General), 115(1):174–176, ???? 1952. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981133.

Anonymous:1952:BMa

[221] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 115(1):??, ???? 1952. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981134.

Anonymous:1952:FMb

[222] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 115(2):??, ???? 1952. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981153.

Rudd:1952:EER

[223] Ernest Rudd. Estimates of expenditure on road transport in Great Britain. Journal of the Royal Statistical Society. Series A (General), 115(2):179– 198, ???? 1952. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2981154.

Peterson:1952:SG

[224] A. W. Peterson. The statistics of gambling. Journal of the Royal Statistical Society. Series A (General), 115(2): 199–218, ???? 1952. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2981155.

Richardson:1952:CDQ

[225] Lewis F. Richardson. Contiguity and deadly quarrels: The local pacifying influence. Journal of the Royal Statistical Society. Series A (General), 115 (2):219–231, ???? 1952. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981156.

Booker:1952:DRO

[226] H. S. Booker and S. T. David. Differences in results obtained by experienced and inexperienced interviewers. Journal of the Royal Statistical Society. Series A (General), 115(2):232– 257, ???? 1952. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2981157.

Abrams:1952:SNS

[227] Mark Abrams. The sources and nature of statistical information in special fields of statistics: Statistics of advertising. Journal of the Royal Statistical Society. Series A (General), 115 (2):258–264, ???? 1952. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981158.

Herbert:1952:SNS

[228] A. G. Herbert and R. D. Clarke. The sources and nature of statistical information in special fields of statistics: The statistics of British insurance. Journal of the Royal Statistical Society. Series A (General), 115(2): 265–280, ???? 1952. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2981159.

Nixon:1952:ISI

[229] J. W. Nixon. International Statistical Institute 27th session, India, 1951. Journal of the Royal Statistical Society. Series A (General), 115(2):281– 282, ???? 1952. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2981160.

Good:1952:BRB

[230] I. J. Good. Book review: A Treatise on Induction and Probability, by G. H. von Wright. Journal of the Royal Statistical Society. Series A (General), 115 (2):283–285, ???? 1952. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981161.

Vajda:1952:BRBc

[231] S. Vajda. Book review: Statistical Design and Analysis of Experiments for Development Research, by D. S. Villars. Journal of the Royal Statistical Society. Series A (General), 115(2):285– 286, ???? 1952. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2981162.

Johnson:1952:BRB

[232] N. L. Johnson. Book review: Confidence Limits for the Hypergeometric Distribution, by J. H. Chung; D. B. Delury. Journal of the Royal Statistical Society. Series A (General), 115 (2):286, ???? 1952. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2981163.

Benjamin:1952:BRBa

[233] B. Benjamin. Book review: Hospital Morbidity Statistics, by Donald Mackay. Journal of the Royal Statistical Society. Series A (General), 115 (2):286–289, ???? 1952. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2981164.

Cox:1952:BRBa

[234] P. R. Cox. Book review: La Recente Inversione nella Tendenza Della Natalità, by Bernardo Colombo. Journal of the Royal Statistical Society. Series A (General), 115(2):289–290, ???? 1952. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2981165.

Kemsley:1952:BRBb

[235] W. F. F. Kemsley. Book review: Social Surveys and Social Action, by M. Abrams. Journal of the Royal Statistical Society. Series A (General), 115 (2):290–291, ???? 1952. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2981166.

Matthews:1952:BRB

[236] R. C. O. Matthews. Book review: What Happens During Business Cycles: A Progress Report, by Wesley C. Mitchell. Journal of the Royal Statistical Society. Series A (General), 115(2):291– 292, ???? 1952. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2981167.

Roy:1952:BRBa

[237] A. D. Roy. Book review: Conference on Research in Income and Wealth: Studies in Income and Wealth, Vol. 13.

Journal of the Royal Statistical Society. Series A (General), 115(2):293– 294, ???? 1952. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2981168.

Butler:1952:BRBb

[238] I. G. Butler. Book review: The Nature and Tax Treatment of Capital Gains and Losses, by Lawrence M. Seltzer. Journal of the Royal Statistical Society. Series A (General), 115(2):294– 295, ???? 1952. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2981169.

Brown:1952:BRB

[239] J. A. C. Brown. Book review: Coffee, Tea and Cocoa, by V. D. Wickizer. Journal of the Royal Statistical Society. Series A (General), 115(2):296, ???? 1952. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981170.

Anonymous:1952:SNb

[240] Anonymous. Statistical notes. Journal of the Royal Statistical Society. Series A (General), 115(2):297–299, ???? 1952. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981171.

Anonymous:1952:CNb

[241] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 115(2):299, ???? 1952. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981172.

Anonymous:1952:SEAb

[242] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society. Series A (General), 115(2):300–304, ???? 1952. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2981173.

Anonymous:1952:LALb

[243] Anonymous. List of additions to the library. Journal of the Royal Statistical Society. Series A (General), 115 (2):305–311, ???? 1952. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2981174.

Anonymous:1952:BMb

[244] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 115(2):??, ???? 1952. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2981175.

Anonymous:1952:FMc

[245] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 115(3):??, ???? 1952. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2980737.

Saunders:1952:CRM

[246] C. T. Saunders. Consumption of raw materials in the United Kingdom: 1851–1950. Journal of the Royal Statistical Society. Series A (General), 115 (3):313–353, ???? 1952. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2980738.

Adelstein:1952:APC

[247] A. M. Adelstein. Accident proneness: A criticism of the concept based upon an analysis of shunters’ accidents.

Journal of the Royal Statistical Society. Series A (General), 115(3):354– 410, ???? 1952. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2980739.

Moser:1952:QS

[248] C. A. Moser. Quota sampling. Journal of the Royal Statistical Society. Series A (General), 115(3):411–423, ???? 1952. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2980740.

Weatherburn:1952:SNS

[249] J. R. Weatherburn. The sources and nature of statistical information in special fields of statistics: Clothing statistics. Journal of the Royal Statistical Society. Series A (General), 115(3):424– 429, ???? 1952. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2980741.

Anonymous:1952:WP

[250] Anonymous. Wholesale prices in 1951. Journal of the Royal Statistical Society. Series A (General), 115(3):430– 434, ???? 1952. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2980742.

Lindley:1952:BRB

[251] D. V. Lindley. Book review: Logical Foundations of Probability, by Rudolf Carnap. Journal of the Royal Statistical Society. Series A (General), 115 (3):435–436, ???? 1952. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2980743.

Armitage:1952:BRBb

[252] P. Armitage. Book review: Probit Analysis, by D. J. Finney. Journal of the Royal Statistical Society. Series A (General), 115(3):436–437, ???? 1952. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2980744.

Healy:1952:BRB

[253] M. J. R. Healy. Book review: Statistical Methods for Chemists, by W. J. Youden. Journal of the Royal Statistical Society. Series A (General), 115 (3):437, ???? 1952. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2980745.

Wishart:1952:BRB

[254] J. Wishart. Book review: Elementary Principles of Statistics, by A. C. Rosander. Journal of the Royal Statistical Society. Series A (General), 115 (3):437–438, ???? 1952. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2980746.

Buckland:1952:BRBa

[255] W. R. Buckland. Book review: Statistics and their Application to Commerce, by A. Lester Boddington; A. R. Ilersic. Journal of the Royal Statistical Society. Series A (General), 115 (3):438–439, ???? 1952. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2980747.

Armitage:1952:BRBc

[256] P. Armitage. Book review: Tables to Facilitate Sequential t-Tests. Journal of the Royal Statistical Society. Series A (General), 115(3):439–440, ????

1952. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2980748.

Hartley:1952:BRBb

[257] H. O. Hartley. Book review: A Purification Method for Computing the Latent Columns of Numerical Matrices and Some Integrals of Differential Equations, by L. F. Richardson. Journal of the Royal Statistical Society. Series A (General), 115(3):440, ???? 1952. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2980749.

Rostas:1952:BRB

[258] L. Rostas. Book review: Some Aspects of the Industrial Structure of Scotland, by C. E. V. Leser. Journal of the Royal Statistical Society. Series A (General), 115(3):440–441, ???? 1952. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2980750.

Hartley:1952:BRBc

[259] H. O. Hartley. Book review: Punched Cards: Their Applications to Science and Industry, by R. S. Casey; J. W. Perry. Journal of the Royal Statistical Society. Series A (General), 115(3): 441, ???? 1952. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2980751.

Kemsley:1952:BRBc

[260] W. F. F. Kemsley. Book review: The Art of Asking Questions, by S. L. Payne. Journal of the Royal Statistical Society. Series A (General), 115 (3):442, ???? 1952. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2980752.

Nixon:1952:BRBa

[261] J. W. Nixon. Book review: Industrial Censuses in Western Countries: Report of a Group of Experts. Journal of the Royal Statistical Society. Series A (General), 115(3):442, ???? 1952. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2980753.

Nixon:1952:BRBb

[262] J. W. Nixon. Book review: United Nations, Department of Economic Affairs. Labour Productivity of the Cotton Textile Industry in Five Latin-American Countries. Journal of the Royal Statistical Society. Series A (General), 115(3):443, ???? 1952. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2980754.

Prais:1952:BRB

[263] S. J. Prais. Book review: Conference on Business Cycles, by Gottfried Haberler. Journal of the Royal Statistical Society. Series A (General), 115 (3):443–445, ???? 1952. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2980755.

Clarke:1952:BRB

[264] R. D. Clarke. Book review: The Economics of National Insurance, by Alan T. Peacock. Journal of the Royal Statistical Society. Series A (General), 115 (3):445–446, ???? 1952. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2980756.

Anonymous:1952:SNc

[265] Anonymous. Statistical notes. Journal of the Royal Statistical Society. Series A (General), 115(3):447–449, ???? 1952. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2980757.

Anonymous:1952:CNc

[266] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 115(3):450, ???? 1952. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2980758.

Anonymous:1952:SEAc

[267] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society. Series A (General), 115(3):451–453, ???? 1952. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2980759.

Anonymous:1952:LALc

[268] Anonymous. List of additions to the library. Journal of the Royal Statistical Society. Series A (General), 115 (3):454–461, ???? 1952. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2980760.

Anonymous:1952:BMc

[269] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 115(3):??, ???? 1952. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2980761.

Anonymous:1952:VI

[270] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (General), 115(4):617– 619, ???? 1952. CODEN JSSAEF.

ISSN 0035-9238. URL http://www.

jstor.org/stable/2980835.

Anonymous:1952:FMd

[271] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 115(4):??, ???? 1952. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2980836.

Ryan:1952:STC

[272] John Ryan. Statistics of tins and cans. Journal of the Royal Statistical Society. Series A (General), 115(4):463– 499, ???? 1952. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2980837.

Bowley:1952:INW

[273] A. L. Bowley. Index-numbers of wagerates and cost of living. Journal of the Royal Statistical Society. Series A (General), 115(4):500–506, ???? 1952. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2980838.

Roy:1952:EE

[274] A. D. Roy. An exercise in errors. Journal of the Royal Statistical Society. Series A (General), 115(4):507–520, ???? 1952. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2980839.

Lomax:1952:CEL

[275] K. S. Lomax. A criterion of efficiency in local administration. Journal of the Royal Statistical Society. Series A (General), 115(4):521–523, ???? 1952. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2980840.

Ross:1952:ESS

[276] N. S. Ross. Employment in shipbuilding and ship-repairing in Great Britain. Journal of the Royal Statistical Society. Series A (General), 115(4):524– 533, ???? 1952. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2980841.

King:1952:SNS

[277] A. L. King. The sources and nature of statistical information in special fields of statistics: Statistics relating to the petroleum industry, with particular reference to the United Kingdom. Journal of the Royal Statistical Society. Series A (General), 115(4):534–565, ???? 1952. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2980842.

Anonymous:1952:RC

[278] Anonymous. Report of the Council. Journal of the Royal Statistical Society. Series A (General), 115(4):566–578 + 580–583, ???? 1952. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2980843.

Anonymous:1952:POH

[279] Anonymous. Proceedings of the one hundred and eighteenth Annual General Meeting of the Royal Statistical Society, held at the London School of Hygiene and Tropical Medicine, on Thursday, June 26th, 1952. Journal of the Royal Statistical Society. Series A (General), 115(4):584, ???? 1952. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2980844.

Buckland:1952:BRBb

[280] W. R. Buckland. Book review: Proceedings of the Second Berkeley Symposium on Mathematical Statistics and Probability, by Jerzy Neyman. Journal of the Royal Statistical Society. Series A (General), 115(4):585, ???? 1952. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2980845.

Cox:1952:BRBb

[281] D. R. Cox. Book review: The Design and Analysis of Experiments, by O. Kempthorne. Journal of the Royal Statistical Society. Series A (General), 115(4):586–587, ???? 1952. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2980846.

Leser:1952:BRB

[282] C. E. V. Leser. Book review: Grundfragen der Ökonometrie, by W. Winkler. Journal of the Royal Statistical Society. Series A (General), 115(4): 587–588, ???? 1952. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2980847.

Morton:1952:BRB

[283] G. Morton. Book review: Social Choice and Individual Values, by Kenneth J. Arrow. Journal of the Royal Statistical Society. Series A (General), 115 (4):588, ???? 1952. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2980848.

Buckland:1952:BRBc

[284] W. R. Buckland. Book review: Quality Control and Industrial Statistics, by Acheson J. Duncan. Journal of the Royal Statistical Society. Series A (General), 115(4):588–589, ???? 1952. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2980849.

Nixon:1952:BRBc

[285] J. W. Nixon. Book review: Application of International Standards to Census Data of the Economically Active Population. Journal of the Royal Statistical Society. Series A (General), 115 (4):589, ???? 1952. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2980850.

Benjamin:1952:BRBb

[286] B. Benjamin. Book review: The Cost of Sickness and the Price of Health, by C. E. A. Winslow. Journal of the Royal Statistical Society. Series A (General), 115(4):589–591, ???? 1952. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2980851.

Benjamin:1952:BRBc

[287] B. Benjamin. Book review: Internal Migration, by Mary P. Newton; James R. Jeffery. Journal of the Royal Statistical Society. Series A (General), 115 (4):591–592, ???? 1952. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2980852.

Knowles:1952:BRB

[288] K. G. J. C. Knowles. Book review: Manpower Resources and Utilization, by A. J. Jaffe; Charles D. Stewart. Journal of the Royal Statistical Society. Series A (General), 115(4):592– 593, ???? 1952. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2980853.

Nicholson:1952:BRB

[289] J. L. Nicholson. Book review: The Urban Working-Class Household Diet 1940 to 1949. Journal of the Royal Statistical Society. Series A (General), 115 (4):593–595, ???? 1952. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2980854.

Benjamin:1952:BRBd

[290] B. Benjamin. Book review: Durée de la Vie Économiquement Active Suivant la Mortalité, by Giorgio Mortara. Journal of the Royal Statistical Society. Series A (General), 115(4):595, ???? 1952. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2980855.

Roy:1952:BRBb

[291] A. D. Roy. Book review: Studies in Income and Wealth, Vol. 14. Conference on Research in Income and Wealth. Journal of the Royal Statistical Society. Series A (General), 115(4):595– 596, ???? 1952. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2980856.

Rix:1952:BRBb

[292] M. S. Rix. Book review: The Sterling Area, by A. R. Conan. Journal of the Royal Statistical Society. Series A (General), 115(4):597, ???? 1952. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2980857.

Anonymous:1952:ONP

[293] Anonymous. Other new publications. Journal of the Royal Statistical Society. Series A (General), 115(4):597– 598, ???? 1952. CODEN JSSAEF.

ISSN 0035-9238. URL http://www.

jstor.org/stable/2980858.

Anonymous:1952:SNd

[294] Anonymous. Statistical notes. Journal of the Royal Statistical Society. Series A (General), 115(4):599–602, ???? 1952. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2980859.

Anonymous:1952:CNd

[295] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 115(4):602, ???? 1952. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2980860.

Anonymous:1952:SEAd

[296] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society. Series A (General), 115(4):603–607, ???? 1952. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2980861.

Anonymous:1952:LALd

[297] Anonymous. List of additions to the library. Journal of the Royal Statistical Society. Series A (General), 115 (4):608–611, ???? 1952. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2980862.

Anonymous:1952:PRb

[298] Anonymous. Periodical returns. Journal of the Royal Statistical Society. Series A (General), 115(4):612–615, ???? 1952. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2980863.

Anonymous:1952:BMd

[299] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 115(4):??, ???? 1952. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/4130702.

Anonymous:1953:FMa

[300] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 116(1):??, ???? 1953. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2980945.

Anonymous:1953:ES

[301] Anonymous. The expansion of statistics. Journal of the Royal Statistical Society. Series A (General), 116(1):1–10, ???? 1953. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2980946.

Kendall:1953:AET

[302] M. G. Kendall and A. Bradford Hill. The analysis of economic time-series-part i: Prices. Journal of the Royal Statistical Society. Series A (General), 116(1):11–34, ???? 1953. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2980947.

Frankel:1953:PGP

[303] H. Frankel. The price of gold and the purchasing power of the pound sterling. Journal of the Royal Statistical Society. Series A (General), 116(1):35–46, ???? 1953. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2980948.

Bellerby:1953:UKC

[304] J. R. Bellerby. United Kingdom: Comparison of urban and rural retail prices. Journal of the Royal Statistical Society. Series A (General), 116(1):47–56, ???? 1953. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2980949.

Jones:1953:SNS

[305] W. D. Stedman Jones. The sources and nature of statistical information in special fields of statistics: Food statistics. Journal of the Royal Statistical Society. Series A (General), 116(1):57–84, ???? 1953. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2980950.

Gaddum:1953:BRB

[306] J. H. Gaddum. Book review: Statistical Method in Biological Assay, by D. J. Finney. Journal of the Royal Statistical Society. Series A (General), 116(1):85– 86, ???? 1953. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2980951.

Lindley:1953:BRBa

[307] D. V. Lindley. Book review: Advanced Statistical Methods in Biometric Research, by C. Radhakrishna Rao. Journal of the Royal Statistical Society. Series A (General), 116(1):86–87, ???? 1953. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2980952.

Jowett:1953:BRB

[308] G. H. Jowett. Book review: Statistical Theory with Engineering Applications, by A. Hald; Statistical Tables and Formulas. by A. Hald. Journal of the Royal Statistical Society. Series A (General), 116(1):87–88, ???? 1953. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2980953.

Newman:1953:BRBa

[309] Dennis Newman. Book review: Technological Applications of Statistics, by L. H. C. Tippett. Journal of the Royal Statistical Society. Series A (General), 116(1):88–89, ???? 1953. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2980954.

Buckland:1953:BRBa

[310] W. R. Buckland. Book review: An Elementary Manual of Statistics, by Arthur Bowley. Journal of the Royal Statistical Society. Series A (General), 116(1):89–90, ???? 1953. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2980955.

Newman:1953:BRBb

[311] D. Newman. Book review: Elements of Statistics, by C. G. Lambe. Journal of the Royal Statistical Society. Series A (General), 116(1):90, ???? 1953. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2980956.

Sutherland:1953:BRB

[312] Ian Sutherland. Book review: Elementary Medical Statistics. The Principles of Quantitative Medicine, by Donald Mainland. Journal of the Royal Statistical Society. Series A (General), 116 (1):90–91, ???? 1953. CODEN JS-SAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2980957.

Wishart:1953:BRBa

[313] J. Wishart. Book review: Lectures and Conferences on Mathematical Statistics and Probability, by Jerzy Neyman. Journal of the Royal Statistical Society. Series A (General), 116(1):91–92, ???? 1953. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2980958.

Hartley:1953:BRBa

[314] H. O. Hartley. Book review: Numerical Analysis, by D. R. Hartree. Journal of the Royal Statistical Society. Series A (General), 116(1):92, ???? 1953. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2980959.

Gorman:1953:BRB

[315] W. M. Gorman. Book review: The Role of Measurement in Economics, by Richard Stone. Journal of the Royal Statistical Society. Series A (General), 116(1):92–93, ???? 1953. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2980960.

Daniell:1953:BRB

[316] P. A. Daniell. Book review: The Social Framework, by J. R. Hicks. Journal of the Royal Statistical Society. Series A (General), 116(1):93, ???? 1953. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2980961.

Roy:1953:BRB

[317] A. D. Roy. Book review: Conference on Research in Business Finance. Journal of the Royal Statistical Society. Series A (General), 116(1):94–95, ????

1953. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2980962.

Corlett:1953:BRB

[318] W. J. Corlett. Book review: The Pattern of United States Import Trade Since 1923: Some New Index Series and their Application, by John H. Alder; Eugene R. Schlesinger; Evelyn van Westerborg. Journal of the Royal Statistical Society. Series A (General), 116(1):95–97, ???? 1953. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2980963.

Cox:1953:BRB

[319] P. R. Cox. Book review: Approaches to Problems of High Fertility in Agrarian Societies. Journal of the Royal Statistical Society. Series A (General), 116 (1):97–98, ???? 1953. CODEN JS-SAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2980964.

Anonymous:1953:SNa

[320] Anonymous. Statistical notes. Journal of the Royal Statistical Society. Series A (General), 116(1):99–103, ???? 1953. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2980965.

Anonymous:1953:CNa

[321] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 116(1):103, ???? 1953. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2980966.

Bowley:1953:ADW

[322] A. L. Bowley. Augustus D. Webb 1880– 1953. Journal of the Royal Statistical Society. Series A (General), 116(1): 104, ???? 1953. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2980967.

Anonymous:1953:TT

[323] Anonymous. Theodore Taylor 1850– 1952. Journal of the Royal Statistical Society. Series A (General), 116(1): 104, ???? 1953. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2980968.

Anonymous:1953:SEAa

[324] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society. Series A (General), 116(1):105–110, ???? 1953. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2980969.

Anonymous:1953:LALa

[325] Anonymous. List of additions to the library. Journal of the Royal Statistical Society. Series A (General), 116 (1):111–118, ???? 1953. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2980970.

Anonymous:1953:PRa

[326] Anonymous. Periodical returns. Journal of the Royal Statistical Society. Series A (General), 116(1):119–121, ???? 1953. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2980971.

Anonymous:1953:BMa

[327] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 116(1):??, ????

1953. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2980972.

Anonymous:1953:FMb

[328] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 116(2):??, ???? 1953. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981056.

Deane:1953:RVU

[329] Phyllis Deane. Regional variations in United Kingdom incomes from employment, 1948. Journal of the Royal Statistical Society. Series A (General), 116 (2):123–139, ???? 1953. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981057.

Peaker:1953:SDU

[330] G. F. Peaker. A sampling design used by the Ministry of Education. Journal of the Royal Statistical Society. Series A (General), 116(2):140–165, ???? 1953. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981058.

Walker:1953:NRB

[331] A. M. Walker. A note on the relation between intelligence and parental income. Journal of the Royal Statistical Society. Series A (General), 116 (2):166–174, ???? 1953. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981059.

Burrowes:1953:PES

[332] W. D. Burrowes. The problem of exceptional samples in agricultural sample surveys. Journal of the Royal Statistical Society. Series A (General), 116 (2):175–176, ???? 1953. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981060.

Scott:1953:BAM

[333] A. D. Scott. Bibliography of applications of mathematical statistics to economics supplement for 1950. Journal of the Royal Statistical Society. Series A (General), 116(2):177–185, ???? 1953. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981061.

Anonymous:1953:WP

[334] Anonymous. Wholesale prices in 1952. Journal of the Royal Statistical Society. Series A (General), 116(2):186– 190, ???? 1953. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2981062.

Hazlewood:1953:SNS

[335] Arthur Hazlewood. The sources and nature of statistical information in special fields of statistics: Telecommunications statistics. Journal of the Royal Statistical Society. Series A (General), 116(2):191–202, ???? 1953. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2981063.

Wishart:1953:BRBb

[336] John Wishart. Book review: Notes on the Teaching of Statistics in Schools, by B. C. Brookes. Journal of the Royal Statistical Society. Series A (General), 116(2):203–204, ???? 1953. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2981064.

Owen:1953:BRB

[337] A. R. G. Owen. Book review: Methods of Statistical Analysis, by Cyril H. Goulden. Journal of the Royal Statistical Society. Series A (General), 116 (2):204–205, ???? 1953. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2981065.

Booker:1953:BRB

[338] H. S. Booker. Book review: Modern Elementary Statistics, by John E. Freund. Journal of the Royal Statistical Society. Series A (General), 116 (2):205–206, ???? 1953. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2981066.

Lindley:1953:BRBb

[339] D. V. Lindley. Book review: Introduction to the Theory of Games, by J. C. C. McKinsey. Journal of the Royal Statistical Society. Series A (General), 116(2):206, ???? 1953. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2981067.

Prest:1953:BRB

[340] A. R. Prest. Book review: Econometrics, by Gerhard Tintner. Journal of the Royal Statistical Society. Series A (General), 116(2):206–207, ???? 1953. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2981068.

Hartley:1953:BRBb

[341] H. O. Hartley. Book review: Mathematical Tables, Vol. X. Bessel Functions, Part II. Functions of Positive Integer Order. Journal of the Royal Statistical Society. Series A (General), 116 (2):207, ???? 1953. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2981069.

Hartley:1953:BRBc

[342] H. O. Hartley. Book review: Society of Actuaries: Report of Committee on New Recording Means and Computing Devices, September, 1952. Journal of the Royal Statistical Society. Series A (General), 116(2):208, ???? 1953. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981070.

Gorvin:1953:BRB

[343] J. H. Gorvin. Book review: Census of Woodlands, 1947–1949, by Forestry Commission. Journal of the Royal Statistical Society. Series A (General), 116 (2):208–209, ???? 1953. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981071.

Wilkins:1953:BRBa

[344] L. T. Wilkins. Book review: Government Information and the Research Worker: Lectures Delivered at a Vacation Course of the University of London School of Librarianship and Archives in April, 1951, by Ronald Staveley. Journal of the Royal Statistical Society. Series A (General), 116 (2):209, ???? 1953. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2981072.

Eddison:1953:BRB

[345] R. T. Eddison. Book review: Operations Research — A Preliminary Annotated Bibliography, by James H. Batchelor. Journal of the Royal Statistical Society. Series A (General), 116(2):

209–210, ???? 1953. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2981073.

Bowley:1953:BRB

[346] A. L. Bowley. Book review: Wages and Salaries in the United Kingdom, 1920– 1938, by Agatha L. Chapman; Rose Knight. Journal of the Royal Statistical Society. Series A (General), 116 (2):210–213, ???? 1953. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2981074.

Moser:1953:BRB

[347] C. A. Moser. Book review: Domestic Food Consumption and Expenditure, 1950. Journal of the Royal Statistical Society. Series A (General), 116(2): 213–214, ???? 1953. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2981075.

Seers:1953:BRBa

[348] Dudley Seers. Book review: Income and Wealth, Series I, by Erik Lundberg. Journal of the Royal Statistical Society. Series A (General), 116(2): 214–216, ???? 1953. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2981076.

Seers:1953:BRBb

[349] Dudley Seers. Book review: First Report of the National Income Committee, April 1951. Journal of the Royal Statistical Society. Series A (General), 116(2):216–217, ???? 1953. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2981077.

Benjamin:1953:BRB

[350] B. Benjamin. Book review: Problem Families: Five Inquiries, by C. P. Blacker. Journal of the Royal Statistical Society. Series A (General), 116 (2):217, ???? 1953. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2981078.

Anonymous:1953:SNb

[351] Anonymous. Statistical notes. Journal of the Royal Statistical Society. Series A (General), 116(2):218–221, ???? 1953. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981079.

Anonymous:1953:CNb

[352] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 116(2):222, ???? 1953. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981080.

Anonymous:1953:SEAb

[353] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society. Series A (General), 116(2):223–228, ???? 1953. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981081.

Anonymous:1953:LALb

[354] Anonymous. List of additions to the library. Journal of the Royal Statistical Society. Series A (General), 116 (2):229–236, ???? 1953. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2981082.

Anonymous:1953:BMb

[355] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 116(2):??, ???? 1953. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2981083.

Anonymous:1953:FMc

[356] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 116(3):??, ???? 1953. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342629.

Rix:1953:CIN

[357] M. S. Rix. Construction of an index number of gilt-edged values. Journal of the Royal Statistical Society. Series A (General), 116(3):237–255, ???? 1953. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342630.

Tippett:1953:SIL

[358] L. H. C. Tippett and P. D. Vincent. Statistical investigations of labour productivity in cotton spinning. Journal of the Royal Statistical Society. Series A (General), 116(3):256–271, ???? 1953. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342631.

Seers:1953:TE

[359] Dudley Seers. The treatment of errors. Journal of the Royal Statistical Society. Series A (General), 116(3):272– 274, ???? 1953. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342632.

Leak:1953:IWB

[360] H. Leak. The Industrial Wages Bill and earnings of operatives in 1935 and 1948. Journal of the Royal Statistical Society. Series A (General), 116(3): 275–282, ???? 1953. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342633.

Lodge:1953:SNS

[361] T. S. Lodge. The sources and nature of statistical information in special fields of statistics: Criminal statistics. Journal of the Royal Statistical Society. Series A (General), 116(3):283–297, ???? 1953. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342634.

MacGregor:1953:SNS

[362] J. J. MacGregor. The sources and nature of statistical information in special fields of statistics: Timber statistics. Journal of the Royal Statistical Society. Series A (General), 116(3):298– 322, ???? 1953. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342635.

Liddell:1953:BRB

[363] F. D. K. Liddell. Book review: Associated Measurements, by M. H. Quenouille. Journal of the Royal Statistical Society. Series A (General), 116 (3):323–324, ???? 1953. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342636.

Camm:1953:BRB

[364] G. L. Camm. Book review: Statistical Astronomy, by Robert J. Trumpler; Harold F. Weaver. Journal of the Royal Statistical Society. Series A (General), 116(3):324, ???? 1953. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342637.

Armitage:1953:BRBa

[365] P. Armitage. Book review: A Statistical Primer, by F. N. David. Journal of the Royal Statistical Society. Series A (General), 116(3):325, ???? 1953. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342638.

Quenouille:1953:BRB

[366] M. H. Quenouille. Book review: Introduction to Statistical Science in Agriculture, by D. J. Finney. Journal of the Royal Statistical Society. Series A (General), 116(3):325–326, ???? 1953. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342639.

Taylor:1953:BRB

[367] J. Taylor. Book review: Field Experimentation with Fruit Trees and Other Perennial Plants, by S. C. Pearce. Journal of the Royal Statistical Society. Series A (General), 116(3):326– 327, ???? 1953. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342640.

Tocher:1953:BRBa

[368] K. D. Tocher. Book review: Statistical Tables and Problems, by Albert E. Waugh. Journal of the Royal Statistical Society. Series A (General), 116 (3):327–329, ???? 1953. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342641.

Gearing:1953:BRBa

[369] Harold Gearing. Book review: Methods of Statistical Analysis in Economics and Business, by Edward E. Lewis. Journal of the Royal Statistical Society. Series A (General), 116(3):329– 330, ???? 1953. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342642.

Buckland:1953:BRBb

[370] W. R. Buckland. Book review: An Introduction to Statistical Calculations, by J. Mounsey. Journal of the Royal Statistical Society. Series A (General), 116(3):331, ???? 1953. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342643.

Wishart:1953:BRBc

[371] J. Wishart. Book review: Table of Arctan X. Journal of the Royal Statistical Society. Series A (General), 116 (3):332, ???? 1953. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342644.

Elphinstone:1953:BRB

[372] M. D. W. Elphinstone. Book review: The Practice of Life Assurance, by N. E. Coe; M. E. Ogborn. Journal of the Royal Statistical Society. Series A (General), 116(3):332, ???? 1953. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342645.

Anonymous:1953:SNc

[373] Anonymous. Statistical notes. Journal of the Royal Statistical Society. Series A (General), 116(3):333–334, ????

1953. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342646.

Anonymous:1953:CNc

[374] Anonymous. Current note. Journal of the Royal Statistical Society. Series A (General), 116(3):334, ???? 1953. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342647.

Anonymous:1953:SEAc

[375] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society. Series A (General), 116(3):335–339, ???? 1953. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342648.

Anonymous:1953:LALc

[376] Anonymous. List of additions to the library. Journal of the Royal Statistical Society. Series A (General), 116 (3):340–344, ???? 1953. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342649.

Anonymous:1953:PRb

[377] Anonymous. Periodical returns. Journal of the Royal Statistical Society. Series A (General), 116(3):345–348, ???? 1953. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342650.

Anonymous:1953:BMc

[378] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 116(3):??, ???? 1953. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342651.

Anonymous:1953:VI

[379] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (General), 116(4):480–iv, ???? 1953. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2343019.

Anonymous:1953:FMd

[380] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 116(4):??, ???? 1953. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343020.

Moser:1953:ESQ

[381] C. A. Moser and A. Stuart. An experimental study of quota sampling. Journal of the Royal Statistical Society. Series A (General), 116(4):349–405, ???? 1953. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343021.

Anonymous:1953:CCE

[382] Anonymous. A comment on a criterion of efficiency in local administration. Journal of the Royal Statistical Society. Series A (General), 116(4): 406–408, ???? 1953. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343022.

Lester:1953:SNS

[383] A. M. Lester. The sources and nature of statistical information in special fields of statistics: International air transport statistics. Journal of the Royal Statistical Society. Series A (General), 116 (4):409–423, ???? 1953. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343023.

Anonymous:1953:SNS

[384] Anonymous. The sources and nature of statistical information in special field of statistics: Statistics relating to the use of land in the United Kingdom. Journal of the Royal Statistical Society. Series A (General), 116(4):424–430, ???? 1953. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343024.

Anonymous:1953:RC

[385] Anonymous. Report of the Council. Journal of the Royal Statistical Society. Series A (General), 116(4):431– 450, ???? 1953. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2343025.

Anonymous:1953:POH

[386] Anonymous. Proceedings of the one hundred and nineteenth Annual General Meeting of the Royal Statistical Society, held at the London School of Hygiene and Tropical Medicine on Wednesday, June 24th, 1953. Journal of the Royal Statistical Society. Series A (General), 116(4):451, ???? 1953. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343026.

Anonymous:1953:RSC

[387] Anonymous. Regulations of the study circle on medical statistics. Journal of the Royal Statistical Society. Series A (General), 116(4):452, ???? 1953. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343027.

Healy:1953:BRB

[388] M. J. R. Healy. Book review: The Design and Analysis of Experiment, by M. H. Quenouille. Journal of the Royal Statistical Society. Series A (General), 116(4):453–454, ???? 1953. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2343028.

Lindley:1953:BRBc

[389] D. V. Lindley. Book review: Stochastic Processes, by J. L. Doob. Journal of the Royal Statistical Society. Series A (General), 116(4):454–456, ???? 1953. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343029.

Tocher:1953:BRBb

[390] K. D. Tocher. Book review: Numerical Solution of Differential Equations, by W. E. Milne. Journal of the Royal Statistical Society. Series A (General), 116(4):456–457, ???? 1953. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2343030.

Tocher:1953:BRBc

[391] K. D. Tocher. Book review: Linear Algebra and Matrix Theory, by Robert R. Stoll. Journal of the Royal Statistical Society. Series A (General), 116 (4):457–459, ???? 1953. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343031.

Armitage:1953:BRBb

[392] P. Armitage. Book review: The Statistics of Bioassay, with Special Reference to the Vitamins, by C. I. Bliss. Journal of the Royal Statistical Society. Series A (General), 116(4):459–460, ????

1953. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343032.

Kendall:1953:BRB

[393] M. G. Kendall. Book review: Handbook of Statistical Methods in Meteorology, by C. F. P. Brooks; N. Carruthers. Journal of the Royal Statistical Society. Series A (General), 116 (4):460–461, ???? 1953. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343033.

Johnson:1953:BRB

[394] N. L. Johnson. Book review: An Introduction to Statistics, by C. E. Clark. Journal of the Royal Statistical Society. Series A (General), 116(4):461– 462, ???? 1953. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343034.

Wilkins:1953:BRBb

[395] Leslie T. Wilkins. Book review: Factor Analysis: An Introduction and Manual for the Psychologist and Social Scientist, by Raymond B. Cattell. Journal of the Royal Statistical Society. Series A (General), 116(4):462–463, ???? 1953. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343035.

Wishart:1953:BRBd

[396] J. Wishart. Book review: Tables of Chebyshev Polynomials S_{n}(x) and C_{n}(x); Tables of the Bessel Functions Y_{0}(x), Y_{1}(x), K_{0}(x), K_{1}(x), 0 ≤ x ≤ 1; Tables of Bessel–Clifford Functions of Orders Zero and One. Journal of the Royal Statistical Society. Series A (General), 116(4):463–464, ????

1953. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343036.

Page:1953:BRB

[397] E. S. Page. Book review: Automatic Digital Calculators, by A. D. Booth; K. H. V. Booth. Journal of the Royal Statistical Society. Series A (General), 116(4):464, ???? 1953. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343037.

Cunliffe:1953:BRB

[398] S. V. Cunliffe. Book review: Statistical Methods for Chemical Experimentation, by W. L. Gore. Journal of the Royal Statistical Society. Series A (General), 116(4):464–465, ???? 1953. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343038.

Gearing:1953:BRBb

[399] Harold Gearing. Book review: Statistics as Applied to Accounting Data, by C. R. Curtis. Journal of the Royal Statistical Society. Series A (General), 116 (4):465, ???? 1953. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343039.

Berman:1953:BRB

[400] L. S. Berman. Book review: Shares of Upper Income Groups in Income and Savings, by Simon Kuznets; Elizabeth Jenks. Journal of the Royal Statistical Society. Series A (General), 116(4): 465–466, ???? 1953. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343040.

Devons:1953:BRB

[401] E. Devons. Book review: The Control of Raw Materials, by J. Hurstfield. Journal of the Royal Statistical Society. Series A (General), 116(4):466– 467, ???? 1953. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2343041.

Richmond:1953:BRB

[402] F. E. Richmond. Book review: Problems of Capital Formation in Underdeveloped Countries, by Ragnar Nurkse. Journal of the Royal Statistical Society. Series A (General), 116(4):467– 469, ???? 1953. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2343042.

Anonymous:1953:SNd

[403] Anonymous. Statistical notes. Journal of the Royal Statistical Society. Series A (General), 116(4):470–471, ???? 1953. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343043.

Anonymous:1953:CNd

[404] Anonymous. Current note. Journal of the Royal Statistical Society. Series A (General), 116(4):471, ???? 1953. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343044.

Anonymous:1953:SEAd

[405] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society. Series A (General), 116(4):472–475, ???? 1953. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343045.

Anonymous:1953:LALd

[406] Anonymous. List of additions to the library. Journal of the Royal Statistical Society. Series A (General), 116 (4):476–479, ???? 1953. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343046.

Anonymous:1953:BMd

[407] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 116(4):i–iv, ???? 1953. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343047.

Anonymous:1954:VIa

[408] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (General), 117(1):??, ???? 1954. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342765.

Anonymous:1954:FM

[409] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 117(1):??, ???? 1954. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342766.

Bauer:1954:SSM

[410] P. T. Bauer. Statistics of statutory marketing in West Africa, 1939–51. Journal of the Royal Statistical Society. Series A (General), 117(1):1–30, ???? 1954. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2342767.

Gabor:1954:EMT

[411] Denis Gabor and Andre Gabor. An essay on the mathematical theory of freedom. Journal of the Royal Statistical Society. Series A (General), 117(1):31– 72, ???? 1954. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2342768.

DaSilvaRodrigues:1954:LLM

[412] Milton Da Silva Rodrigues. Length of life of musicians. Journal of the Royal Statistical Society. Series A (General), 117(1):73–79, ???? 1954. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342769.

Anonymous:1954:ISI

[413] Anonymous. International Statistical Institute, 28th session, Rome, 1953. Journal of the Royal Statistical Society. Series A (General), 117(1):80–84, ???? 1954. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2342770.

Lomax:1954:SNS

[414] K. S. Lomax. The sources and nature of statistical information in special fields of statistics regional economic statistics. Journal of the Royal Statistical Society. Series A (General), 117 (1):85–101, ???? 1954. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342771.

Lindley:1954:BRB

[415] D. V. Lindley. Book review: Statistical Inference, by Helen M. Walker; Joseph Lev. Journal of the Royal Statistical Society. Series A (General), 117 (1):102–103, ???? 1954. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342772.

Vajda:1954:BRB

[416] S. Vajda. Book review: Contributions to the Theory of Games. Volume II, by H. W. Kuhn; A. W. Tucker. Journal of the Royal Statistical Society. Series A (General), 117(1):103–104, ???? 1954. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342773.

Bartlett:1954:BRB

[417] M. S. Bartlett. Book review: Communication Theory, by Willis Jackson. Journal of the Royal Statistical Society. Series A (General), 117(1):104– 105, ???? 1954. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342774.

Ehrenberg:1954:BRB

[418] A. S. C. Ehrenberg. Book review: Design and Analysis of Experiments in Psychology and Education, by E. F. Lindquist. Journal of the Royal Statistical Society. Series A (General), 117 (1):105–106, ???? 1954. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342775.

Tocher:1954:BRB

[419] K. D. Tocher. Book review: Faster Than Thought. A Symposium on Digital Computing Machines, by B. V. Bowden. Journal of the Royal Statistical Society. Series A (General), 117 (1):106–107, ???? 1954. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342776.

Booker:1954:BRB

[420] H. S. Booker. Book review: Charting Statistics, by Mary E. Spear. Journal of the Royal Statistical Society. Series A (General), 117(1):107, ???? 1954. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2342777.

Cox:1954:BRBa

[421] D. R. Cox. Book review: Statistics for Technologists, by C. G. Paradine; B. H. P. Rivett. Journal of the Royal Statistical Society. Series A (General), 117 (1):107–108, ???? 1954. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2342778.

VanRest:1954:BRB

[422] E. D. Van Rest. Book review: Techniques Modernes de Controle des Fabrications, by J. Mothes. Journal of the Royal Statistical Society. Series A (General), 117(1):108–109, ???? 1954. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2342779.

Wishart:1954:BRBa

[423] J. Wishart. Book review: Tables of Normal Probability Functions; Probability Tables for the Analysis of Extreme-Value Data. Journal of the Royal Statistical Society. Series A (General), 117(1):109–110, ???? 1954. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2342780.

Wishart:1954:BRBb

[424] J. Wishart. Book review: Cambridge Elementary Statistical Tables, by D.

V. Lindley; J. C. P. Miller. Journal of the Royal Statistical Society. Series A (General), 117(1):110, ???? 1954. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342781.

Clark:1954:BRBa

[425] Colin Clark. Book review: The Economic Impact on Under-Developed Societies, by S. H. Frankel. Journal of the Royal Statistical Society. Series A (General), 117(1):111, ???? 1954. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342782.

Rostas:1954:BRB

[426] L. Rostas. Book review: Productivity and Economic Progress, by Frederick C. Mills. Journal of the Royal Statistical Society. Series A (General), 117 (1):112, ???? 1954. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342783.

Clark:1954:BRBb

[427] Colin Clark. Book review: World Population and Production, by W. S. Woytinsky; E. S. Woytinsky. Journal of the Royal Statistical Society. Series A (General), 117(1):112–113, ???? 1954. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342784.

Cox:1954:BRBb

[428] P. R. Cox. Book review: Typenlehre der Demographie, by Wilhelm Winkler. Journal of the Royal Statistical Society. Series A (General), 117(1):114, ???? 1954. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342785.

Ridley:1954:BRB

[429] T. M. Ridley. Book review: Income and Wealth, Series II, by Simon Kuznets; Income and Wealth, Series III., by Milton Gilbert. Journal of the Royal Statistical Society. Series A (General), 117 (1):114–115, ???? 1954. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2342786.

Saunders:1954:BRBa

[430] C. T. Saunders. Book review: The Sterling Area: An American Analysis, by John M. Cassels. Journal of the Royal Statistical Society. Series A (General), 117(1):116–117, ???? 1954. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2342787.

Gorvin:1954:BRB

[431] J. H. Gorvin. Book review: Standard of Living in India and Pakistan, 1931–32 to 1940–41, by R. C. Desai. Journal of the Royal Statistical Society. Series A (General), 117(1):117, ???? 1954. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2342788.

Anonymous:1954:SNa

[432] Anonymous. Statistical notes. Journal of the Royal Statistical Society. Series A (General), 117(1):118–119, ???? 1954. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2342789.

Anonymous:1954:CNa

[433] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 117(1):120, ????

1954. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342790.

Sutherland:1954:WR

[434] Ian Sutherland. W. T. Russell. Journal of the Royal Statistical Society. Series A (General), 117(1):121, ???? 1954. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342791.

Anonymous:1954:SEAa

[435] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society. Series A (General), 117(1):122–126, ???? 1954. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342792.

Anonymous:1954:LALa

[436] Anonymous. List of additions to the library. Journal of the Royal Statistical Society. Series A (General), 117 (1):127–131, ???? 1954. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342793.

Browning:1954:CCG

[437] H. E. Browning and A. A. Sorrell. Cinemas and cinema-going in Great Britain. Journal of the Royal Statistical Society. Series A (General), 117(2):133– 170, ???? 1954. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343336.

Farrell:1954:DMC

[438] M. J. Farrell. The demand for motorcars in the United States. Journal of the Royal Statistical Society. Series A (General), 117(2):171–201, ????

1954. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343337.

Read:1954:POO

[439] D. R. Read, P. G. Smith, and J. H. Ince. Pairing of observations in ordered samples from: A normal population. Journal of the Royal Statistical Society. Series A (General), 117 (2):202–204, ???? 1954. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343338.

Anonymous:1954:WP

[440] Anonymous. Wholesale prices in 1953. Journal of the Royal Statistical Society. Series A (General), 117(2):205– 209, ???? 1954. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343339.

Benjamin:1954:NSS

[441] B. Benjamin. The nature and sources of statistics in special fields of statistics: Medicine and medical services in the United Kingdom. Journal of the Royal Statistical Society. Series A (General), 117(2):210–232, ???? 1954. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2343340.

Anscombe:1954:BRB

[442] F. J. Anscombe. Book review: Scientific Explanation: A Study of the Function of Theory, Probability and Law in Science, by R. B. Braithwaite. Journal of the Royal Statistical Society. Series A (General), 117(2):233–234, ???? 1954. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343341.

Corlett:1954:BRB

[443] W. J. Corlett. Book review: Demand Analysis: A Study in Econometrics, by Herman Wold; Lars Jureen. Journal of the Royal Statistical Society. Series A (General), 117(2):235, ???? 1954. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343342.

Morton:1954:BRB

[444] G. Morton. Book review: An Introduction to Linear Programming, by A. Charnes; W. W. Cooper; A. Henderson. Journal of the Royal Statistical Society. Series A (General), 117(2):236, ???? 1954. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2343343.

Gray:1954:BRB

[445] P. G. Gray. Book review: Sample Survey Methods and Theory, by M. Hansen; W. Hurwitz; W. Madow. Journal of the Royal Statistical Society. Series A (General), 117(2):236–237, ???? 1954. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343344.

Wishart:1954:BRBc

[446] J. Wishart. Book review: Exercises in Theoretical Statistics, by M. G. Kendall. Journal of the Royal Statistical Society. Series A (General), 117 (2):237, ???? 1954. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343345.

Wishart:1954:BRBd

[447] J. Wishart. Book review: 50–100 Binomial Tables, by Harry G. Romig. Journal of the Royal Statistical Society. Series A (General), 117(2):237–238, ???? 1954. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343346.

Wishart:1954:BRBe

[448] J. Wishart. Book review: Tables of 10^{x} (Antilogarithms to the Base 10); Table of Natural Logarithms for Arguments Between Zero and Five to Sixteen Decimal Places. Journal of the Royal Statistical Society. Series A (General), 117 (2):238, ???? 1954. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343347.

Nixon:1954:BRBa

[449] J. W. Nixon. Book review: External Migration: A Study of the Available Statistics, 1815–1950, by N. H. Carrier; J. R. Jeffery. Journal of the Royal Statistical Society. Series A (General), 117 (2):238–239, ???? 1954. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343348.

Lomax:1954:BRB

[450] K. S. Lomax. Book review: Social Accounts of Local Authorities: An Experiment in Sampling Methods, by J. E. G. Utting. Journal of the Royal Statistical Society. Series A (General), 117 (2):239, ???? 1954. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343349.

Maizels:1954:BRB

[451] A. Maizels. Book review: British Overseas Trade from 1700 to the 1930’s, by Werner Schlote; W. O. Henderson; W. H. Chaloner. Journal of the Royal Statistical Society. Series A (General), 117 (2):240–242, ???? 1954. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343350.

Glenday:1954:BRBa

[452] R. Glenday. Book review: A Study in Trade Cycle History, 1833–42, by R. C. O. Matthews. Journal of the Royal Statistical Society. Series A (General), 117(2):242, ???? 1954. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343351.

George:1954:BRB

[453] Percy George. Book review: Welsh Economic Statistics: A Handbook of Sources, by Anne Martin; J. Parry Lewis. Journal of the Royal Statistical Society. Series A (General), 117(2): 243, ???? 1954. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343352.

Wilkins:1954:BRB

[454] Leslie T. Wilkins. Book review: Bristol on the Move. Journal of the Royal Statistical Society. Series A (General), 117(2):243–244, ???? 1954. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2343353.

Anonymous:1954:SNb

[455] Anonymous. Statistical notes. Journal of the Royal Statistical Society. Series A (General), 117(2):245–248, ???? 1954. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343354.

Anonymous:1954:CNb

[456] Anonymous. Current note. Journal of the Royal Statistical Society. Series A (General), 117(2):248, ????

1954. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343355.

Anonymous:1954:SEAb

[457] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society. Series A (General), 117(2):249–252, ???? 1954. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343356.

Anonymous:1954:LALb

[458] Anonymous. List of additions to the library. Journal of the Royal Statistical Society. Series A (General), 117 (2):253–259, ???? 1954. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2343357.

Anonymous:1954:PRa

[459] Anonymous. Periodical returns. Journal of the Royal Statistical Society. Series A (General), 117(2):260–261, ???? 1954. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343358.

Clark:1954:WSR

[460] Colin Clark. World supply and requirements of farm products. Journal of the Royal Statistical Society. Series A (General), 117(3):263–296, ???? 1954. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343382.

Glover:1954:ORG

[461] K. F. Glover and D. N. Miller. The outlines of the road goods transport industry. Journal of the Royal Statistical Society. Series A (General), 117 (3):297–330, ???? 1954. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343383.

Rosenbaum:1954:HWA

[462] S. Rosenbaum. Heights and weights of the Army intake, 1951. Journal of the Royal Statistical Society. Series A (General), 117(3):331–347, ???? 1954. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343384.

Lee:1954:SNS

[463] Doris M. Lee. The sources and nature of statistical information in special fields of statistics: The statistics of education. Journal of the Royal Statistical Society. Series A (General), 117 (3):348–360, ???? 1954. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343385.

Cox:1954:BRBc

[464] D. R. Cox. Book review: Biomathematics. The Principles of Mathematics for Students of Biological Science, by C. A. B. Smith. Journal of the Royal Statistical Society. Series A (General), 117(3):361–362, ???? 1954. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2343386.

Wishart:1954:BRBf

[465] J. Wishart. Book review: Tests of Significance in a 2 × 2 Contingency Table: Extension of Finney’s Table, by R. Latscha; Tables of Symmetric Functions. Part IV, by F. N. David; M. G. Kendall; Bessel Functions and Formulae, by W. G. Bickley. Journal of the Royal Statistical Society. Series A (General), 117(3):362, ????

1954. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343387.

Liddell:1954:BRB

[466] F. D. K. Liddell. Book review: Statistical Methods in Electrical Engineering, by D. A. Bell. Journal of the Royal Statistical Society. Series A (General), 117(3):362–363, ???? 1954. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2343388.

Cox:1954:BRBd

[467] P. R. Cox. Book review: Population Statistics and their Compilation, by Hugh H. Wolfenden. Journal of the Royal Statistical Society. Series A (General), 117(3):363–364, ???? 1954. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343389.

Nicholson:1954:BRB

[468] J. L. Nicholson. Book review: A Study of Aggregate Consumption Functions, by R. Ferber. Journal of the Royal Statistical Society. Series A (General), 117 (3):364–365, ???? 1954. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2343390.

Deane:1954:BRB

[469] P. M. Deane. Book review: An International Comparison of National Products and the Purchasing Power of Currencies, by Milton Gilbert; Irving B. Kravis. Journal of the Royal Statistical Society. Series A (General), 117 (3):365–366, ???? 1954. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2343391.

Glenday:1954:BRBb

[470] R. Glenday. Book review: The Frontiers of Economic Knowledge, by Arthur F. Burns. Journal of the Royal Statistical Society. Series A (General), 117(3):366–367, ???? 1954. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2343392.

Anonymous:1954:SNc

[471] Anonymous. Statistical notes. Journal of the Royal Statistical Society. Series A (General), 117(3):368–370, ???? 1954. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343393.

Anonymous:1954:CNc

[472] Anonymous. Current note. Journal of the Royal Statistical Society. Series A (General), 117(3):370, ???? 1954. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343394.

Anonymous:1954:SEAc

[473] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society. Series A (General), 117(3):371–375, ???? 1954. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343395.

Anonymous:1954:LALc

[474] Anonymous. List of additions to the library. Journal of the Royal Statistical Society. Series A (General), 117 (3):376–381, ???? 1954. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343396.

Anonymous:1954:PRb

[475] Anonymous. Periodical returns. Journal of the Royal Statistical Society. Series A (General), 117(3):382–385, ???? 1954. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343397.

Anonymous:1954:VIb

[476] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (General), 117(4):513– 516, ???? 1954. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2342678.

Durbin:1954:CCS

[477] J. Durbin and A. Stuart. Callbacks and clustering in sample surveys: An experimental study. Journal of the Royal Statistical Society. Series A (General), 117(4):387–428, ???? 1954. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2342679.

Silcock:1954:PLT

[478] H. Silcock. The phenomenon of labour turnover. Journal of the Royal Statistical Society. Series A (General), 117 (4):429–440, ???? 1954. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2342680.

Karmel:1954:RBC

[479] P. H. Karmel. The relations between chained indexes of input, Gross output and net output. Journal of the Royal Statistical Society. Series A (General), 117(4):441–458, ???? 1954. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2342681.

Finucan:1954:MGD

[480] H. M. Finucan. The mode of a grouped distribution. Journal of the Royal Statistical Society. Series A (General), 117 (4):459, ???? 1954. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342682.

L:1954:FIC

[481] W. P. D. L. First International Conference of National Committees on Vital and Health Statistics. Journal of the Royal Statistical Society. Series A (General), 117(4):460–462, ???? 1954. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342683.

Anonymous:1954:RC

[482] Anonymous. Report of the Council. Journal of the Royal Statistical Society. Series A (General), 117(4):463– 480, ???? 1954. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342684.

Anonymous:1954:POH

[483] Anonymous. Proceedings of the one hundred and Twentieth Annual General Meeting of the Royal Statistical Society, held at the London School of Hygiene and Tropical Medicine on Wednesday, June 23rd, 1954. Journal of the Royal Statistical Society. Series A (General), 117(4):481, ???? 1954. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342685.

Kendall:1954:BRBa

[484] Maurice Kendall. Book review: Uppsala Symposium on Psychological Factor Analysis, 17–19 March, 1953. Journal of the Royal Statistical Society. Series A (General), 117(4):482–483, ???? 1954. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342686.

Kendall:1954:BRBb

[485] Maurice Kendall. Book review: A Study in the Analysis of Stationary Time-Series, by Herman Wold. Journal of the Royal Statistical Society. Series A (General), 117(4):484, ???? 1954. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342687.

Wishart:1954:BRBg

[486] J. Wishart. Book review: Sampling Theory of Surveys with Applications, by Pandurang V. Sukhatme. Journal of the Royal Statistical Society. Series A (General), 117(4):484–485, ???? 1954. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342688.

Brookes:1954:BRB

[487] B. C. Brookes. Book review: Handbook of Graphic Presentation, by Calvin F. Schmid. Journal of the Royal Statistical Society. Series A (General), 117 (4):485, ???? 1954. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342689.

Cox:1954:BRBe

[488] D. R. Cox. Book review: Elements of Statistics, by H. C. Fryer. Journal of the Royal Statistical Society. Series A (General), 117(4):486, ???? 1954. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342690.

Glover:1954:BRB

[489] K. F. Glover. Book review: Transport and the State of Trade in Britain, by Thos. Hultgren; William I. Greenwald. Journal of the Royal Statistical Society. Series A (General), 117(4): 486–487, ???? 1954. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342691.

Mason:1954:BRB

[490] J. I. Mason. Book review: Advertising Expenditure 1952, by Rodney Silverman. Journal of the Royal Statistical Society. Series A (General), 117 (4):488–489, ???? 1954. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342692.

Copland:1954:BRB

[491] B. D. Copland. Book review: Marketing Research and Management, by A. G. Irvine. Journal of the Royal Statistical Society. Series A (General), 117 (4):489–490, ???? 1954. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342693.

Douglas:1954:BRB

[492] J. I. Douglas. Book review: The Scottish Economy: A Statistical Account of Scottish Life, by Members of the Staff of Glasgow University, by A. K. Cairncross. Journal of the Royal Statistical Society. Series A (General), 117(4): 490–491, ???? 1954. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342694.

Nixon:1954:BRBb

[493] J. W. Nixon. Book review: Migration and Economic Growth: A Study of Great Britain and the Atlantic Economy, by Brinley Thomas. Journal of the Royal Statistical Society. Series A (General), 117(4):491–492, ???? 1954. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342695.

Roy:1954:BRB

[494] A. D. Roy. Book review: Economic Change: Selected Essays in Business Cycles, National Income, and Economic Growth, by Simon Kuznets. Journal of the Royal Statistical Society. Series A (General), 117(4):492– 493, ???? 1954. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342696.

Nixon:1954:BRBc

[495] J. W. Nixon. Book review: International Definition and Measurement of Standards and Levels of Living. Journal of the Royal Statistical Society. Series A (General), 117(4):493–494, ???? 1954. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342697.

Saunders:1954:BRBb

[496] C. T. Saunders. Book review: The Trend of Government Activity in the United States since 1900, by Solomon Fabricant; Robert E. Lipsey. Journal of the Royal Statistical Society. Series A (General), 117(4):494–495, ???? 1954. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342698.

Saunder:1954:BRB

[497] C. T. Saunder. Book review: International Insurance Transactions: Insurances in the Balances-of-Payments, by S. J. Lengyel. Journal of the Royal Statistical Society. Series A (General), 117 (4):495–496, ???? 1954. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342699.

Irwin:1954:BRB

[498] J. O. Irwin. Book review: How to Lie with Statistics, by Darrell Huff. Journal of the Royal Statistical Society. Series A (General), 117(4):496, ???? 1954. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342700.

Anonymous:1954:SNd

[499] Anonymous. Statistical notes. Journal of the Royal Statistical Society. Series A (General), 117(4):497–499, ???? 1954. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342701.

Anonymous:1954:CNd

[500] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 117(4):499, ???? 1954. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342702.

Anonymous:1954:SHC

[501] Anonymous. Sir Henry Clay. Journal of the Royal Statistical Society. Series A (General), 117(4):500–501, ???? 1954. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342703.

Anonymous:1954:MS

[502] Anonymous. A. M. Southall. Journal of the Royal Statistical Society. Series A (General), 117(4):501, ????

1954. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342704.

Anonymous:1954:SEAd

[503] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society. Series A (General), 117(4):502–505, ???? 1954. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342705.

Anonymous:1954:LALd

[504] Anonymous. List of additions to the library. Journal of the Royal Statistical Society. Series A (General), 117 (4):506–511, ???? 1954. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342706.

Anonymous:1954:PRc

[505] Anonymous. Periodical returns. Journal of the Royal Statistical Society. Series A (General), 117(4):512, ???? 1954. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342707.

Anonymous:1955:VIa

[506] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (General), 118(1):??, ???? 1955. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342515.

Anonymous:1955:FM

[507] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 118(1):??, ???? 1955. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342516.

Piercy:1955:MGS

[508] Lord Piercy. The MacMillan gap and the shortage of risk capital. Journal of the Royal Statistical Society. Series A (General), 118(1):1–7, ???? 1955. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342517.

Anonymous:1955:PM

[509] Anonymous. Proceedings of the meeting. Journal of the Royal Statistical Society. Series A (General), 118(1):7–11, ???? 1955. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2342518.

Benjamin:1955:CPP

[510] B. Benjamin, E. Grebenik, W. P. D. Logan, Brinley Thomas, J. Hajnal, and J. A. Fraser Roberts. Current population problems: With particular reference to the United Nations World Population Conference. Journal of the Royal Statistical Society. Series A (General), 118(1):12–27, ???? 1955. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342519.

Turner:1955:MU

[511] H. A. Turner. Measuring unemployment. Journal of the Royal Statistical Society. Series A (General), 118(1):28– 50, ???? 1955. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2342520.

Khandker:1955:PDE

[512] R. H. Khandker. The pattern of a divided economy: A national income analysis of India and Pakistan. Journal of the Royal Statistical Society. Series A (General), 118(1):51–55, ???? 1955. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342521.

Prais:1955:MSM

[513] S. J. Prais. Measuring social mobility. Journal of the Royal Statistical Society. Series A (General), 118(1):56–66, ???? 1955. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2342522.

Mason:1955:SNS

[514] J. I. Mason and D. Sachs. The sources and nature of statistical information in special fields of statistics: Statistics bearing on financial investment. Journal of the Royal Statistical Society. Series A (General), 118(1):67–100, ???? 1955. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342523.

Cox:1955:BRBa

[515] D. R. Cox. Book review: Design and Analysis of Industrial Experiments, by O. L. Davies. Journal of the Royal Statistical Society. Series A (General), 118 (1):101–102, ???? 1955. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342524.

Vajda:1955:BRBa

[516] S. Vajda. Book review: Theory of Games and Statistical Decisions, by David Blackwell; M. A. Girshick. Journal of the Royal Statistical Society. Series A (General), 118(1):102–103, ???? 1955. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342525.

Cox:1955:BRBb

[517] D. R. Cox. Book review: Introduction to Mathematical Statistics, by P. G. Hoel. Journal of the Royal Statistical Society. Series A (General), 118 (1):103, ???? 1955. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342526.

Welch:1955:BRB

[518] B. L. Welch. Book review: Biometrika Tables for Statisticians, Vol. I, by E. S. Pearson; H. O. Hartley. Journal of the Royal Statistical Society. Series A (General), 118(1):103–105, ???? 1955. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342527.

Tocher:1955:BRBa

[519] K. D. Tocher. Book review: Tables of Langrangian Coefficients for Sexagesimal Interpolation; Tables of Secants and Cosecants to Nine Significant Figures at Hundredths of a Degree. Journal of the Royal Statistical Society. Series A (General), 118(1):105–106, ???? 1955. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342528.

Tocher:1955:BRBb

[520] K. D. Tocher. Book review: Statistical Theory of Extreme Values and Some Practical Applications, by E. J. Gumbel; Probability Tables for the Analysis of Extreme Value Data. Journal of the Royal Statistical Society. Series A (General), 118(1):106, ???? 1955. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342529.

Tocher:1955:BRBc

[521] K. D. Tocher. Book review: A Short Table for Bessel Functions of Integer Orders and Large Arguments, by L. Fox; Table of Binomial Coefficients, by J. C. P. Miller. Journal of the Royal Statistical Society. Series A (General), 118(1):106–107, ???? 1955. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2342530.

Armitage:1955:BRBa

[522] P. Armitage. Book review: An Outline of Biometry, by C. I. Bliss; D. W. Calhoun. Journal of the Royal Statistical Society. Series A (General), 118 (1):107–108, ???? 1955. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2342531.

Chambers:1955:BRB

[523] E. G. Chambers. Book review: The Psychology of Politics, by H. J. Eysenck. Journal of the Royal Statistical Society. Series A (General), 118 (1):108–109, ???? 1955. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2342532.

Swan:1955:BRB

[524] A. W. Swan. Book review: Quality Control, by H. L. Edrick. Journal of the Royal Statistical Society. Series A (General), 118(1):109–110, ???? 1955. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2342533.

Vajda:1955:BRBb

[525] S. Vajda. Book review: Introduction a los Metodos de la Estadistica.Part 2, by Sixto Rios. Journal of the Royal Statistical Society. Series A (General), 118 (1):110–111, ???? 1955. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342534.

Lomax:1955:BRB

[526] K. S. Lomax. Book review: British Economic Statistics, by C. F. Carter; A. D. Roy. Journal of the Royal Statistical Society. Series A (General), 118 (1):111–112, ???? 1955. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342535.

Goldie:1955:BRB

[527] N. Goldie. Book review: Proceedings of the Toronto Meterological Conference 1953. Journal of the Royal Statistical Society. Series A (General), 118 (1):112–113, ???? 1955. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342536.

Daniell:1955:BRB

[528] P. A. Daniell. Book review: The Economics of Recession and Revival, by K. D. Roose. Journal of the Royal Statistical Society. Series A (General), 118 (1):113, ???? 1955. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342537.

Ady:1955:BRB

[529] P. Ady. Book review: The International Flow of Private Capital 1946– 52. Journal of the Royal Statistical Society. Series A (General), 118(1):113– 114, ???? 1955. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342538.

Bowley:1955:BRB

[530] A. L. Bowley. Book review: The Measurement of Consumers’ Expenditure and Behaviour in the United Kingdom, 1920–1938. Vol. I., by Richard Stone; D. A. Rowe; W. J. Corlett; Consumers’ Expenditure in the United Kingdom, 1900–1919, by A. R. Prest; A. A. Adams. Journal of the Royal Statistical Society. Series A (General), 118 (1):114–116, ???? 1955. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342539.

Menzler:1955:BRB

[531] F. A. A. Menzler. Book review: An Introduction to the Study of Industrial Relations, by J. Henry Richardson. Journal of the Royal Statistical Society. Series A (General), 118(1):116– 118, ???? 1955. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342540.

Gorvin:1955:BRB

[532] J. H. Gorvin. Book review: Wartime Agriculture in Australia and New Zealand, 1939–50, by J. G. Crawford. Journal of the Royal Statistical Society. Series A (General), 118(1):118, ???? 1955. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342541.

Ramm:1955:BRB

[533] B. T. Ramm. Book review: Retail Trading in Britain 1850–1950, by James B. Jefferys. Journal of the Royal Statistical Society. Series A (General), 118(1):118–120, ???? 1955. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2342542.

Anonymous:1955:SNa

[534] Anonymous. Statistical notes. Journal of the Royal Statistical Society. Series A (General), 118(1):121–123, ???? 1955. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342543.

Anonymous:1955:CNa

[535] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 118(1):123, ???? 1955. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342544.

Reddaway:1955:DR

[536] W. B. Reddaway. Dr. L. Rostas. Journal of the Royal Statistical Society. Series A (General), 118(1):124–125, ???? 1955. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342545.

Bowley:1955:BSR

[537] A. L. Bowley. B. Seebohm Rowntree, C.H. 1871–1954. Journal of the Royal Statistical Society. Series A (General), 118(1):125–126, ???? 1955. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2342546.

Anonymous:1955:SEAa

[538] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society. Series A (General), 118(1):127–130, ???? 1955. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342547.

Anonymous:1955:LALa

[539] Anonymous. List of additions to the library. Journal of the Royal Statistical Society. Series A (General), 118 (1):131–137, ???? 1955. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342548.

Anonymous:1955:PRa

[540] Anonymous. Periodical returns. Journal of the Royal Statistical Society. Series A (General), 118(1):138–139, ???? 1955. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342549.

Redfern:1955:NIF

[541] Philip Redfern. Net investment in fixed assets in the United Kingdom, 1938– 1953. Journal of the Royal Statistical Society. Series A (General), 118(2): 141–192, ???? 1955. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343128.

Anonymous:1955:DTM

[542] Anonymous. A discussion on the teaching of mathematical statistics at university level. Journal of the Royal Statistical Society. Series A (General), 118 (2):193–223, ???? 1955. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343129.

Houston:1955:FWC

[543] George Houston. Farm wages in central Scotland from 1814 to 1870. Journal of the Royal Statistical Society. Series A (General), 118(2):224–228, ???? 1955. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343130.

Anonymous:1955:WP

[544] Anonymous. Wholesale prices in 1954. Journal of the Royal Statistical Society. Series A (General), 118(2):229– 233, ???? 1955. CODEN JSSAEF.

ISSN 0035-9238. URL http://www.

jstor.org/stable/2343131.

Brocks:1955:SNS

[545] B. J. Brocks. The sources and nature of statistical information in special fields of statistics: Statistics available on the local government services. Journal of the Royal Statistical Society. Series A (General), 118(2):234–244, ???? 1955. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343132.

Good:1955:BRB

[546] I. J. Good. Book review: The Foundations of Statistics, by L. J. Savage. Journal of the Royal Statistical Society. Series A (General), 118(2):245– 246, ???? 1955. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343133.

Booker:1955:BRB

[547] H. S. Booker. Book review: Research Methods in the Behavioral Sciences, by Leon Festinger; Daniel Katz. Journal of the Royal Statistical Society. Series A (General), 118(2):246–247, ???? 1955. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343134.

Sampford:1955:BRB

[548] M. R. Sampford. Book review: Statistics and Mathematics in Biology, by Oscar Kempthorne; T. A. Bancroft; J. W. Owen; J. L. Lush. Journal of the Royal Statistical Society. Series A (General), 118(2):247–248, ???? 1955. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343135.

Lindley:1955:BRBa

[549] D. V. Lindley. Book review: Limit Distributions for Sums of Independent Random Variables, by B. V. Gnedenko; A. N. Kolmogorov; K. L. Chung. Journal of the Royal Statistical Society. Series A (General), 118(2):248–249, ???? 1955. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343136.

Gabor:1955:BRB

[550] Andre Gabor. Book review: Numerical Methods, by Andrew D. Booth. Journal of the Royal Statistical Society. Series A (General), 118(2):249–250, ???? 1955. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343137.

Farrell:1955:BRB

[551] M. J. Farrell. Book review: The Multiplier Theory, by H. Hegeland. Journal of the Royal Statistical Society. Series A (General), 118(2):250, ???? 1955. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343138.

Hammersley:1955:BRBa

[552] J. M. Hammersley. Book review: The Compleat Strategyst, by J. D. Williams. Journal of the Royal Statistical Society. Series A (General), 118(2):250– 251, ???? 1955. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2343139.

Cox:1955:BRBc

[553] P. R. Cox. Book review: The Trend and Pattern of Fertility in Great Britain: A Report on the Family Census of 1946, by D. V. Glass; E. Grebenik. Journal of the Royal Statistical Society. Series A (General), 118(2):251–252, ???? 1955. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343140.

Adams:1955:BRB

[554] A. A. Adams. Book review: Statistics of National Product and Expenditure, 1938, 1947 and 1952. Journal of the Royal Statistical Society. Series A (General), 118(2):252–253, ???? 1955. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343141.

Nixon:1955:BRBa

[555] J. W. Nixon. Book review: British Post-War Migration, by Julius Isaac. Journal of the Royal Statistical Society. Series A (General), 118(2):253– 255, ???? 1955. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343142.

Nixon:1955:BRBb

[556] J. W. Nixon. Book review: National Incomes and International Trade: A Quantitative Analysis, by Hans Neisser; Franco Modigliani. Journal of the Royal Statistical Society. Series A (General), 118(2):255–256, ???? 1955. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343143.

Glenday:1955:BRBa

[557] R. Glenday. Book review: A History of Economic Thought, by Eric Roll. Journal of the Royal Statistical Society. Series A (General), 118(2):256–257, ???? 1955. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343144.

Anonymous:1955:SNb

[558] Anonymous. Statistical notes. Journal of the Royal Statistical Society. Series A (General), 118(2):258–262, ???? 1955. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343145.

Anonymous:1955:SEAb

[559] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society. Series A (General), 118(2):263–267, ???? 1955. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343146.

Anonymous:1955:LALb

[560] Anonymous. List of additions to the library. Journal of the Royal Statistical Society. Series A (General), 118 (2):268–274, ???? 1955. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2343147.

Thom:1955:SEM

[561] A. Thom. A statistical examination of the megalithic sites in Britain. Journal of the Royal Statistical Society. Series A (General), 118(3):275–295, ???? 1955. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2342494.

Lane:1955:MLT

[562] K. F. Lane and J. E. Andrew. A method of labour turnover analysis. Journal of the Royal Statistical Society. Series A (General), 118(3):296– 323, ???? 1955. CODEN JSSAEF.

ISSN 0035-9238. URL http://www.

jstor.org/stable/2342495.

Cox:1955:SAC

[563] D. R. Cox. The statistical analysis of congestion. Journal of the Royal Statistical Society. Series A (General), 118 (3):324–335, ???? 1955. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342496.

Bellerby:1955:AI

[564] J. R. Bellerby. Agricultural income. Journal of the Royal Statistical Society. Series A (General), 118(3):336– 344, ???? 1955. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342497.

Wilson:1955:SNS

[565] Ross Wilson. The sources and nature of statistical information in special fields of statistics: The Scotch whisky industry. Journal of the Royal Statistical Society. Series A (General), 118(3):345– 362, ???? 1955. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342498.

Gorman:1955:BRB

[566] W. M. Gorman. Book review: Linear Aggregation of Economic Relations, by H. Theil. Journal of the Royal Statistical Society. Series A (General), 118 (3):363–364, ???? 1955. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342499.

Smith:1955:BRB

[567] Cedric A. B. Smith. Book review: From Simple Numbers to the Calculus: Mathematics for Everyman, by Egmont Colerus; B. C. Brookes; H. F. Brookes.

Journal of the Royal Statistical Society. Series A (General), 118(3):364, ???? 1955. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342500.

Hammersley:1955:BRBb

[568] J. M. Hammersley. Book review: A Million Random Digits with 100,000 Normal Deviates. Journal of the Royal Statistical Society. Series A (General), 118(3):364–365, ???? 1955. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2342501.

Wishart:1955:BRBa

[569] J. Wishart. Book review: Tables of Sine and Cosine Integrals for Arguments from 10 to 100; Table of the Gamma Function for Complex Arguments; Tables of Generalized k-Statistics, by S. H. Abdel-Aty. Journal of the Royal Statistical Society. Series A (General), 118(3):365–366, ???? 1955. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342502.

Wishart:1955:BRBb

[570] J. Wishart. Book review: Computer Development (SEAC and DYSEAC) at the National Bureau of Standards, Washington, DC. Journal of the Royal Statistical Society. Series A (General), 118(3):366, ???? 1955. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342503.

Benjamin:1955:BRBa

[571] B. Benjamin. Book review: Addendum to Demography–Developments during 1949–54, by Peter R. Cox. Journal of the Royal Statistical Society.

Series A (General), 118(3):367, ???? 1955. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342504.

Benjamin:1955:BRBb

[572] B. Benjamin. Book review: A Theory of Economic-Demographic Development, by Harvey Leibenstein. Journal of the Royal Statistical Society. Series A (General), 118(3):367–368, ???? 1955. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342505.

Brown:1955:BRB

[573] J. A. C. Brown. Book review: Short-Term Economic Forecasting. Journal of the Royal Statistical Society. Series A (General), 118(3):368–369, ???? 1955. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342506.

Glenday:1955:BRBb

[574] R. Glenday. Book review: Sixth Report of the O.E.E.C., From Recovery Towards Economic Strength. Journal of the Royal Statistical Society. Series A (General), 118(3):369, ???? 1955. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342507.

Deane:1955:BRB

[575] P. M. Deane. Book review: A Study of Saving in the United States. Volume I: Introduction–Tables of Annual Estimates of Saving 1897 to 1949, by Raymond W. Goldsmith; A Study of Saving in the United States. Volume II: Nature and Derivation of Annual Estimates of Saving 1897 to 1949 by Raymond W. Goldsmith. Journal of the Royal Statistical Society. Series A (General), 118(3):370–371, ???? 1955. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2342508.

Frankel:1955:BRB

[576] S. Herbert Frankel. Book review: The National Income of Nigeria 1950–51, by A. R. Prest; I. G. Stewart. Journal of the Royal Statistical Society. Series A (General), 118(3):371–372, ???? 1955. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2342509.

Anonymous:1955:SNc

[577] Anonymous. Statistical notes. Journal of the Royal Statistical Society. Series A (General), 118(3):373–374, ???? 1955. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2342510.

Anonymous:1955:CNb

[578] Anonymous. Current note. Journal of the Royal Statistical Society. Series A (General), 118(3):375, ???? 1955. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2342511.

Anonymous:1955:SEAc

[579] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society. Series A (General), 118(3):376–379, ???? 1955. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2342512.

Anonymous:1955:LALc

[580] Anonymous. List of additions to the library. Journal of the Royal Statistical Society. Series A (General), 118 (3):380–383, ???? 1955. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342513.

Anonymous:1955:PRb

[581] Anonymous. Periodical returns. Journal of the Royal Statistical Society. Series A (General), 118(3):384–387, ???? 1955. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342514.

Anonymous:1955:VIb

[582] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (General), 118(4):515– 517, ???? 1955. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342708.

Irwin:1955:UDS

[583] J. O. Irwin. A unified derivation of some well-known frequency distributions of interest in biometry and statistics. Journal of the Royal Statistical Society. Series A (General), 118(4):389– 404, ???? 1955. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342709.

Liddell:1955:CS

[584] F. D. K. Liddell. Colliery statistics. Journal of the Royal Statistical Society. Series A (General), 118(4):405– 416, ???? 1955. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342710.

West:1955:ALN

[585] John H. West. An analysis of 162,332 lottery numbers. Journal of the Royal Statistical Society. Series A (General), 118(4):417–426, ???? 1955. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2342711.

Anonymous:1955:SAN

[586] Anonymous. Statistical assessment of the need for clinical services. Journal of the Royal Statistical Society. Series A (General), 118(4):427–433, ???? 1955. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342712.

Utting:1955:SNS

[587] J. E. G. Utting. The sources and nature of statistical information in special fields of statistics: National income and related statistics. Journal of the Royal Statistical Society. Series A (General), 118(4):434–451, ???? 1955. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2342713.

Grebenik:1955:SNS

[588] E. Grebenik. The sources and nature of statistical information in special fields of statistics: Population and vital statistics. Journal of the Royal Statistical Society. Series A (General), 118 (4):452–464, ???? 1955. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342714.

Anonymous:1955:RC

[589] Anonymous. Report of the Council. Journal of the Royal Statistical Society. Series A (General), 118(4):465– 482, ???? 1955. CODEN JSSAEF.

ISSN 0035-9238. URL http://www.

jstor.org/stable/2342715.

Anonymous:1955:POH

[590] Anonymous. Proceedings of the one hundred and twenty-first Annual General Meeting of the Royal Statistical Society held at the London School of Hygiene and Tropical Medicine on Wednesday, June 22nd, 1955, at 5.30. Journal of the Royal Statistical Society. Series A (General), 118(4):m, ???? 1955. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342716.

Bailey:1955:BRB

[591] Norman T. J. Bailey. Book review: An Introduction to Stochastic Processes: With Special Reference to Methods and Application, by M. S. Bartlett. Journal of the Royal Statistical Society. Series A (General), 118(4):484–485, ???? 1955. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342717.

Johnson:1955:BRBa

[592] N. L. Johnson. Book review: Chance and Choice, by Cardpack and Chessboard. … Vol. II, by Lancelot Hogben. Journal of the Royal Statistical Society. Series A (General), 118(4):485– 486, ???? 1955. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342718.

Johnson:1955:BRBb

[593] N. L. Johnson. Book review: The Elements of Probability Theory and Some of Its Applications, by H. Cramér. Journal of the Royal Statistical Society. Series A (General), 118(4):486, ????

1955. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2342719.

Anscombe:1955:BRB

[594] F. J. Anscombe. Book review: Statistical Methods for Research Workers, by Ronald A. Fisher. Journal of the Royal Statistical Society. Series A (General), 118(4):486–487, ???? 1955. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2342720.

Oldham:1955:BRB

[595] P. D. Oldham. Book review: Principles of Medical Statistics, by A. Bradford Hill. Journal of the Royal Statistical Society. Series A (General), 118 (4):487, ???? 1955. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2342721.

Lindley:1955:BRBb

[596] D. V. Lindley. Book review: Selected Papers in Statistics and Probability, by Abraham Wald. Journal of the Royal Statistical Society. Series A (General), 118(4):488–489, ???? 1955. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2342722.

Armitage:1955:BRBb

[597] P. Armitage. Book review: Graphical Method of Statistical Inference, by M. Masuyama. Journal of the Royal Statistical Society. Series A (General), 118 (4):489–490, ???? 1955. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2342723.

Foster:1955:BRB

[598] F. G. Foster. Book review: A Distribution Model Applicable to Economics, by Arne Jensen. Journal of the Royal Statistical Society. Series A (General), 118(4):490, ???? 1955. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342724.

Wilkins:1955:BRB

[599] Leslie T. Wilkins. Book review: Interviewing in Social Research, by Herbert H. Hyman; William J. Cobb; Jacob J. Feldman; Clyde W. Hart; Charles Herbert Stember. Journal of the Royal Statistical Society. Series A (General), 118 (4):490–491, ???? 1955. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342725.

Milne:1955:BRB

[600] J. M. N. Milne. Book review: Industrial Censuses in the United States. Journal of the Royal Statistical Society. Series A (General), 118(4):491–492, ???? 1955. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342726.

Redfern:1955:BRB

[601] Philip Redfern. Book review: Private United States Investment in Europe and the Overseas Territories. Journal of the Royal Statistical Society. Series A (General), 118(4):492–493, ???? 1955. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342727.

Anonymous:1955:SNd

[602] Anonymous. Statistical notes. Journal of the Royal Statistical Society. Series A (General), 118(4):494–496, ???? 1955. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342728.

Anonymous:1955:CNc

[603] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 118(4):496, ???? 1955. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2342729.

Tippett:1955:PL

[604] L. H. C. Tippett. Philip Lyle. Journal of the Royal Statistical Society. Series A (General), 118(4):497–498, ???? 1955. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2342730.

Anonymous:1955:SJ

[605] Anonymous. Stanley Jevons. Journal of the Royal Statistical Society. Series A (General), 118(4):498–499, ???? 1955. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2342731.

Nixon:1955:GFS

[606] J. W. Nixon. George Findlay Shirras. Journal of the Royal Statistical Society. Series A (General), 118(4):499– 500, ???? 1955. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2342732.

Anonymous:1955:SEAd

[607] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society. Series A (General), 118(4):501–505, ???? 1955. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2342733.

Anonymous:1955:LALd

[608] Anonymous. List of additions to the library. Journal of the Royal Statistical Society. Series A (General), 118 (4):506–513, ???? 1955. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342734.

Anonymous:1955:PRc

[609] Anonymous. Periodical returns. Journal of the Royal Statistical Society. Series A (General), 118(4):514, ???? 1955. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342735.

Anonymous:1956:FMa

[610] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 119(1):??, ???? 1956. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342959.

Finney:1956:SPF

[611] D. J. Finney. The statistician and the planning of field experiments. Journal of the Royal Statistical Society. Series A (General), 119(1):1–27, ???? 1956. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342960.

Wold:1956:CIO

[612] Herman Wold. Causal inference from observational data: A review of end and means. Journal of the Royal Statistical Society. Series A (General), 119 (1):28–61, ???? 1956. CODEN JS-SAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2342961.

Anstey:1956:SAS

[613] Vera Anstey and Russi J. Taraporevala. Some aspects of the structure of Indian industry. Journal of the Royal Statistical Society. Series A (General), 119 (1):62–82, ???? 1956. CODEN JS-SAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2342962.

Kendall:1956:PFA

[614] M. G. Kendall and D. N. Lawley. The principles of factor analysis. Journal of the Royal Statistical Society. Series A (General), 119(1):83–84, ???? 1956. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342963.

Anonymous:1956:MS

[615] Anonymous. Medical section. Journal of the Royal Statistical Society. Series A (General), 119(1):85, ???? 1956. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342964.

Lindley:1956:BRBa

[616] D. V. Lindley. Book review: Theory of Games as a Tool for the Moral Philosopher, by R. B. Braithwaite. Journal of the Royal Statistical Society. Series A (General), 119(1):86–87, ???? 1956. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342965.

Vajda:1956:BRB

[617] S. Vajda. Book review: Stochastic Models for Learning, by Robert R. Bush; Frederick Mosteller. Journal of the Royal Statistical Society. Series A (General), 119(1):87, ????

1956. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342966.

Wishart:1956:BRBa

[618] J. Wishart. Book review: Statistics in Research, by Bernard Ostle. Journal of the Royal Statistical Society. Series A (General), 119(1):87–88, ???? 1956. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342967.

Martin:1956:BRBa

[619] Leslie V. Martin. Book review: Mortality and Other Investigations, by H. W. Haycocks; W. Perks. Journal of the Royal Statistical Society. Series A (General), 119(1):88–89, ???? 1956. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342968.

Healy:1956:BRB

[620] M. J. R. Healy. Book review: Principles and Practice of Field Experimentation, by J. Wishart; H. G. Sanders. Journal of the Royal Statistical Society. Series A (General), 119(1):89–90, ???? 1956. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2342969.

David:1956:BRB

[621] F. N. David. Book review: Rank Correlation Methods, by M. G. Kendall. Journal of the Royal Statistical Society. Series A (General), 119(1):90, ???? 1956. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342970.

Wishart:1956:BRBb

[622] J. Wishart. Book review: Piano degli Esperimenti, by Giuseppe Pompilj; Diego Napolitani. Journal of the Royal Statistical Society. Series A (General), 119(1):90, ???? 1956. CODEN JS-SAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2342971.

Burt:1956:BRB

[623] Cyril Burt. Book review: Prediction Methods in Relation to Borstal Training, by Hermann Mannheim; Leslie T. Wilkins. Journal of the Royal Statistical Society. Series A (General), 119(1): 91, ???? 1956. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2342972.

Norris:1956:BRB

[624] V. Norris. Book review: Suicide in London: An Ecological Study, by Peter Sainsbury. Journal of the Royal Statistical Society. Series A (General), 119 (1):92–93, ???? 1956. CODEN JS-SAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2342973.

Cox:1956:BRB

[625] P. R. Cox. Book review: World Population and Resources, by P. E. P. Journal of the Royal Statistical Society. Series A (General), 119(1):93–94, ???? 1956. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2342974.

Rivett:1956:BRB

[626] B. H. P. Rivett. Book review: Operations Research for Management, by J. F. McCloskey; F. N. Trefethen. Journal of the Royal Statistical Society. Series A (General), 119(1):94–95, ????

1956. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342975.

Deane:1956:BRBa

[627] P. M. Deane. Book review: The Redistribution of Income in Postwar Britain: A Study in the Effects of the Central Government Fiscal Program in 1948– 49, by Allan M. Cartter. Journal of the Royal Statistical Society. Series A (General), 119(1):95–96, ???? 1956. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342976.

Clark:1956:BRBa

[628] Colin Clark. Book review: Labor Productivity in Soviet and American Industry, by Walter Galenson. Journal of the Royal Statistical Society. Series A (General), 119(1):96–97, ???? 1956. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342977.

vanRest:1956:BRB

[629] E. D. van Rest. Book review: Control Chart Technique when Manufacturing to a Specification, by B. P. Dudding; W. J. Jennett. Journal of the Royal Statistical Society. Series A (General), 119(1):97–98, ???? 1956. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342978.

Anonymous:1956:SNa

[630] Anonymous. Statistical notes. Journal of the Royal Statistical Society. Series A (General), 119(1):99–101, ???? 1956. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342979.

Anonymous:1956:CNa

[631] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 119(1):101, ???? 1956. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2342980.

Bowley:1956:HWM

[632] A. L. Bowley. H. W. Methorst — 1868– 1955. Journal of the Royal Statistical Society. Series A (General), 119(1): 102, ???? 1956. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2342981.

Murray:1956:CSO

[633] Keith A. H. Murray. Charles Stewart Orwin. Journal of the Royal Statistical Society. Series A (General), 119 (1):103–104, ???? 1956. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2342982.

Anonymous:1956:SEAa

[634] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society. Series A (General), 119(1):105–111, ???? 1956. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2342983.

Anonymous:1956:LALa

[635] Anonymous. List of additions to the library. Journal of the Royal Statistical Society. Series A (General), 119 (1):112–121, ???? 1956. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2342984.

Anonymous:1956:PRa

[636] Anonymous. Periodical returns. Journal of the Royal Statistical Society. Series A (General), 119(1):122–123, ???? 1956. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342985.

Anonymous:1956:FMb

[637] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 119(2):??, ???? 1956. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342879.

Pearson:1956:SAG

[638] E. S. Pearson. Some aspects of the geometry of statistics: The use of visual presentation in understanding the theory and application of mathematical statistics. Journal of the Royal Statistical Society. Series A (General), 119 (2):125–146, ???? 1956. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342880.

Anonymous:1956:PM

[639] Anonymous. Proceedings of the meeting. Journal of the Royal Statistical Society. Series A (General), 119(2):146– 149, ???? 1956. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342881.

Hart:1956:ABC

[640] P. E. Hart and S. J. Prais. The analysis of business concentration: A statistical approach. Journal of the Royal Statistical Society. Series A (General), 119 (2):150–191, ???? 1956. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342882.

Glenday:1956:RJR

[641] R. Glenday. Richard Jones: A reappraisal. Journal of the Royal Statistical Society. Series A (General), 119 (2):192–193, ???? 1956. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2342883.

Anonymous:1956:WP

[642] Anonymous. Wholesale prices in 1955. Journal of the Royal Statistical Society. Series A (General), 119(2):194– 198, ???? 1956. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2342884.

Dacey:1956:SNS

[643] W. Manning Dacey. The sources and nature of statistical information in special fields of statistics: Banking statistics. Journal of the Royal Statistical Society. Series A (General), 119(2):199– 209, ???? 1956. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2342885.

Bailey:1956:BRB

[644] Norman T. J. Bailey. Book review: Experimental Design and its Statistical Basis, by D. J. Finney. Journal of the Royal Statistical Society. Series A (General), 119(2):210–211, ???? 1956. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2342886.

Johnson:1956:BRBa

[645] N. L. Johnson. Book review: Basic Statistical Concepts, by J. K. Adams. Journal of the Royal Statistical Society. Series A (General), 119(2):211, ????

1956. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342887.

Chambers:1956:BRB

[646] E. G. Chambers. Book review: Risk and Gambling, by J. Cohen; M. Hansel. Journal of the Royal Statistical Society. Series A (General), 119(2):211– 212, ???? 1956. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342888.

Tocher:1956:BRBa

[647] K. D. Tocher. Book review: Numerical Analysis, by Zdenek Kopal. Journal of the Royal Statistical Society. Series A (General), 119(2):212–213, ???? 1956. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342889.

Good:1956:BRB

[648] I. J. Good. Book review: Fondements d’une Theorie Positive des Choix Comportant un Risque et Critique des Postulats et Axiomes de l’Ecole Americaine, by M. Allais. Journal of the Royal Statistical Society. Series A (General), 119(2):213–214, ???? 1956. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342890.

Benjamin:1956:BRB

[649] B. Benjamin. Book review: The Cost of the National Health Service in England and Wales, by Brian Abel-Smith; Richard M. Titmuss. Journal of the Royal Statistical Society. Series A (General), 119(2):214–215, ???? 1956. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342891.

Barr:1956:BRBa

[650] A. Barr. Book review: The Bethlem Royal Hospital and the Maudsley Hospital Triennial Statistical Report, 1949–1951, by C. P. Blacker; A. T. Gore. Journal of the Royal Statistical Society. Series A (General), 119(2): 215–216, ???? 1956. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2342892.

Thomas:1956:BRBa

[651] Brinley Thomas. Book review: Migration Probabilities, by G. Kulldorff. Journal of the Royal Statistical Society. Series A (General), 119(2):216, ???? 1956. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2342893.

Thomas:1956:BRBb

[652] Brinley Thomas. Book review: Planned Migration: The Social Determinants of the Dutch-Canadian Movement, by William Petersen. Journal of the Royal Statistical Society. Series A (General), 119(2):216–217, ???? 1956. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2342894.

Adams:1956:BRBa

[653] A. A. Adams. Book review: Income and Wealth, Series IV, by Milton Gilbert; Richard Stone. Journal of the Royal Statistical Society. Series A (General), 119(2):217–218, ???? 1956. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2342895.

Prais:1956:BRB

[654] S. J. Prais. Book review: Structure of Indian Industries, by M. M. Mehta.

Journal of the Royal Statistical Society. Series A (General), 119(2):218– 219, ???? 1956. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342896.

Lydall:1956:BRBa

[655] Harold Lydall. Book review: Income of the American People, by Herman P. Miller. Journal of the Royal Statistical Society. Series A (General), 119 (2):219–220, ???? 1956. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342897.

Anonymous:1956:SNb

[656] Anonymous. Statistical notes. Journal of the Royal Statistical Society. Series A (General), 119(2):221–224, ???? 1956. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342898.

Anonymous:1956:SEAb

[657] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society. Series A (General), 119(2):225–230, ???? 1956. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342899.

Anonymous:1956:LALb

[658] Anonymous. List of additions to the library. Journal of the Royal Statistical Society. Series A (General), 119 (2):231–237, ???? 1956. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342900.

Anonymous:1956:BMa

[659] Anonymous. Back matter. Journal of the Royal Statistical Society.

Series A (General), 119(2):??, ???? 1956. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342901.

Anonymous:1956:FMc

[660] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 119(3):??, ???? 1956. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342736.

Phillips:1956:UKI

[661] H. S. Phillips. United Kingdom indices of wholesale prices, 1949–1955. Journal of the Royal Statistical Society. Series A (General), 119(3):239–283, ???? 1956. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342737.

Emmett:1956:TAU

[662] B. P. Emmett. The television audience in the United Kingdom. Journal of the Royal Statistical Society. Series A (General), 119(3):284–311, ???? 1956. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342738.

Seng:1956:STIa

[663] You Poh Seng. Some theory of index numbers. Journal of the Royal Statistical Society. Series A (General), 119 (3):312–332, ???? 1956. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342739.

Cox:1956:GDA

[664] C. P. Cox. A geometrical derivation of the analyses of covariance and variance. Journal of the Royal Statistical Society. Series A (General), 119(3):

333–335, ???? 1956. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342740.

Zarkovic:1956:NHS

[665] S. S. Zarkovic. Note on the history of sampling methods in Russia. Journal of the Royal Statistical Society. Series A (General), 119(3):336–338, ???? 1956. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342741.

Briggs:1956:BRB

[666] F. E. A. Briggs. Book review: Elementary Statistics, by R. Clay Sprowls. Journal of the Royal Statistical Society. Series A (General), 119(3):339– 340, ???? 1956. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342742.

Mauldon:1956:BRB

[667] J. G. Mauldon. Book review: The Theory of Games and Linear Programming, by S. Vajda. Journal of the Royal Statistical Society. Series A (General), 119(3):340, ???? 1956. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342743.

Nelder:1956:BRB

[668] J. A. Nelder. Book review: Experimental Design, Theory and Application, by Walter T. Federer. Journal of the Royal Statistical Society. Series A (General), 119(3):340–341, ???? 1956. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2342744.

Maurice:1956:BRB

[669] R. J. Maurice. Book review: A Primer of Social Statistics, by S. M.

Dornbusch; C. F. Schmid. Journal of the Royal Statistical Society. Series A (General), 119(3):341–342, ???? 1956. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342745.

Booker:1956:BRB

[670] H. S. Booker. Book review: Statistics, by L. H. C. Tippett. Journal of the Royal Statistical Society. Series A (General), 119(3):342–343, ???? 1956. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342746.

Moore:1956:BRB

[671] P. G. Moore. Book review: Methods in Numerical Analysis, by K. L. Nielsen. Journal of the Royal Statistical Society. Series A (General), 119 (3):343–344, ???? 1956. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342747.

Tocher:1956:BRBb

[672] K. D. Tocher. Book review: Table of Hyperbolic Sines and Cosines x = 2 to x = 10. Journal of the Royal Statistical Society. Series A (General), 119 (3):344, ???? 1956. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342748.

Tocher:1956:BRBc

[673] K. D. Tocher. Book review: Table of the Descending Exponential x = 2.5 to x = 10. Journal of the Royal Statistical Society. Series A (General), 119 (3):344, ???? 1956. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342749.

Tocher:1956:BRBd

[674] K. D. Tocher. Book review: Tables of the Cumulate Binomial Probability Distribution. Journal of the Royal Statistical Society. Series A (General), 119 (3):344–345, ???? 1956. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342750.

Tocher:1956:BRBe

[675] K. D. Tocher. Book review: Correlated Random Normal Deviates, by E. C. Fieller; T. Lewis; E. S. Pearson. Journal of the Royal Statistical Society. Series A (General), 119(3):345–346, ???? 1956. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342751.

Tocher:1956:BRBf

[676] K. D. Tocher. Book review: Rectangular-Polar Conversion Tables, by E. H. Neville. Journal of the Royal Statistical Society. Series A (General), 119(3):346, ???? 1956. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342752.

Lydall:1956:BRBb

[677] H. F. Lydall. Book review: Report on Proposals for a New Index of Retail Prices. Journal of the Royal Statistical Society. Series A (General), 119 (3):346–347, ???? 1956. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342753.

Adams:1956:BRBb

[678] A. A. Adams. Book review: British Industry, 1700–1950, by W. Hoffman; W. H. Chaloner; A. O. Henderson. Journal of the Royal Statistical Society. Series A (General), 119(3):347–348, ????

1956. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342754.

Barr:1956:BRBb

[679] A. Barr. Book review: Migration and Mental Disease: A Study of First Admissions to Hospitals for Mental Disease, New York, 1939–1941, by Benjamin Malzburg; Everett M. Lee. Journal of the Royal Statistical Society. Series A (General), 119(3):348–349, ???? 1956. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342755.

Heady:1956:BRB

[680] J. A. Heady. Book review: Health in Industry. Journal of the Royal Statistical Society. Series A (General), 119 (3):349, ???? 1956. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342756.

Nixon:1956:BRB

[681] J. W. Nixon. Book review: The European Housing Situation. Journal of the Royal Statistical Society. Series A (General), 119(3):349–350, ???? 1956. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342757.

Clark:1956:BRBb

[682] Colin Clark. Book review: World Commerce and Governments, by W. S. Woytinsky; E. S. Woytinsky. Journal of the Royal Statistical Society. Series A (General), 119(3):350–351, ???? 1956. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342758.

Elphinstone:1956:BRB

[683] M. D. W. Elphinstone. Book review: The History of the Faculty of Actuaries in Scotland 1856–1956, by A. R. Davidson. Journal of the Royal Statistical Society. Series A (General), 119(3):351, ???? 1956. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2342759.

Anonymous:1956:SNc

[684] Anonymous. Statistical notes. Journal of the Royal Statistical Society. Series A (General), 119(3):352–354, ???? 1956. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342760.

Anonymous:1956:SEAc

[685] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society. Series A (General), 119(3):355–360, ???? 1956. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342761.

Anonymous:1956:LALc

[686] Anonymous. List of additions to the library. Journal of the Royal Statistical Society. Series A (General), 119 (3):361–366, ???? 1956. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342762.

Anonymous:1956:PRb

[687] Anonymous. Periodical returns. Journal of the Royal Statistical Society. Series A (General), 119(3):367–370, ???? 1956. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342763.

Anonymous:1956:BMb

[688] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 119(3):??, ???? 1956. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2342764.

Anonymous:1956:VI

[689] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (General), 119(4):506– 508, ???? 1956. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2342574.

Anonymous:1956:FMd

[690] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 119(4):??, ???? 1956. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2342575.

Cole:1956:EES

[691] Dorothy Cole and J. E. G. Utting. Estimating expenditure, saving and income from household budgets. Journal of the Royal Statistical Society. Series A (General), 119(4):371–392, ???? 1956. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/

stable/2342576. See notes [723].

Reynolds:1956:CRA

[692] D. J. Reynolds. The cost of road accidents. Journal of the Royal Statistical Society. Series A (General), 119 (4):393–408, ???? 1956. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2342577.

Taylor:1956:NZA

[693] W. B. Taylor and D. V. P. Clement. The New Zealand Agricultural Sample Survey. Journal of the Royal Statistical Society. Series A (General), 119 (4):409–424, ???? 1956. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342578.

Seng:1956:STIb

[694] You Poh Seng. Some theory of index numbers — II: Maintenance of continuity by the assumption of proportionality of the new index number to the old index. Journal of the Royal Statistical Society. Series A (General), 119 (4):425–455, ???? 1956. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342579.

Bartlett:1956:NVM

[695] M. S. Bartlett. Note on a visit to Moscow for the third Soviet mathematical Congress. Journal of the Royal Statistical Society. Series A (General), 119 (4):456–458, ???? 1956. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342580.

Anonymous:1956:RC

[696] Anonymous. Report of the Council. Journal of the Royal Statistical Society. Series A (General), 119(4):459– 476, ???? 1956. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342581.

Anonymous:1956:POH

[697] Anonymous. Proceedings of the one hundred and twenty-second Annual General Meeting of the Royal Statistical Society held at the London School of Hygiene and Tropical Medicine on Wednesday, June 20th, 1956, at 5.30. Journal of the Royal Statistical Society. Series A (General), 119(4):m, ???? 1956. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342582.

Corlett:1956:BRB

[698] W. J. Corlett. Book review: An Introduction to British Economic Statistics, by E. Devons. Journal of the Royal Statistical Society. Series A (General), 119(4):478–479, ???? 1956. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2342583.

Lindley:1956:BRBb

[699] D. V. Lindley. Book review: Statistics: A New Approach, by W. Allen Wallis; Harry V. Roberts. Journal of the Royal Statistical Society. Series A (General), 119(4):479–480, ???? 1956. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2342584.

Johnson:1956:BRBb

[700] N. L. Johnson. Book review: The Essentials of Educational Statistics, by F. G. Cornell. Journal of the Royal Statistical Society. Series A (General), 119 (4):480, ???? 1956. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342585.

Martin:1956:BRBb

[701] J. P. Martin. Book review: Juvenile Offenders before the Courts, by Max Grunhut. Journal of the Royal Statistical Society. Series A (General), 119 (4):480–481, ???? 1956. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342586.

Roy:1956:BRB

[702] A. D. Roy. Book review: Individuals, Groups, and Economic Behavior, by C. Addison Hickman; Manford H. Kuhn. Journal of the Royal Statistical Society. Series A (General), 119(4): 481–482, ???? 1956. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342587.

Plymen:1956:BRB

[703] J. Plymen. Book review: Urban Mortgage Lending. Comparative Markets and Experiences, by J. E. Morton. Journal of the Royal Statistical Society. Series A (General), 119(4):482– 484, ???? 1956. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342588.

Deane:1956:BRBb

[704] P. M. Deane. Book review: Capital Formation and Economic Growth. A Conference of the Universities, National Bureau Committee for Economic Research. Journal of the Royal Statistical Society. Series A (General), 119 (4):484, ???? 1956. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342589.

Martin:1956:BRBc

[705] W. J. Martin. Book review: Annual Epidemiological and Vital Statistics, 1953. Journal of the Royal Statistical Society. Series A (General), 119 (4):484–486, ???? 1956. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342590.

Deane:1956:BRBc

[706] Marjorie Deane. Book review: Books for All, by R. E. Barker. Journal of the Royal Statistical Society. Series A (General), 119(4):486, ???? 1956. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2342591.

Gorman:1956:BRB

[707] W. M. Gorman. Book review: Cahiers du Seminaire d’Econometrie No. 3. Les Modeles Econometriques, by Rene Roy. Journal of the Royal Statistical Society. Series A (General), 119(4):486– 487, ???? 1956. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2342592.

Anonymous:1956:SNd

[708] Anonymous. Statistical notes. Journal of the Royal Statistical Society. Series A (General), 119(4):488–490, ???? 1956. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2342593.

Anonymous:1956:CNb

[709] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 119(4):490–491, ???? 1956. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2342594.

Bartlett:1956:JWD

[710] M. S. Bartlett. John Wishart, D.Sc., F.R.S.E. Journal of the Royal Statistical Society. Series A (General), 119 (4):492–493, ???? 1956. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2342595.

Reddaway:1956:AAA

[711] W. B. Reddaway. Arthur Avery Adams. Journal of the Royal Statistical Society. Series A (General), 119 (4):493–494, ???? 1956. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342596.

Hopkin:1956:MJR

[712] W. A. B. Hopkin. Mrs. J. R. N. Stone. Journal of the Royal Statistical Society. Series A (General), 119(4):494, ???? 1956. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342597.

Anonymous:1956:SEAd

[713] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society. Series A (General), 119(4):495–499, ???? 1956. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342598.

Anonymous:1956:LALd

[714] Anonymous. List of additions to the library. Journal of the Royal Statistical Society. Series A (General), 119 (4):500–504, ???? 1956. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342599.

Anonymous:1956:PRc

[715] Anonymous. Periodical returns. Journal of the Royal Statistical Society. Series A (General), 119(4):505, ???? 1956. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342600.

Anonymous:1956:BMc

[716] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 119(4):??, ???? 1956. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342601.

Anonymous:1957:VIa

[717] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (General), 120(1):??, ???? 1957. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2342550.

Anonymous:1957:FM

[718] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 120(1):??, ???? 1957. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2342551.

Barna:1957:RCF

[719] T. Barna. The replacement cost of fixed assets in British manufacturing industry in 1955. Journal of the Royal Statistical Society. Series A (General), 120 (1):1–47, ???? 1957. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2342552.

Bartlett:1957:MPC

[720] M. S. Bartlett. Measles periodicity and community size. Journal of the Royal Statistical Society. Series A (General), 120(1):48–70, ???? 1957. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2342553.

Parkinson:1957:SWR

[721] J. R. Parkinson. Ship wastage rates. Journal of the Royal Statistical Society. Series A (General), 120(1):71–83, ???? 1957. CODEN JSSAEF. ISSN

                                        0035-9238. URL http://www.jstor.
                                        org/stable/2342554.

Nicholson:1957:GFA

[722] J. L. Nicholson. The general form of the adding-up criterion. Journal of the Royal Statistical Society. Series A (General), 120(1):84–85, ???? 1957. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342555.

Durant:1957:FNP

[723] Henry Durant, Dorothy Cole, and J. E. G. Utting. Further notes on the paper by J. E. G. Utting and Dorothy Cole on estimating expenditure, saving and income from household budgets. Journal of the Royal Statistical Society. Series A (General), 120(1):86– 87, ???? 1957. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2342556. See [691].

Bailey:1957:BRB

[724] Norman T. J. Bailey. Book review: Statistical Methods and Scientific Inference, by Ronald Fisher. Journal of the Royal Statistical Society. Series A (General), 120(1):88–89, ???? 1957. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342557.

Good:1957:BRB

[725] I. J. Good. Book review: Language as Choice and Chance, by G. Herdan. Journal of the Royal Statistical Society. Series A (General), 120(1):89–91, ???? 1957. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2342558.

Hammersley:1957:BRB

[726] J. M. Hammersley. Book review: Symposium on Monte Carlo Methods, by Herbert A. Meyer. Journal of the Royal Statistical Society. Series A (General), 120(1):91, ???? 1957. CODEN JS-SAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2342559.

Tocher:1957:BRBa

[727] K. D. Tocher. Book review: Numerical Analysis. Proceedings of Symposia in Applied Mathematics. Journal of the Royal Statistical Society. Series A (General), 120(1):91–92, ???? 1957. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342560.

Tocher:1957:BRBb

[728] K. D. Tocher. Book review: A New Form of Table for Significance Tests in a 2 × 2 Contingency Table, by P. Armsen. Journal of the Royal Statistical Society. Series A (General), 120(1):92, ???? 1957. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2342561.

Wilkins:1957:BRB

[729] Leslie T. Wilkins. Book review: CIBA Foundation Symposium on Extrasensory Perception, by G. E. W. Wolstenholme; Elaine C. P. Millar. Journal of the Royal Statistical Society. Series A (General), 120(1):93, ???? 1957. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342562.

Booker:1957:BRB

[730] H. S. Booker. Book review: Dynamic Factors in Industrial Productivity, by Seymour Melman. Journal of the Royal Statistical Society. Series A (General), 120(1):93–94, ???? 1957. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342563.

Saunders:1957:BRB

[731] C. T. Saunders. Book review: National Income Visualised, by Arthur O. Dahlberg. Journal of the Royal Statistical Society. Series A (General), 120 (1):94–95, ???? 1957. CODEN JS-SAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2342564.

Stewart:1957:BRB

[732] C. M. Stewart. Book review: Introduction to Demography, by Mortimer Spiegelman. Journal of the Royal Statistical Society. Series A (General), 120 (1):95–96, ???? 1957. CODEN JS-SAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2342565.

Rivett:1957:BRB

[733] B. H. P. Rivett. Book review: Operations Research for Management, by Joseph F. McCloskey; John M. Coppinger. Journal of the Royal Statistical Society. Series A (General), 120(1):96– 97, ???? 1957. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2342566.

Knapp:1957:BRB

[734] John Knapp. Book review: Distribution’s Place in the American Economy since 1869, by Harold Barger. Journal of the Royal Statistical Society. Series A (General), 120(1):97–98, ???? 1957. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342567.

Maurice:1957:BRB

[735] Rita J. Maurice. Book review: Method of Construction and Calculation of the Index of Retail Prices. Journal of the Royal Statistical Society. Series A (General), 120(1):98–99, ???? 1957. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342568.

Lewis:1957:BRB

[736] J. Parry Lewis. Book review: The Social Accounts of the Welsh Economy, 1948 to 1952, by Edward Nevin. Journal of the Royal Statistical Society. Series A (General), 120(1):99–100, ???? 1957. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342569.

Hewitt:1957:BRB

[737] D. Hewitt. Book review: Childbirth, Complications and Mortality of Obstetrical Patients in British Columbia (Health Statistics Research Survey), by Zygmunt Lehoczky. Journal of the Royal Statistical Society. Series A (General), 120(1):100–101, ???? 1957. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342570.

Anonymous:1957:SNa

[738] Anonymous. Statistical notes. Journal of the Royal Statistical Society. Series A (General), 120(1):102–104, ???? 1957. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342571.

Anonymous:1957:SEAa

[739] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society. Series A (General), 120(1):105–110, ???? 1957. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342572.

Anonymous:1957:LALa

[740] Anonymous. List of additions to the library. Journal of the Royal Statistical Society. Series A (General), 120 (1):111–120, ???? 1957. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342573.

Gales:1957:ICI

[741] Kathleen Gales and M. G. Kendall. An inquiry concerning interviewer variability. Journal of the Royal Statistical Society. Series A (General), 120(2):121– 147, ???? 1957. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342821.

Horsnell:1957:EAS

[742] G. Horsnell. Economical acceptance sampling schemes. Journal of the Royal Statistical Society. Series A (General), 120(2):148–201, ???? 1957. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2342822.

Lea:1957:CPM

[743] A. J. Lea. Changes in the proportions of male colour types during the period 1914–39. Journal of the Royal Statistical Society. Series A (General), 120 (2):202–208, ???? 1957. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342823.

Sarkar:1957:MES

[744] N. K. Sarkar. A method of estimating surplus labour in peasant agriculture in over-populated under-developed countries. Journal of the Royal Statistical Society. Series A (General), 120 (2):209–214, ???? 1957. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342824.

Maizels:1957:UVV

[745] A. Maizels. Unit value and volume index numbers of inter-area trade. Journal of the Royal Statistical Society. Series A (General), 120(2):215–219, ???? 1957. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342825.

Lindley:1957:BRBa

[746] D. V. Lindley. Book review: Scientific Inference, by Harold Jeffreys. Journal of the Royal Statistical Society. Series A (General), 120(2):220–221, ???? 1957. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342826.

Healy:1957:BRB

[747] M. J. R. Healy. Book review: Statistical Methods, by George W. Snedecor. Journal of the Royal Statistical Society. Series A (General), 120(2):221, ???? 1957. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342827.

Whitfield:1957:BRB

[748] J. W. Whitfield. Book review: Nonparametric Statistics for the Behavioral Sciences, by Sidney Siegel. Journal of the Royal Statistical Society. Series A (General), 120(2):221–222, ???? 1957. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342828.

Kendall:1957:BRB

[749] M. G. Kendall. Book review: Variabilita e Concentrazione, by Corrado Gini. Journal of the Royal Statistical Society. Series A (General), 120(2): 222–223, ???? 1957. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342829.

Tocher:1957:BRBc

[750] K. D. Tocher. Book review: Petrographic Modal Analysis, by Felix Chayes. Journal of the Royal Statistical Society. Series A (General), 120 (2):223–224, ???? 1957. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342830.

Vajda:1957:BRBa

[751] S. Vajda. Book review: Wahrscheinlichkeitstheorie, by Hans Richter. Journal of the Royal Statistical Society. Series A (General), 120(2):224, ???? 1957. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342831.

Corlett:1957:BRB

[752] W. J. Corlett. Book review: Statistical Sampling for Auditors and Accountants, by Lawrence L. Vance; John Neter. Journal of the Royal Statistical Society. Series A (General), 120(2): 225, ???? 1957. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342832.

Gore:1957:BRB

[753] A. T. Gore. Book review: Local Expenditure and Exchequer Grants, by D. S. Lees; W. Appleyard; D. A. Hogg; H. Keeling; T. Watson; S. P. Roberts.

Journal of the Royal Statistical Society. Series A (General), 120(2):225– 226, ???? 1957. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342833.

Clarke:1957:BRB

[754] R. D. Clarke. Book review: The Registrar-General’s Decennial Supplement, England and Wales, 1951, Life Tables. Journal of the Royal Statistical Society. Series A (General), 120 (2):226–228, ???? 1957. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342834.

Cox:1957:BRBa

[755] P. R. Cox. Book review: Methods for Population Projection, by Sex and Age. Journal of the Royal Statistical Society. Series A (General), 120(2):228, ???? 1957. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342835.

Thomas:1957:BRB

[756] Brinley Thomas. Book review: Immigrants and Their Children 1850– 1950, by E. P. Hutchinson. Journal of the Royal Statistical Society. Series A (General), 120(2):229, ???? 1957. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342836.

Carrier:1957:BRBa

[757] N. H. Carrier. Book review: European Refugees, 1939–52, by Malcolm J. Proudfoot. Journal of the Royal Statistical Society. Series A (General), 120 (2):229–230, ???? 1957. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342837.

Cox:1957:BRBb

[758] P. R. Cox. Book review: Social Characteristics of Urban and Rural Communities, 1950, by Otis Dudley Duncan; Albert J. Reiss. Journal of the Royal Statistical Society. Series A (General), 120(2):230–231, ???? 1957. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2342838.

Abrams:1957:BRB

[759] Mark Abrams. Book review: Readings in Market Research, by Frederick Edwards. Journal of the Royal Statistical Society. Series A (General), 120(2): 231, ???? 1957. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342839.

Mason:1957:BRB

[760] John Mason. Book review: Modern Market Research: A Guide for Business Executives, by Max K. Adler. Journal of the Royal Statistical Society. Series A (General), 120(2):231– 232, ???? 1957. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342840.

Anonymous:1957:SNb

[761] Anonymous. Statistical notes. Journal of the Royal Statistical Society. Series A (General), 120(2):233–235, ???? 1957. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342841.

Anonymous:1957:CNa

[762] Anonymous. Current note. Journal of the Royal Statistical Society. Series A (General), 120(2):235, ????

1957. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342842.

Allen:1957:PSA

[763] R. G. D. Allen and R. F. George. Professor Sir Arthur Lyon Bowley: November 6th, 1869–January 21st, 1957. Journal of the Royal Statistical Society. Series A (General), 120(2):236– 241, ???? 1957. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342843.

Anonymous:1957:SEAb

[764] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society. Series A (General), 120(2):242–245, ???? 1957. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342844.

Anonymous:1957:LALb

[765] Anonymous. List of additions to the library. Journal of the Royal Statistical Society. Series A (General), 120 (2):246–252, ???? 1957. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342845.

Farrell:1957:MPE

[766] M. J. Farrell. The measurement of productive efficiency. Journal of the Royal Statistical Society. Series A (General), 120(3):253–290, ???? 1957. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2343100.

Merriman:1957:DUE

[767] J. H. H. Merriman, D. T. Caminer, H. L. Barman, D. Evans, and H. Durant. A discussion on the use of electronic data processing equipment. Journal of the Royal Statistical Society. Series A (General), 120(3):291– 307, ???? 1957. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343101.

Land:1957:ALP

[768] A. H. Land. An application of linear programming to the transport of coking coal. Journal of the Royal Statistical Society. Series A (General), 120 (3):308–319, ???? 1957. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343102.

Herdan:1957:MRB

[769] G. Herdan. The mathematical relation between the number of diseases and the number of patients in a community. Journal of the Royal Statistical Society. Series A (General), 120 (3):320–330, ???? 1957. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343103.

Wake:1957:SLD

[770] William C. Wake. Sentence-length distributions of Greek authors. Journal of the Royal Statistical Society. Series A (General), 120(3):331–346, ???? 1957. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343104.

Anonymous:1957:WP

[771] Anonymous. Wholesale prices in 1956. Journal of the Royal Statistical Society. Series A (General), 120(3):347– 351, ???? 1957. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343105.

Bartlett:1957:BRB

[772] M. S. Bartlett. Book review: Statistical Analysis of Stationary Time-Series, by U. Grenander; Murray Rosenblatt. Journal of the Royal Statistical Society. Series A (General), 120(3):352– 353, ???? 1957. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2343106.

Johnson:1957:BRBa

[773] N. L. Johnson. Book review: The Mathematical Theory of Sampling, by Walter A. Hendricks. Journal of the Royal Statistical Society. Series A (General), 120(3):353, ???? 1957. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343107.

Roy:1957:BRB

[774] A. D. Roy. Book review: Mathematical Economics, by R. G. D. Allen. Journal of the Royal Statistical Society. Series A (General), 120(3):353–355, ???? 1957. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343108.

Vajda:1957:BRBb

[775] S. Vajda. Book review: Chapters from the Classical Theory of Probability, by C. Jordan. Journal of the Royal Statistical Society. Series A (General), 120 (3):355, ???? 1957. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2343109.

Hill:1957:BRBa

[776] I. D. Hill. Book review: Statistical Methods in Quality Control, by Dudley J. Cowden. Journal of the Royal Statistical Society. Series A (General), 120(3):355–356, ???? 1957. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2343110.

Tocher:1957:BRBd

[777] K. D. Tocher. Book review: Electronic Data Processing for Business and Industry, by Richard G. Canning. Journal of the Royal Statistical Society. Series A (General), 120(3):356–357, ???? 1957. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343111.

Hill:1957:BRBb

[778] I. D. Hill. Book review: Introduction to Operations Research, by C. West Churchman; Russell L. Ackoff; E. Leonard Arnoff. Journal of the Royal Statistical Society. Series A (General), 120(3):357–358, ???? 1957. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2343112.

Tippett:1957:BRB

[779] L. H. C. Tippett. Book review: A Dictionary of Statistical Terms, 1957, by M. G. Kendall; W. R. Buckland. Journal of the Royal Statistical Society. Series A (General), 120(3):358, ???? 1957. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343113.

Winsten:1957:BRB

[780] C. B. Winsten. Book review: A Revision of Demand Theory, by J. R. Hicks. Journal of the Royal Statistical Society. Series A (General), 120(3):359– 360, ???? 1957. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343114.

Phillips:1957:BRB

[781] R. H. S. Phillips. Book review: Bibliography of Index Numbers. Journal of the Royal Statistical Society. Series A (General), 120(3):360–361, ???? 1957. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343115.

Benjamin:1957:BRB

[782] B. Benjamin. Book review: The Interpretation of Accounts, by F. Sewell Bray. Journal of the Royal Statistical Society. Series A (General), 120(3): 361–362, ???? 1957. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2343116.

Ridley:1957:BRB

[783] T. M. Ridley. Book review: Consumer Instalment Credit, by Board of Governors of the Federal Reserve System; National Bureau of Economic Research; George D. Bailey. Journal of the Royal Statistical Society. Series A (General), 120(3):362–363, ???? 1957. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343117.

Berman:1957:BRB

[784] L. S. Berman. Book review: Problems in the International Comparison of Economic Accounts. Journal of the Royal Statistical Society. Series A (General), 120(3):363–364, ???? 1957. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343118.

Ainsworth:1957:BRBa

[785] R. B. Ainsworth. Book review: The Measurement and Behavior of Unemployment. Journal of the Royal Statistical Society. Series A (General), 120 (3):364–365, ???? 1957. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343119.

Grebenik:1957:BRB

[786] E. Grebenik. Book review: Rural Depopulation in England and Wales: 1851–1951, by John Saville. Journal of the Royal Statistical Society. Series A (General), 120(3):365, ???? 1957. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343120.

Ainsworth:1957:BRBb

[787] R. B. Ainsworth. Book review: Geographical Differentials of Agricultural Wages in the United States, by W. D. Weatherford. Journal of the Royal Statistical Society. Series A (General), 120 (3):365–366, ???? 1957. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343121.

Horsnell:1957:BRB

[788] G. Horsnell. Book review: New Technology and the Supply of Petroleum: The Treatment of Uncertainty in Resources Planning, by Miller B. Spangler. Journal of the Royal Statistical Society. Series A (General), 120(3): 366, ???? 1957. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343122.

Anonymous:1957:SNc

[789] Anonymous. Statistical notes. Journal of the Royal Statistical Society. Series A (General), 120(3):367–368, ???? 1957. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343123.

Anonymous:1957:CNb

[790] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 120(3):368, ???? 1957. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343124.

Anonymous:1957:SEAc

[791] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society. Series A (General), 120(3):369–376, ???? 1957. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343125.

Anonymous:1957:LALc

[792] Anonymous. List of additions to the library. Journal of the Royal Statistical Society. Series A (General), 120 (3):377–383, ???? 1957. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2343126.

Anonymous:1957:PRa

[793] Anonymous. Periodical returns. Journal of the Royal Statistical Society. Series A (General), 120(3):384–385, ???? 1957. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343127.

Anonymous:1957:VIb

[794] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (General), 120(4):532– 534, ???? 1957. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2342846.

Craddock:1957:AST

[795] J. M. Craddock. An analysis of the slower temperature variations at Kew Observatory by means of mutually exclusive band pass filters. Journal of the Royal Statistical Society. Series A (General), 120(4):387–397, ???? 1957. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342847.

Charnock:1957:NSA

[796] H. Charnock. Notes on the specification of atmospheric turbulence. Journal of the Royal Statistical Society. Series A (General), 120(4):398–408, ???? 1957. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342848.

Rushton:1957:SAT

[797] S. Rushton and J. Neumann. Some applications of time series analysis of atmospheric turbulence and oceanography. Journal of the Royal Statistical Society. Series A (General), 120(4): 409–425, ???? 1957. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342849.

Lewis:1957:DP

[798] R. P. Waldo Lewis, G. M. Jenkins, Professor Bartlett, S. K. Zaremba, G. A. Tunnell, R. G. Veryard, P. A. Sheppard, Mr. Craddock, A. M. Walker, M. B. Priestley, H. Wold, M. J. Tucker, and Mr. Charnock. Discussion on the papers. Journal of the Royal Statistical Society. Series A (General), 120 (4):425–439, ???? 1957. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342850.

Sandon:1957:RNA

[799] Frank Sandon. The relative numbers and abilities of some ten-year-old twins. Journal of the Royal Statistical Society. Series A (General), 120(4):440– 450, ???? 1957. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2342851.

Burton:1957:APS

[800] Harry Burton. Alternative presentations of social accounting. Journal of the Royal Statistical Society. Series A (General), 120(4):451–456, ???? 1957. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2342852.

Champernowne:1957:GFA

[801] D. G. Champernowne and J. L. Nicholson. The general form of the addingup criterion: A rejoinder. Journal of the Royal Statistical Society. Series A (General), 120(4):457–458, ???? 1957. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2342853.

Anonymous:1957:RC

[802] Anonymous. Report of the Council. Journal of the Royal Statistical Society. Series A (General), 120(4):459– 478, ???? 1957. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2342854.

Anonymous:1957:POH

[803] Anonymous. Proceedings of the one hundred and twenty-third Annual General Meeting of the Royal Statistical Society held at the London school of hygiene tropical medicine on Wednesday, June 19th. Journal of the Royal Statistical Society. Series A (General), 120(4):479, ???? 1957. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342855.

Lindley:1957:BRBb

[804] D. V. Lindley. Book review: Nonparametric Methods in Statistics, by D. A. S. Fraser. Journal of the Royal Statistical Society. Series A (General), 120 (4):480–481, ???? 1957. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342856.

Durbin:1957:BRB

[805] J. Durbin. Book review: The Lognormal Distribution with Special Reference to its Uses in Economics, by J. Aitchison; J. A. C. Brown. Journal of the Royal Statistical Society. Series A (General), 120(4):481–482, ???? 1957. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342857.

Martin:1957:BRB

[806] Leslie V. Martin. Book review: Probability: An Intermediate Text-Book, by M. T. L. Bizley. Journal of the Royal Statistical Society. Series A (General), 120(4):482–483, ???? 1957. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2342858.

Moore:1957:BRB

[807] P. G. Moore. Book review: Random Processes in Automatic Control, by J. H. Laning; R. H. Battin. Journal of the Royal Statistical Society. Series A (General), 120(4):483–484, ???? 1957. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342859.

Tocher:1957:BRBe

[808] K. D. Tocher. Book review: Contributions on Partially Balanced Incomplete Block Designs with Two Associate Classes, by W. H. Clatworthy. Journal of the Royal Statistical Society. Series A (General), 120(4):484, ???? 1957. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2342860.

Johnson:1957:BRBb

[809] N. L. Johnson. Book review: Proceedings of the First International Congress on the Application of the Theory of Probability in Telephone Engineering and Administration 1955. Journal of the Royal Statistical Society. Series A (General), 120(4):484–485, ???? 1957. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2342861.

Barr:1957:BRB

[810] A. Barr. Book review: Teach Yourself Statistics, by Richard Goodman. Journal of the Royal Statistical Society. Series A (General), 120(4):485, ???? 1957. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2342862.

Barna:1957:BRB

[811] T. Barna. Book review: National Income Statistics: Sources and Methods. Journal of the Royal Statistical Society. Series A (General), 120(4):485– 486, ???? 1957. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2342863.

Deane:1957:BRBa

[812] Marjorie Deane. Book review: The Pattern of Financial Asset Ownership: Wisconsin Individuals, 1949, by Thomas R. Atkinson. Journal of the Royal Statistical Society. Series A (General), 120(4):486–487, ???? 1957. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342864.

Deane:1957:BRBb

[813] Marjorie Deane. Book review: Economic Issues: A Financial and Economic Debate in the Critical Years 1954–57, by Stephen Frowen; H. C. Hillmann. Journal of the Royal Statistical Society. Series A (General), 120 (4):487, ???? 1957. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342865.

Ainsworth:1957:BRBc

[814] R. B. Ainsworth. Book review: The Demand and Supply of Scientific Personnel, by David M. Blank; George J. Stigler. Journal of the Royal Statistical Society. Series A (General), 120 (4):487–488, ???? 1957. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342866.

Clark:1957:BRB

[815] Colin Clark. Book review: The Growth of Public Employment in Great Britain, by Moses Abramovitz; Vera Eliasberg. Journal of the Royal Statistical Society. Series A (General), 120(4):488– 489, ???? 1957. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342867.

Carrier:1957:BRBb

[816] N. H. Carrier. Book review: Family and Neighbourhood, by J. M. Mogey. Journal of the Royal Statistical Society. Series A (General), 120(4):489, ???? 1957. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2342868.

Rutland:1957:BRB

[817] E. H. Rutland. Book review: American Housing and Its Use: The Demand for Shelter Space, by Louis Winnick. Journal of the Royal Statistical Society. Series A (General), 120(4):489–490, ???? 1957. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2342869.

Westergaard:1957:BRB

[818] J. H. Westergaard. Book review: London Travel Survey, 1954. Journal of the Royal Statistical Society. Series A (General), 120(4):490–491, ???? 1957. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2342870.

Deane:1957:BRBc

[819] Marjorie Deane. Book review: A National Survey of Readership of Newspapers and Magazines on a Continuing Basis. Journal of the Royal Statistical Society. Series A (General), 120 (4):491–492, ???? 1957. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2342871.

Menzler:1957:BRB

[820] F. A. A. Menzler. Book review: Company Reports and Their Presentation. Journal of the Royal Statistical Society. Series A (General), 120(4):492, ????

1957. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342872.

Anonymous:1957:SNd

[821] Anonymous. Statistical notes. Journal of the Royal Statistical Society. Series A (General), 120(4):493–495, ???? 1957. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342873.

Anonymous:1957:CNc

[822] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 120(4):495–496, ???? 1957. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342874.

Anonymous:1957:SEAd

[823] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society. Series A (General), 120(4):497–503, ???? 1957. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342875.

Anonymous:1957:LALd

[824] Anonymous. List of additions to the library. Journal of the Royal Statistical Society. Series A (General), 120 (4):504–509, ???? 1957. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342876.

Anonymous:1957:LSP

[825] Anonymous. List of serial publications currently received in the library. Journal of the Royal Statistical Society. Series A (General), 120(4):510–526, ????

1957. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2342877.

Anonymous:1957:PRb

[826] Anonymous. Periodical returns. Journal of the Royal Statistical Society. Series A (General), 120(4):527–531, ???? 1957. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2342878.

Anonymous:1958:VIa

[827] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (General), 121(1):??, ???? 1958. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2342986.

Anonymous:1958:FM

[828] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 121(1):??, ???? 1958. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2342987.

Campion:1958:RDE

[829] Harry Campion. Recent developments in economic statistics. Journal of the Royal Statistical Society. Series A (General), 121(1):1–17, ???? 1958. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2342988.

Cramer:1958:DMM

[830] J. S. Cramer. The depreciation and mortality of motor-cars. Journal of the Royal Statistical Society. Series A (General), 121(1):18–59, ????

1958. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342989.

Klein:1958:BPS

[831] L. R. Klein. The British propensity to save. Journal of the Royal Statistical Society. Series A (General), 121(1):60– 96, ???? 1958. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2342990.

Geary:1958:NCE

[832] R. C. Geary. A note on the comparison of exchange rates and purchasing power between countries. Journal of the Royal Statistical Society. Series A (General), 121(1):97–99, ???? 1958. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342991.

Lah:1958:GYF

[833] Ivo Lah. Generalization of Yastremsky’s formula for analytical graduation of fertility rates. Journal of the Royal Statistical Society. Series A (General), 121(1):100–104, ???? 1958. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2342992.

Wise:1958:BRBa

[834] J. Wise. Book review: The Analysis of Multiple Time Series, by M. H. Quenouille. Journal of the Royal Statistical Society. Series A (General), 121 (1):105–106, ???? 1958. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342993.

Lindley:1958:BRB

[835] D. V. Lindley. Book review: Probability and Scientific Inference, by G. Spencer Brown. Journal of the Royal Statistical Society. Series A (General), 121 (1):106–107, ???? 1958. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342994.

Johnson:1958:BRBa

[836] N. L. Johnson. Book review: Proceedings of the Third Berkeley Symposium on Mathematical Statistics and Probability. Vol. I: The Theory of Statistics; Proceedings of the Third Berkeley Symposium on Mathematical Statistics and Probability. Vol. II: Probability Theory; Proceedings of the Third Berkeley Symposium on Mathematical Statistics and Probability. Vol. III: Astronomy and Physics; Proceedings of the Third Berkeley Symposium on Mathematical Statistics and Probability. Vol IV Biology and Problems of Health; Proceedings of the Third Berkeley Symposium on Mathematical Statistics and Probability. Vol. V: Econometrics Industrial Research and Psychometry. Journal of the Royal Statistical Society. Series A (General), 121(1):107–109, ???? 1958. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342995.

Maurice:1958:BRBa

[837] Rita J. Maurice. Book review: Introduction to Modern Statistics, by Werner Z. Hirsch. Journal of the Royal Statistical Society. Series A (General), 121(1):109–110, ???? 1958. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2342996.

Wise:1958:BRBb

[838] M. E. Wise. Book review: Mathematische Statistik, by B. L. van der Waerden. Journal of the Royal Statistical Society. Series A (General), 121 (1):110–111, ???? 1958. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342997.

Smith:1958:BRBa

[839] Cedric A. B. Smith. Book review: Mathematics and Statistics for use in Pharmacy, Biology and Chemistry, by L. Saunders; R. Fleming. Journal of the Royal Statistical Society. Series A (General), 121(1):112, ???? 1958. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342998.

Smith:1958:BRBb

[840] Cedric A. B. Smith. Book review: The Fascination of Numbers, by W. J. Reichmann. Journal of the Royal Statistical Society. Series A (General), 121 (1):112, ???? 1958. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342999.

Lipton:1958:BRB

[841] S. Lipton. Book review: Digital Computer Programming, by D. D. Mc-Cracken. Journal of the Royal Statistical Society. Series A (General), 121 (1):113, ???? 1958. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343000.

Hill:1958:BRBa

[842] I. D. Hill. Book review: Regression Analysis of Production Costs and Factory Operations, by Philip Lyle; Leonard H. C. Tippett. Journal of the Royal Statistical Society. Series A (General), 121(1):113–114, ????

1958. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343001.

Martin:1958:BRBa

[843] W. J. Martin. Book review: Annual Epidemiological and Vital Statistics, 1954, by World Health Organization. Journal of the Royal Statistical Society. Series A (General), 121 (1):114–115, ???? 1958. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2343002.

Barr:1958:BRBa

[844] A. Barr. Book review: Social Aspects of Prescribing, by J. P. Martin. Journal of the Royal Statistical Society. Series A (General), 121(1):115–116, ???? 1958. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343003.

Benjamin:1958:BRBa

[845] B. Benjamin. Book review: Population Theories and the Economic Interpretation, by Sydney H. Coontz. Journal of the Royal Statistical Society. Series A (General), 121(1):116–117, ???? 1958. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343004.

Carrier:1958:BRBa

[846] N. H. Carrier. Book review: American Families, by Paul C. Glick. Journal of the Royal Statistical Society. Series A (General), 121(1):117–119, ???? 1958. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343005.

Nicholson:1958:BRB

[847] J. L. Nicholson. Book review: The Analysis of Family Budgets, by S. J. Prais; H. S. Houthakker. Journal of the Royal Statistical Society. Series A (General), 121(1):119–123, ???? 1958. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343006.

Berman:1958:BRB

[848] L. S. Berman. Book review: Regional Income, by Conference on Research in Income and Wealth. Journal of the Royal Statistical Society. Series A (General), 121(1):123–124, ???? 1958. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343007.

Champernowne:1958:BRB

[849] D. G. Champernowne. Book review: A Theory of the Consumption Function, by Milton Friedman. Journal of the Royal Statistical Society. Series A (General), 121(1):124–126, ???? 1958. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343008.

Searle:1958:BRB

[850] W. F. Searle. Book review: The Investigation of National Income in British Tropical Dependencies, by A. R. Prest. Journal of the Royal Statistical Society. Series A (General), 121(1):126– 127, ???? 1958. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343009.

Barr:1958:BRBb

[851] A. Barr. Book review: An Economic Survey of Northern Ireland, by K. S.

Isles; N. Cuthbert. Journal of the Royal Statistical Society. Series A (General), 121(1):127–128, ???? 1958. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2343010.

Maurice:1958:BRBb

[852] Rita Maurice. Book review: Elementary Business and Economic Statistics, by Alva M. Tuttle. Journal of the Royal Statistical Society. Series A (General), 121(1):128–129, ???? 1958. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2343011.

Lloyd:1958:BRB

[853] E. H. Lloyd. Book review: Analysis for Production Management, by E. H. Bowman; R. B. Fetter. Journal of the Royal Statistical Society. Series A (General), 121(1):129–130, ???? 1958. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343012.

Anonymous:1958:SNa

[854] Anonymous. Statistical notes. Journal of the Royal Statistical Society. Series A (General), 121(1):131–133, ???? 1958. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343013.

Anonymous:1958:CNa

[855] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 121(1):133, ???? 1958. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343014.

Snow:1958:RGG

[856] E. C. Snow. Roy Goncalves Glenday, 1889–1957. Journal of the Royal Statistical Society. Series A (General), 121 (1):134, ???? 1958. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343015.

Anonymous:1958:SEAa

[857] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society. Series A (General), 121(1):135–146, ???? 1958. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343016.

Anonymous:1958:LALa

[858] Anonymous. List of additions to the library. Journal of the Royal Statistical Society. Series A (General), 121 (1):147–156, ???? 1958. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343017.

Anonymous:1958:PRa

[859] Anonymous. Periodical returns. Journal of the Royal Statistical Society. Series A (General), 121(1):157–159, ???? 1958. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343018.

Anonymous:1958:BM

[860] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (General), 121(1):??, ???? 1958. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/4290495.

Brichler:1958:CPS

[861] M. Brichler. Classification of the population by social and economic characteristics: The French experience and international recommendations. Journal of the Royal Statistical Society. Series A (General), 121(2):161–195, ???? 1958. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343359.

Kendall:1958:RS

[862] M. G. Kendall. Research in statistics. Journal of the Royal Statistical Society. Series A (General), 121(2):196– 215, ???? 1958. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343360.

Kopanitsas:1958:DPS

[863] D. E. Kopanitsas. Differential prices for the services of ships belonging to the same general class — a case of oligopoly. Journal of the Royal Statistical Society. Series A (General), 121 (2):216–222, ???? 1958. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343361.

Kosambi:1958:ERC

[864] D. D. Kosambi and U. V. Ramamohana Rao. The efficiency of randomization by card shuffling. Journal of the Royal Statistical Society. Series A (General), 121(2):223–233, ???? 1958. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343362. See note [994].

Smith:1958:BRBc

[865] Cedric A. B. Smith. Book review: Statistical Theory: The Relationship of Probability, Credibility and Error, by Lancelot Hogben. Journal of the Royal Statistical Society. Series A (General), 121(2):234–236, ???? 1958. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2343363.

Bartlett:1958:BRB

[866] M. S. Bartlett. Book review: The Mathematical Theory of Epidemics, by N. T. J. Bailey. Journal of the Royal Statistical Society. Series A (General), 121(2):236–237, ???? 1958. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2343364.

Cox:1958:BRBa

[867] D. R. Cox. Book review: Experimental Designs, by W. G. Cochran; G. M. Cox. Journal of the Royal Statistical Society. Series A (General), 121 (2):237–238, ???? 1958. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343365.

Smith:1958:BRBd

[868] Cedric A. B. Smith. Book review: Techniques of Mathematical Analysis, by C. J. Tranter. Journal of the Royal Statistical Society. Series A (General), 121(2):238, ???? 1958. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343366.

Good:1958:BRB

[869] I. J. Good. Book review: Probability, Statistics and Truth, by Richard von Mises; Hilda Geiringer. Journal of the Royal Statistical Society. Series A (General), 121(2):238–240, ???? 1958. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343367.

Tocher:1958:BRBa

[870] K. D. Tocher. Book review: Statistical Tables for Biological, Agricultural and Medical Research, by R. A. Fisher; F. Yates. Journal of the Royal Statistical Society. Series A (General), 121 (2):240, ???? 1958. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343368.

Tocher:1958:BRBb

[871] K. D. Tocher. Book review: Fractional Factorial Experiment Designs for Factors at Two Levels, by The Statistical Engineering Laboratory. Journal of the Royal Statistical Society. Series A (General), 121(2):240–241, ???? 1958. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343369.

Tocher:1958:BRBc

[872] K. D. Tocher. Book review: Tables of the Non-Central t-Distribution, by George J. Resnikoff; Gerald J. Lieberman. Journal of the Royal Statistical Society. Series A (General), 121(2): 241–242, ???? 1958. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343370.

Benjamin:1958:BRBb

[873] B. Benjamin. Book review: The Measurement of Levels of Living with Special Reference to Jamaica, by C. A. Moser. Journal of the Royal Statistical Society. Series A (General), 121 (2):242, ???? 1958. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343371.

Utting:1958:BRB

[874] J. E. G. Utting. Book review: Report of an Enquiry into Household Expenditure in Great Britain. Journal of the Royal Statistical Society. Series A (General), 121(2):242–245, ???? 1958. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2343372.

Carrier:1958:BRBb

[875] N. H. Carrier. Book review: The Registrar-General’s Statistical Review of England and Wales for the Year 1955, Part III Commentary. Journal of the Royal Statistical Society. Series A (General), 121(2):245–250, ???? 1958. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343373.

Barr:1958:BRBc

[876] A. Barr. Book review: Tuberculosis Statistics for England and Wales, 1938–1955: An Analysis of Trends and Geographical Distribution, by W. P. D. Logan; B. Benjamin. Journal of the Royal Statistical Society. Series A (General), 121(2):250–251, ???? 1958. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343374.

Benjamin:1958:BRBc

[877] B. Benjamin. Book review: Chronic Bronchitis in Newcastle-Upon-Tyne, by A. G. Ogilvie; D. J. Newell. Journal of the Royal Statistical Society. Series A (General), 121(2):251, ???? 1958. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343375.

Heady:1958:BRB

[878] J. A. Heady. Book review: Morbidity and Mortality in the First Year of Life, by Fred Grundy; E. Lewis-Faning. Journal of the Royal Statistical Society. Series A (General), 121(2):251– 253, ???? 1958. CODEN JSSAEF.

ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2343376.

Cox:1958:BRBb

[879] P. R. Cox. Book review: The Aging of Populations and Its Economic and Social Implications. Journal of the Royal Statistical Society. Series A (General), 121(2):253–254, ???? 1958. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2343377.

Carrier:1958:BRBc

[880] N. H. Carrier. Book review: The Demography of Ceylon, by N. K. Sarkar. Journal of the Royal Statistical Society. Series A (General), 121(2):254– 255, ???? 1958. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2343378.

Anonymous:1958:SNb

[881] Anonymous. Statistical notes. Journal of the Royal Statistical Society. Series A (General), 121(2):256–258, ???? 1958. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343379.

Anonymous:1958:SEAb

[882] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society. Series A (General), 121(2):259–263, ???? 1958. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343380.

Anonymous:1958:LALb

[883] Anonymous. List of additions to the library. Journal of the Royal Statistical Society. Series A (General), 121 (2):264–268, ???? 1958. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343381.

Shone:1958:IPS

[884] Robert Shone and H. R. Fisher. Industrial production and steel consumption. Journal of the Royal Statistical Society. Series A (General), 121(3): 269–311, ???? 1958. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342902.

Prais:1958:SPM

[885] S. J. Prais. Some problems in the measurement of price changes with special reference to the cost of living. Journal of the Royal Statistical Society. Series A (General), 121(3):312–332, ???? 1958. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342903.

Ashford:1958:DLT

[886] J. R. Ashford. The design of a longterm sampling programme to measure the hazard associated with an industrial environment. Journal of the Royal Statistical Society. Series A (General), 121(3):333–347, ???? 1958. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2342904.

Anonymous:1958:WP

[887] Anonymous. Wholesale prices in 1957. Journal of the Royal Statistical Society. Series A (General), 121(3):348– 352, ???? 1958. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342905.

Vajda:1958:BRB

[888] S. Vajda. Book review: Games and Decisions: Introduction and Critical Survey, by R. Duncan Luce; Howard Raiffa. Journal of the Royal Statistical Society. Series A (General), 121 (3):353–354, ???? 1958. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342906.

Cox:1958:BRBc

[889] D. R. Cox. Book review: An Introduction to Probability Theory and Its Applications, by W. Feller. Journal of the Royal Statistical Society. Series A (General), 121(3):354–355, ???? 1958. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342907.

Johnson:1958:BRBb

[890] N. L. Johnson. Book review: Introduction to Statistical Reasoning, by Philip J. McCarthy. Journal of the Royal Statistical Society. Series A (General), 121 (3):355, ???? 1958. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342908.

Cox:1958:BRBd

[891] D. R. Cox. Book review: Queues, Inventories and Maintenance, by P. M. Morse. Journal of the Royal Statistical Society. Series A (General), 121 (3):355–356, ???? 1958. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342909.

Johnson:1958:BRBc

[892] N. L. Johnson. Book review: Combination of Observations, by W. M. Smart. Journal of the Royal Statistical Society. Series A (General), 121 (3):356–357, ???? 1958. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342910.

Creasy:1958:BRB

[893] M. A. Creasy. Book review: Experimental Design in Psychology and the Medical Sciences, by A. E. Maxwell. Journal of the Royal Statistical Society. Series A (General), 121(3):357– 358, ???? 1958. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342911.

Smith:1958:BRBe

[894] Cedric A. B. Smith. Book review: An Introduction to Genetic Statistics, by Oscar Kempthorne. Journal of the Royal Statistical Society. Series A (General), 121(3):358–359, ???? 1958. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342912.

Hill:1958:BRBb

[895] I. D. Hill. Book review: Statistical Exercises, Part II, by N. L. Johnson. Journal of the Royal Statistical Society. Series A (General), 121(3):359– 360, ???? 1958. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342913.

Gray:1958:BRB

[896] P. G. Gray. Book review: Survey Methods in Social Investigation, by C. A. Moser. Journal of the Royal Statistical Society. Series A (General), 121 (3):360–361, ???? 1958. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342914.

Cottrell:1958:BRB

[897] H. C. Cottrell. Book review: Economic Models: An Exposition, by E. F. Beach. Journal of the Royal Statistical Society. Series A (General), 121 (3):361–362, ???? 1958. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2342915.

Deane:1958:BRBa

[898] Marjorie Deane. Book review: Economic Concentration and the Monopoly Problem, by Edward S. Mason. Journal of the Royal Statistical Society. Series A (General), 121(3):362–363, ???? 1958. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2342916.

Roy:1958:BRB

[899] A. D. Roy. Book review: Time in Economics, by G. L. S. Shackle. Journal of the Royal Statistical Society. Series A (General), 121(3):363–364, ???? 1958. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2342917.

Cox:1958:BRBe

[900] P. R. Cox. Book review: Economic Backwardness and Economic Growth, by Harvey Leibenstein. Journal of the Royal Statistical Society. Series A (General), 121(3):364, ???? 1958. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2342918.

Martin:1958:BRBb

[901] W. J. Martin. Book review: Measurement of Levels of Health. Journal of the Royal Statistical Society. Series A (General), 121(3):365, ????

1958. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342919.

Lewis-Faning:1958:BRB

[902] E. Lewis-Faning. Book review: The Survey of Sickness 1943 to 1952, by W. P. D. Logan; Eileen M. Brooke. Journal of the Royal Statistical Society. Series A (General), 121(3):365–367, ???? 1958. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342920.

Barr:1958:BRBd

[903] A. Barr. Book review: The Bethlem Royal Hospital and The Maudsley Hospital. Triennial Statistical Report, 1952–54, by C. P. Blacker. Journal of the Royal Statistical Society. Series A (General), 121(3):367, ???? 1958. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342921.

Martin:1958:BRBc

[904] W. J. Martin. Book review: Registrar General’s Decennial Supplement, England and Wales, 1951, Area Mortality. Journal of the Royal Statistical Society. Series A (General), 121(3): 367–368, ???? 1958. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342922.

Martin:1958:BRBd

[905] W. J. Martin. Book review: Registrar General’s Decennial Supplement, England and Wales, 1951, Occupational Mortality. Journal of the Royal Statistical Society. Series A (General), 121 (3):368–369, ???? 1958. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342923.

Parkinson:1958:BRB

[906] J. R. Parkinson. Book review: Shipbuilding in Britain, Mainly between the Two World Wars, by L. Jones Cardiff. Journal of the Royal Statistical Society. Series A (General), 121(3):369– 370, ???? 1958. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2342924.

Anonymous:1958:SNc

[907] Anonymous. Statistical notes. Journal of the Royal Statistical Society. Series A (General), 121(3):371–373, ???? 1958. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2342925.

Anonymous:1958:SEAc

[908] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society. Series A (General), 121(3):374–381, ???? 1958. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2342926.

Anonymous:1958:LALc

[909] Anonymous. List of additions to the library. Journal of the Royal Statistical Society. Series A (General), 121 (3):382–389, ???? 1958. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2342927.

Anonymous:1958:VIb

[910] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (General), 121(4):519– 521, ???? 1958. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2343309.

Murray:1958:DUG

[911] Keith Murray. The development of the universities in Great Britain. Journal of the Royal Statistical Society. Series A (General), 121(4):391–419, ???? 1958. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343310.

Chandler:1958:ETM

[912] K. N. Chandler and J. C. Tanner. Estimates of the total miles run by road vehicles in Great Britain in 1952 and 1956. Journal of the Royal Statistical Society. Series A (General), 121(4): 420–437, ???? 1958. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343311.

Gales:1958:SF

[913] Kathleen Gales. A survey of Fellows — I. Journal of the Royal Statistical Society. Series A (General), 121(4):438– 457, ???? 1958. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343312.

Anonymous:1958:RC

[914] Anonymous. Report of the Council. Journal of the Royal Statistical Society. Series A (General), 121(4):458– 478, ???? 1958. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343313.

Anonymous:1958:POH

[915] Anonymous. Proceedings of the one hundred and twenty-fourth Annual General Meeting of the Royal Statistical Society held at the London School of Hygiene and Tropical Medicine on Wednesday, June 18th, 1958, at 5.30.

Journal of the Royal Statistical Society. Series A (General), 121(4):m, ???? 1958. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343314.

Barton:1958:BRB

[916] D. E. Barton. Book review: A Course in Multivariate Analysis, by M. G. Kendall. Journal of the Royal Statistical Society. Series A (General), 121 (4):480–481, ???? 1958. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343315.

Johnson:1958:BRBd

[917] N. L. Johnson. Book review: Some Aspects of Multivariate Analysis, by S. N. Roy. Journal of the Royal Statistical Society. Series A (General), 121 (4):481–482, ???? 1958. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343316.

Johnson:1958:BRBe

[918] N. L. Johnson. Book review: Introduction to Multivariate Statistical Analysis, by T. W. Anderson. Journal of the Royal Statistical Society. Series A (General), 121(4):482–483, ???? 1958. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343317.

Wise:1958:BRBc

[919] J. Wise. Book review: Readings in Linear Programming, by S. Vajda. Journal of the Royal Statistical Society. Series A (General), 121(4):483–484, ???? 1958. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343318.

Land:1958:BRB

[920] A. H. Land. Book review: Programme Linéaire — Agrégation et Nombres Indices, by R. Roy. Journal of the Royal Statistical Society. Series A (General), 121(4):484, ???? 1958. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343319.

Brookes:1958:BRB

[921] B. C. Brookes. Book review: A First Course in Statistics, by Robert Loveday. Journal of the Royal Statistical Society. Series A (General), 121(4):484– 485, ???? 1958. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343320.

Jenkins:1958:BRB

[922] G. M. Jenkins. Book review: Numerical Methods, by R. A. Buckingham. Journal of the Royal Statistical Society. Series A (General), 121(4):485–486, ???? 1958. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343321.

Benjamin:1958:BRBd

[923] B. Benjamin. Book review: Programming for an Automatic Digital Calculator, by K. H. V. Booth. Journal of the Royal Statistical Society. Series A (General), 121(4):486–487, ???? 1958. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343322.

Haycocks:1958:BRB

[924] H. W. Haycocks. Book review: The Investment Decision: An Empirical Study, by John R. Meyer; Edwin Kuh. Journal of the Royal Statistical Society.

Series A (General), 121(4):487, ???? 1958. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343323.

Deane:1958:BRBb

[925] Marjorie Deane. Book review: Women’s Measurements and Sizes. Journal of the Royal Statistical Society. Series A (General), 121(4):487– 488, ???? 1958. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343324.

Stein:1958:BRB

[926] L. Stein. Book review: A Study of the Major Psychoses in an English County, by Michael Shepherd. Journal of the Royal Statistical Society. Series A (General), 121(4):488–489, ???? 1958. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343325.

Barr:1958:BRBe

[927] A. Barr. Book review: Cancer Statistics for England and Wales, 1901– 1955, A Summary of Data Relating to Mortality and Morbidity, by Alan McKenzie. Journal of the Royal Statistical Society. Series A (General), 121 (4):490, ???? 1958. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343326.

Corlett:1958:BRB

[928] W. J. Corlett. Book review: A Survey of Social Conditions in England and Wales as Illustrated, by Statistics. by A. M. Carr-Saunders; D. Caradog Jones; C. A. Moser. Journal of the Royal Statistical Society. Series A (General), 121(4):491–492, ????

1958. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343327.

Boreham:1958:BRB

[929] A. J. Boreham. Book review: The Economics of International Migration, by Brinley Thomas. Journal of the Royal Statistical Society. Series A (General), 121(4):492–493, ???? 1958. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2343328.

Deane:1958:BRBc

[930] Marjorie Deane. Book review: The Control of Inflation, by J. E. Meade. Journal of the Royal Statistical Society. Series A (General), 121(4):493, ???? 1958. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343329.

Mars:1958:BRB

[931] John Mars. Book review: Economic Analysis and Policy in Underdeveloped Countries, by P. T. Bauer. Journal of the Royal Statistical Society. Series A (General), 121(4):493–496, ???? 1958. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343330.

Anonymous:1958:SNd

[932] Anonymous. Statistical notes. Journal of the Royal Statistical Society. Series A (General), 121(4):497–499, ???? 1958. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343331.

Anonymous:1958:CNb

[933] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 121(4):499–500, ????

1958. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343332.

Anonymous:1958:SEAd

[934] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society. Series A (General), 121(4):501–507, ???? 1958. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343333.

Anonymous:1958:LALd

[935] Anonymous. List of additions to the library. Journal of the Royal Statistical Society. Series A (General), 121 (4):508–513, ???? 1958. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2343334.

Anonymous:1958:PRb

[936] Anonymous. Periodical returns. Journal of the Royal Statistical Society. Series A (General), 121(4):514–518, ???? 1958. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343335.

Anonymous:1959:VIa

[937] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (General), 122(1):??, ???? 1959. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343048.

Anonymous:1959:FM

[938] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (General), 122(1):??, ???? 1959. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343049.

Lydall:1959:LTT

[939] H. F. Lydall. The long-term trend in the size distribution of income. Journal of the Royal Statistical Society. Series A (General), 122(1):1–46, ???? 1959. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343050.

Pearson:1959:RCS

[940] E. S. Pearson. Report of the committee on the supply of and demand for statisticians. Journal of the Royal Statistical Society. Series A (General), 122(1):47– 76, ???? 1959. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2343051.

Florence:1959:SCO

[941] P. Sargant Florence. Size of company and other factors in dividend policy. Journal of the Royal Statistical Society. Series A (General), 122(1):77–98, ???? 1959. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2343052.

Welch:1959:BRB

[942] B. L. Welch. Book review: The Advanced Theory of Statistics: Vol. I– Distribution Theory, by M. G. Kendall; A. Stuart. Journal of the Royal Statistical Society. Series A (General), 122 (1):99–100, ???? 1959. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343053.

Lindley:1959:BRB

[943] D. V. Lindley. Book review: Fundamentals of Statistical Reasoning, by M. H. Quenouille. Journal of the Royal Statistical Society. Series A (General), 122(1):100–101, ???? 1959. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2343054.

Barton:1959:BRB

[944] D. E. Barton. Book review: Introduction to Combinatorial Analysis, by John Riordan. Journal of the Royal Statistical Society. Series A (General), 122(1):102–103, ???? 1959. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2343055.

Ehrenberg:1959:BRB

[945] A. S. C. Ehrenberg. Book review: Multivariate Correlational Analysis, by Philip H. DuBois. Journal of the Royal Statistical Society. Series A (General), 122(1):103, ???? 1959. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2343056.

Maurice:1959:BRBa

[946] Rita J. Maurice. Book review: Applied Statistics for Economists: A Course in Statistical Methods, by P. H. Karmel. Journal of the Royal Statistical Society. Series A (General), 122(1):103– 104, ???? 1959. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2343057.

Cox:1959:BRBa

[947] D. R. Cox. Book review: Experimental Designs in Industry, by V. Chew. Journal of the Royal Statistical Society. Series A (General), 122(1):104–105, ???? 1959. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343058.

Vajda:1959:BRBa

[948] S. Vajda. Book review: Studies in the Mathematical Theory of Inventory and Production, by K. J. Arrow; S. Karlin; H. Scarf. Journal of the Royal Statistical Society. Series A (General), 122 (1):106–107, ???? 1959. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343059.

Gray:1959:BRB

[949] P. G. Gray. Book review: Sampling Opionions: An Analysis of Survey Procedure, by Frederick F. Stephan; Philip J. McCarthy. Journal of the Royal Statistical Society. Series A (General), 122 (1):107, ???? 1959. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343060.

Tocher:1959:BRBa

[950] K. D. Tocher. Book review: Further Contributions to the Solution of Simultaneous Linear Equations and the Determination of Eigenvalues. Journal of the Royal Statistical Society. Series A (General), 122(1):107–108, ???? 1959. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343061.

Tocher:1959:BRBb

[951] K. D. Tocher. Book review: Integrals of Airy Functions. Journal of the Royal Statistical Society. Series A (General), 122(1):108–109, ???? 1959. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2343062.

Tocher:1959:BRBc

[952] K. D. Tocher. Book review: Table of Natural Logarithms for Arguments between Five and Ten to Sixteen Decimal Places. Journal of the Royal Statistical Society. Series A (General), 122 (1):109, ???? 1959. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343063.

Benson:1959:BRBa

[953] F. Benson. Book review: Finite Queueing Tables, by L. G. Peck; R. N. Hazelwood. Journal of the Royal Statistical Society. Series A (General), 122(1): 109–110, ???? 1959. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343064. See cor-

rigendum [1043].

Benson:1959:BRBb

[954] F. Benson. Book review: A Comprehensive Bibliography on Operations Research, through 1956 with Supplement for 1957. Journal of the Royal Statistical Society. Series A (General), 122 (1):110–111, ???? 1959. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343065.

Thomas:1959:BRBa

[955] Brinley Thomas. Book review: Selected Studies of Migration since World War II. Journal of the Royal Statistical Society. Series A (General), 122(1):111– 112, ???? 1959. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343066.

Cox:1959:BRBb

[956] P. R. Cox. Book review: The Older Population of the United States, by Henry D. Sheldon; Clark Tibbitts. Journal of the Royal Statistical Society. Series A (General), 122(1):112– 113, ???? 1959. CODEN JSSAEF.

ISSN 0035-9238. URL http://www.

jstor.org/stable/2343067.

Haley:1959:BRB

[957] K. B. Haley. Book review: The Efficiency of the Coal Industry, by James M. Henderson. Journal of the Royal Statistical Society. Series A (General), 122(1):113–114, ???? 1959. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2343068.

Anonymous:1959:SNa

[958] Anonymous. Statistical notes. Journal of the Royal Statistical Society. Series A (General), 122(1):115–117, ???? 1959. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343069.

Anonymous:1959:CNa

[959] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 122(1):118, ???? 1959. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343070.

Anonymous:1959:SEAa

[960] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society. Series A (General), 122(1):119–130, ???? 1959. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343071.

Anonymous:1959:LALa

[961] Anonymous. List of additions to the library. Journal of the Royal Statistical Society. Series A (General), 122 (1):131–141, ???? 1959. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343072.

Anonymous:1959:PRa

[962] Anonymous. Periodical returns. Journal of the Royal Statistical Society. Series A (General), 122(1):142–143, ???? 1959. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2343073.

Dicks-Mireaux:1959:DWI

[963] L. A. Dicks-Mireaux and J. C. R. Dow. The determinants of wage inflation: United Kingdom, 1946–56. Journal of the Royal Statistical Society. Series A (General), 122(2):145–184, ???? 1959. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2342602.

Lomax:1959:PPM

[964] K. S. Lomax. Production and productivity movements in the United Kingdom since 1900. Journal of the Royal Statistical Society. Series A (General), 122(2):185–220, ???? 1959. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2342603.

Hollingsworth:1959:UEC

[965] T. H. Hollingsworth. Using an electronic computer in a problem of medical diagnosis. Journal of the Royal Statistical Society. Series A (General), 122 (2):221–231, ???? 1959. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2342604.

Bartholomew:1959:NMP

[966] D. J. Bartholomew. Note on the measurement and prediction of labour turnover. Journal of the Royal Statistical Society. Series A (General), 122 (2):232–239, ???? 1959. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342605.

Anonymous:1959:PSG

[967] Anonymous. Proceedings of a special general meeting of the Royal Statistical Society held at the London School of Hygiene and Tropical Medicine on Thursday, December 18th, 1958, immediately on the conclusion of the ordinary meeting convened for the same day. Journal of the Royal Statistical Society. Series A (General), 122(2):240, ???? 1959. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2342606.

Armitage:1959:BRB

[968] P. Armitage. Book review: Planning of Experiments, by D. R. Cox. Journal of the Royal Statistical Society. Series A (General), 122(2):241–242, ???? 1959. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342607.

Brookes:1959:BRB

[969] B. C. Brookes. Book review: Principles of Statistical Techniques: A First Course from the Beginnings for Schools and Universities, by P. G. Moore. Journal of the Royal Statistical Society. Series A (General), 122(2):242–243, ???? 1959. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342608.

James:1959:BRB

[970] G. S. James. Book review: Statistics of Extremes, by E. J. Gumbel. Journal of the Royal Statistical Society. Series A (General), 122(2):243–245, ???? 1959. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2342609.

Creasy:1959:BRB

[971] Monica A. Creasy. Book review: Measurement and Statistics, by Virginia L. Senders. Journal of the Royal Statistical Society. Series A (General), 122 (2):245–246, ???? 1959. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2342610.

Durbin:1959:BRB

[972] J. Durbin. Book review: Sampling in Sweden: Contributions to the Methods and Theories of Sample Survey Practice, by Tore Dalenius. Journal of the Royal Statistical Society. Series A (General), 122(2):246–248, ???? 1959. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2342611.

Maurice:1959:BRBb

[973] Rita J. Maurice. Book review: Economic Arithmetic, by Robin Marris. Journal of the Royal Statistical Society. Series A (General), 122(2):248– 249, ???? 1959. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2342612.

Duckworth:1959:BRB

[974] W. E. Duckworth. Book review: Linear Programming: Methods and Application, by Saul I. Gass. Journal of the Royal Statistical Society. Series A (General), 122(2):249–250, ???? 1959. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2342613.

Vajda:1959:BRBb

[975] S. Vajda. Book review: Allocation in Space. Production, Transport and Industrial Location, by Louis Lefeber. Journal of the Royal Statistical Society. Series A (General), 122(2):250– 251, ???? 1959. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342614.

Cox:1959:BRBc

[976] P. R. Cox. Book review: Elements of Vital Statistics, by B. Benjamin. Journal of the Royal Statistical Society. Series A (General), 122(2):251–252, ???? 1959. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342615.

Martin:1959:BRBa

[977] W. J. Martin. Book review: Morbidity Statistics from General Practice, by W. P. D. Logan; A. A. Cushion. Journal of the Royal Statistical Society. Series A (General), 122(2):252–253, ???? 1959. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342616.

Benjamin:1959:BRBa

[978] B. Benjamin. Book review: Population Problem in India, by P. K. Wattal. Journal of the Royal Statistical Society. Series A (General), 122(2):253– 254, ???? 1959. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342617.

Lodge:1959:BRB

[979] T. S. Lodge. Book review: Statistics Essential for Police Efficiency, by John I. Griffin. Journal of the Royal Statistical Society. Series A (General), 122 (2):254, ???? 1959. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2342618.

Paine:1959:BRB

[980] G. Paine. Book review: An Appraisal of the 1950 Census Income Data. Journal of the Royal Statistical Society. Series A (General), 122(2):254–255, ???? 1959. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2342619.

Bransby:1959:BRB

[981] E. R. Bransby. Book review: Social Economics, by Walter Hagenbuch. Journal of the Royal Statistical Society. Series A (General), 122(2):255– 257, ???? 1959. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2342620.

Ainsworth:1959:BRB

[982] R. B. Ainsworth. Book review: The Labour Force under Changing Income and Employment, by Clarence D. Long. Journal of the Royal Statistical Society. Series A (General), 122(2):257– 258, ???? 1959. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2342621.

Carrier:1959:BRB

[983] N. H. Carrier. Book review: Economic Characteristics of International Migrants: Statistics for Selected Countries 1918–1954. Journal of the Royal Statistical Society. Series A (General), 122(2):258, ???? 1959. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2342622.

Cottrell:1959:BRB

[984] H. C. Cottrell. Book review: Corporate Bond Quality and Investor Experience, by W. Braddock Hickman. Journal of the Royal Statistical Society. Series A (General), 122(2):258–259, ???? 1959. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342623.

Anonymous:1959:SNb

[985] Anonymous. Statistical notes. Journal of the Royal Statistical Society. Series A (General), 122(2):260–262, ???? 1959. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342624.

Champernowne:1959:FB

[986] D. G. Champernowne. F. A. Burchardt 1902–1958. Journal of the Royal Statistical Society. Series A (General), 122 (2):263, ???? 1959. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342625.

Champernowne:1959:ACP

[987] D. G. Champernowne. Arthur Cecil Pigou 1877–1959. Journal of the Royal Statistical Society. Series A (General), 122(2):263–265, ???? 1959. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2342626.

Anonymous:1959:SEAb

[988] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society. Series A (General), 122(2):266–272, ???? 1959. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342627.

Anonymous:1959:LALb

[989] Anonymous. List of additions to the library. Journal of the Royal Statistical Society. Series A (General), 122 (2):273–278, ???? 1959. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2342628.

Armitage:1959:CSC

[990] P. Armitage. The comparison of survival curves. Journal of the Royal Statistical Society. Series A (General), 122 (3):279–300, ???? 1959. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2342794.

Thomson:1959:EMS

[991] W. E. Thomson. Ernie — a mathematical and statistical analysis. Journal of the Royal Statistical Society. Series A (General), 122(3):301–333, ???? 1959. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2342795.

Cramer:1959:PMD

[992] J. S. Cramer. Private motoring and the demand for petrol. Journal of the Royal Statistical Society. Series A (General), 122(3):334–347, ???? 1959. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2342796.

Gales:1959:SFI

[993] Kathleen Gales. A survey of Fellows– II. Journal of the Royal Statistical Society. Series A (General), 122(3):348– 372, ???? 1959. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2342797.

Joseph:1959:NPD

[994] A. W. Joseph. A note on the paper by D. D. Kosambi and U. V. Ramamohana Rao on “The Efficiency of Randomization by Card Shuffling”. Journal of the Royal Statistical Society. Series A (General), 122(3):373–374, ???? 1959. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342798. See [864].

Anonymous:1959:WP

[995] Anonymous. Wholesale prices in 1958. Journal of the Royal Statistical Society. Series A (General), 122(3):375– 379, ???? 1959. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342799.

Mallows:1959:BRB

[996] C. L. Mallows. Book review: Information Theory and Statistics, by Solomon Kullback. Journal of the Royal Statistical Society. Series A (General), 122 (3):380–381, ???? 1959. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342800.

Beale:1959:BRB

[997] E. M. L. Beale. Book review: Studies in Linear and Non-Linear Programming, by Kenneth J. Arrow; Leonid Hurwicz; Hirofumi Uzawa; Hollis B. Chenery; Selmer M. Johnson; Samuel Karlin; Thomas Marschak; Robert M. Solow. Journal of the Royal Statistical Society. Series A (General), 122(3): 381–382, ???? 1959. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2342801.

Tocher:1959:BRBd

[998] K. D. Tocher. Book review: Operations Research for Industrial Management, by Dimitris N. Chorafas. Journal of the Royal Statistical Society. Series A (General), 122(3):382–384, ???? 1959. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2342802.

Tocher:1959:BRBe

[999] K. D. Tocher. Book review: Sampled-Data Control Systems, by John R. Ragazzini; Gene F. Franklin. Journal of the Royal Statistical Society. Series A (General), 122(3):384–385, ???? 1959. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2342803.

Hill:1959:BRB

[1000] I. D. Hill. Book review: Statistical Quality Control: An Introduction for Management, by D. H. W. Allan. Journal of the Royal Statistical Society. Series A (General), 122(3):385–386, ???? 1959. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2342804.

Winsten:1959:BRB

[1001] C. B. Winsten. Book review: Introduction to Difference Equations, by Samuel Goldberg. Journal of the Royal Statistical Society. Series A (General), 122(3):386–387, ???? 1959. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2342805.

Corlett:1959:BRB

[1002] W. J. Corlett. Book review: An Introduction to Mathematics for Students of Economics, by J. Parry Lewis. Journal of the Royal Statistical Society. Series A (General), 122(3):387, ???? 1959. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342806.

Shankleman:1959:BRB

[1003] E. Shankleman. Book review: Modern Business Statistics, by J. E. Freund; F. J. Williams. Journal of the Royal Statistical Society. Series A (General), 122(3):388, ???? 1959. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342807.

Lomax:1959:BRB

[1004] K. S. Lomax. Book review: The Structure of British Industry: A Symposium, by Duncan Burn. Journal of the Royal Statistical Society. Series A (General), 122(3):389–390, ???? 1959. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2342808.

Marks:1959:BRB

[1005] M. Marks. Book review: The Industrial Efficiency of Rural Labour, by C. D. Harbury; A. D. Smith. Journal of the Royal Statistical Society. Series A (General), 122(3):390–391, ???? 1959. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342809.

Stringer:1959:BRB

[1006] J. Stringer. Book review: Efficiency in Government Through Systems Analysis, with Emphasis on Water Resources Development, by Roland N. McKean. Journal of the Royal Statistical Society. Series A (General), 122(3):391– 392, ???? 1959. CODEN JSSAEF.

ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2342810.

Clark:1959:BRB

[1007] Colin Clark. Book review: The Income of Nations: Theory, Measurement and Analysis, Past and Present; A Study in Applied Economics and Statistics, by Paul Studenski. Journal of the Royal Statistical Society. Series A (General), 122(3):392–393, ???? 1959. CODEN JSSAEF. ISSN 0035-9238. URL http:

                                        //www.jstor.org/stable/2342811.

Barr:1959:BRBa

[1008] A. Barr. Book review: Mental Illness in London, by Vera Norris. Journal of the Royal Statistical Society. Series A (General), 122(3):393–394, ???? 1959. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2342812.

Brooke:1959:BRB

[1009] Eileen M. Brooke. Book review: Economics of Mental Illness, by Rashi Fein. Journal of the Royal Statistical Society. Series A (General), 122(3): 394–395, ???? 1959. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2342813.

Gore:1959:BRB

[1010] A. T. Gore. Book review: Social and Biological Factors in Infant Mortality, by J. A. Heady; M. A. Heasman. Journal of the Royal Statistical Society. Series A (General), 122(3):395–396, ???? 1959. CODEN JSSAEF. ISSN 0035-

                                        9238. URL http://www.jstor.org/
                                        stable/2342814.

Thomas:1959:BRBb

[1011] C. J. Thomas. Book review: Census 1951. England and Wales. Fertility Report. Journal of the Royal Statistical Society. Series A (General), 122 (3):396–398, ???? 1959. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342815.

Anonymous:1959:SNc

[1012] Anonymous. Statistical notes. Journal of the Royal Statistical Society. Series A (General), 122(3):399–401, ???? 1959. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342816.

Anonymous:1959:CNb

[1013] Anonymous. Current notes. Journal of the Royal Statistical Society. Series A (General), 122(3):401–402, ???? 1959. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342817.

Yates:1959:WLS

[1014] F. Yates. Wilfred Leslie Stevens. Journal of the Royal Statistical Society. Series A (General), 122(3):403–404, ???? 1959. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342818.

Anonymous:1959:SEAc

[1015] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society. Series A (General), 122(3):405–412, ???? 1959. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2342819.

Anonymous:1959:LALc

[1016] Anonymous. List of additions to the library. Journal of the Royal Statistical Society. Series A (General), 122 (3):413–418, ???? 1959. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2342820.

Anonymous:1959:VIb

[1017] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (General), 122(4):577– 579, ???? 1959. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343074.

Stone:1959:EHU

[1018] P. A. Stone. The economics of housing and urban development. Journal of the Royal Statistical Society. Series A (General), 122(4):417–483, ???? 1959. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343075.

Youle:1959:SSI

[1019] P. V. Youle, K. D. Tocher, W. N. Jessop, and F. I. Musk. Simulation studies of industrial operations. Journal of the Royal Statistical Society. Series A (General), 122(4):484–510, ???? 1959. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343076.

Anonymous:1959:ARS

[1020] Anonymous. The 125th anniversary of the Royal Statistical Society. Journal of the Royal Statistical Society. Series A (General), 122(4):511–516, ???? 1959. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343077.

Anonymous:1959:RLG

[1021] Anonymous. Regulations of the local groups and of the industrial applications section. Journal of the Royal Statistical Society. Series A (General), 122 (4):517–521, ???? 1959. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343078.

Anonymous:1959:ARC

[1022] Anonymous. Annual report of the Council. Journal of the Royal Statistical Society. Series A (General), 122 (4):522–544, ???? 1959. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343079.

Anonymous:1959:POH

[1023] Anonymous. Proceedings of the one hundred and twenty-fifth Annual General Meeting of The Royal Statistical Society, held at the London School of Hygiene and Tropical Medicine, on Wednesday on, June 17th, 1959, at 5.30. Journal of the Royal Statistical Society. Series A (General), 122(4):m, ???? 1959. CODEN JSSAEF. ISSN

0035-9238. URL http://www.jstor.
org/stable/2343080.

Sampford:1959:BRB

[1024] M. R. Sampford. Book review: Statistical Methods in Biology, by N. T. J. Bailey. Journal of the Royal Statistical Society. Series A (General), 122 (4):546–547, ???? 1959. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343081.

Barr:1959:BRBb

[1025] A. Barr. Book review: Elementary Statistics with Applications in Medicine and the Biological Sciences, by Frederick E. Croxton. Journal of the Royal Statistical Society. Series A (General), 122(4):547, ???? 1959. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343082.

Cox:1959:BRBd

[1026] D. R. Cox. Book review: Elementary Decision Theory, by H. Chernoff; L. E. Moses. Journal of the Royal Statistical Society. Series A (General), 122 (4):547–548, ???? 1959. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343083.

vanRest:1959:BRB

[1027] E. D. van Rest. Book review: Quality Control and Industrial Statistics, by Acheson J. Duncan. Journal of the Royal Statistical Society. Series A (General), 122(4):548–549, ???? 1959. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343084.

Vajda:1959:BRBc

[1028] S. Vajda. Book review: Probability and Statistics for Business Decisions: An Introduction to Managerial Economics Under Uncertainty, by Robert Schlaifer. Journal of the Royal Statistical Society. Series A (General), 122 (4):549, ???? 1959. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

jstor.org/stable/2343085.

Vajda:1959:BRBd

[1029] S. Vajda. Book review: Operations Research: Methods and Problems, by M. Sasieni; A. Yaspan; L. Friedman. Journal of the Royal Statistical Society. Series A (General), 122(4):549–550, ????

1959. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343086.

Tocher:1959:BRBf

[1030] K. D. Tocher. Book review: Punched Cards: Their Applications to Science and Industry, by Robert S. Casey; James W. Perry; Madeline M. Berry; Allen Kent. Journal of the Royal Statistical Society. Series A (General), 122 (4):550–551, ???? 1959. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343087.

Tocher:1959:BRBg

[1031] K. D. Tocher. Book review: Handbook of Automation, Computation and Control. Vol. I. Control Fundamentals, by Eugene M. Grabbe; Simon Ramo; Dean E. Wooldridge. Journal of the Royal Statistical Society. Series A (General), 122(4):551, ???? 1959. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343088.

Vajda:1959:BRBe

[1032] S. Vajda. Book review: Strategy and Market Structure: Competition, Oligopoly, and the Theory of Games, by Martin Shubik. Journal of the Royal Statistical Society. Series A (General), 122(4):551–552, ???? 1959. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2343089.

Fisk:1959:BRB

[1033] P. R. Fisk. Book review: A Time Series Analysis of Interindustry Demands, by K. J. Arrow; M. Hoffenberg. Journal of the Royal Statistical Society. Series A (General), 122(4):552–553, ????

1959. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343090.

Cox:1959:BRBe

[1034] P. R. Cox. Book review: Social Theories of Fertility and the Malthusian Debate, by D. E. C. Eversley. Journal of the Royal Statistical Society. Series A (General), 122(4):553–554, ???? 1959. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343091.

Spicer:1959:BRB

[1035] C. C. Spicer. Book review: Smoking: The Cancer Controversy. Some Attempts to Assess the Evidence, by Ronald A. Fisher. Journal of the Royal Statistical Society. Series A (General), 122(4):554–556, ???? 1959. CODEN JSSAEF. ISSN 0035-9238. URL http:

//www.jstor.org/stable/2343092.

Martin:1959:BRBb

[1036] Leslie V. Martin. Book review: Tobacco Manufacturers’ Standing Committee Research Papers, No. 1. Statistics of Smoking, by G. F. Todd. Journal of the Royal Statistical Society. Series A (General), 122(4):556, ???? 1959. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343093.

Benjamin:1959:BRBb

[1037] B. Benjamin. Book review: Economics for Pleasure, by G. L. S. Shackle. Journal of the Royal Statistical Society. Series A (General), 122(4):556–557, ???? 1959. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343094.

Anonymous:1959:SNd

[1038] Anonymous. Statistical notes. Journal of the Royal Statistical Society. Series A (General), 122(4):558–560, ???? 1959. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343095.

Anonymous:1959:CNc

[1039] Anonymous. Current note. Journal of the Royal Statistical Society. Series A (General), 122(4):560, ???? 1959. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343096.

Anonymous:1959:SEAd

[1040] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society. Series A (General), 122(4):561–568, ???? 1959. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343097.

Anonymous:1959:LALd

[1041] Anonymous. List of additions to the library. Journal of the Royal Statistical Society. Series A (General), 122 (4):569–574, ???? 1959. CODEN JS-SAEF. ISSN 0035-9238. URL http://

www.jstor.org/stable/2343098.

Anonymous:1959:PRb

[1042] Anonymous. Periodical returns. Journal of the Royal Statistical Society. Series A (General), 122(4):575, ???? 1959. CODEN JSSAEF. ISSN 0035-

9238. URL http://www.jstor.org/
stable/2343099.

Muir:1960:BRB

[1043] A. Muir. Book review: Probability and Statistical Inference for Engineers: A First Course, by Cyrus Derman; Merton Klein. Journal of the Royal Statistical Society. Series A (General), 123 (2):203, ???? 1960. CODEN JSSAEF. ISSN 0035-9238. URL http://www.

                                        jstor.org/stable/2342660.

Bofinger:1961:NPW

[1044] Eve Bofinger and V. J. Bofinger. A note on the paper by W. E. Thomson on “Ernie — A Mathematical and Statistical Analysis”. Journal of the Royal Statistical Society. Series A (General), 124 (2):240–243, ???? 1961. CODEN JS-SAEF. ISSN 0035-9238. URL http://

                                        www.jstor.org/stable/2984133. See

[991].